### v3

# Theory of Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Using Super-Resolution Operators as Background Error Covariance Matrices

Abstract
This study proposes a theory of unsupervised super-resolution data assimilation (SRDA) using conditional variational autoencoders (CVAEs). This framework employs an evidence lower bound as the objective function and encompasses the three-dimensional variational (3D-Var) formalism. Similar to 3D-Var, it is essential to use a non-diagonal background error covariance matrix to assimilate distant observations. Instead of using such a non-diagonal matrix, we extend the proposed framework to employ the non-local nature of super-resolution operations. In the extended framework, super-resolution operators can serve as background error covariance matrices. Since the loss function for CVAEs is generally an evidence lower bound, SRDA can be performed using CVAEs. By incorporating the SR model into the CVAE, the encoder estimates the high-resolution (HR) analysis from the input of HR observations and low-resolution (LR) forecasts, while the decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis. The effectiveness of the proposed method was evaluated through numerical experiments using an idealized barotropic ocean jet system. Compared to inference by an ensemble Kalman filter, the proposed SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times because the SRDA does not require HR numerical integration or ensemble calculations. The findings of this study provide a theoretical basis for the fusion of SR and DA, which will stimulate further research on their integration.
preprint: AIP/123-QED
I Introduction
Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD).Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time. Uncertainties in model parameters can also lead to discrepancies from reality. DA reduces these errors by correcting forecast values or model parameters based on observational data, thereby being essential for accurate predictions.
Recent advances in deep learning are impacting various scientific fields and are also found in DA. One reason is that both DA and deep learning use backpropagation based on adjoint equations.Abarbanel, Rozdeba, and Shirman (2018); Chen et al. (2018); Geer (2021) Another reason stems from the use of the Bayesian approach.Carrassi et al. (2018); Bocquet et al. (2020); Geer (2021) In this approach, computational processes for target variables are modeled by probability distributions based on known data. For instance, in image generation,Kingma and Welling (2019); Croitoru et al. (2023) image data distributions are modeled and learned from existing images. Once trained, the neural network can generate a new image by sampling from the learned distribution. In DA, the true states of fluid systems are estimated through Bayesian modeling, namely by integrating known data, such as observations and fluid model outputs.
A typical neural network based on the Bayesian approach is the variational autoencoder (VAE).Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014) Like standard autoencoders,Dong et al. (2018) a VAE consists of an encoder and a decoder, both of which are implemented by neural networks. Typically, the encoder transforms the input data into a compressed representation in the latent space, while the decoder reconstructs the original data from these latent variables. The integration of Bayesian inference introduces two key characteristics into VAEs.Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014) First, a regularization term is added to the loss function, which structures the latent space. For instance, in a problem of generating handwritten digits, a specific dimension in the latent space may govern the digit styles.Kingma and Welling (2014); Rezende, Mohamed, and Wierstra (2014) Second, the VAE acts as a generative model and can produce new data by sampling a latent variable and decoding it. The Conditional VAE (CVAE) is an extension of VAEs that is based on conditional probability modeling.Kingma et al. (2014); Sohn, Lee, and Yan (2015) In the problem of handwritten digits, for example, CVAEs make it feasible to designate handwritten digits to be generated through this conditioning.Kingma et al. (2014); Sohn, Lee, and Yan (2015)
Applications of VAEs in DA can be categorized into two groups. The first group employs VAEs for dimensionality reduction.Liu, Grana, and de Figueiredo (2021); Mohd Razak et al. (2022); Melinc, Boštjan and Zaplotnik, Žiga (2024) In this approach, the encoder maps fluid model states into a low-dimensional latent space. DA methods, such as Kalman filters, are then applied in this space. Due to the reduced dimensionality, the computational cost is significantly lowered. This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks.Mack et al. (2020); Cheng et al. (2022, 2024) The second group employs VAEs as generators that learn the probability distribution of flow fields.Grooms (2021); Yang and Grooms (2021); Grooms et al. (2023) In this scheme, a fluid model state (i.e., a background state) is first encoded. Noise is then added in the latent space, creating various potential latent variables. Ensemble members are finally obtained by decoding these latent variables back into the physical space. This process facilitates efficient computation of covariance matrices used in DA methods, such as Kalman filters. A common characteristic in both groups is that VAEs do not directly perform DA; rather, VAEs are used to make DA methods more efficient. The present study proposes a new method, categorized into the third group, where VAEs directly perform DA.
In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state.Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) Typically, observations represent a state close to reality, but they are spatially sparse. In contrast, fluid model outputs (i.e., background states) include larger errors, but they are spatially dense. In DA, these background states are corrected by observations in a non-local manner using covariance matrices. For instance, when an observed value exists at a grid point, this observation is used to correct surrounding background values that are correlated with the background at that point.
The estimation of covariance matrices is also important for DA using autoencoders (including VAEs). In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown.Cheng et al. (2022); Melinc, Boštjan and Zaplotnik, Žiga (2024) More generally, the background error distributions are non-Gaussian, and such statistics can be estimated by VAEs.Xiao et al. (2024) The distributions learned by VAEs can be incorporated into a traditional DA method,Xiao et al. (2024) namely the three-dimensional variational (3D-Var)Lorenc (1986) formalism. In the present study, we focus on the non-local nature of super-resolution (SR) and utilize SR operators as covariance matrices.
Originally, SR is a technique to enhance the resolution of images and has been studied in computer vision as an application of neural networks.Ha et al. (2019); Anwar, Khan, and Barnes (2020) The success of such neural networks has led to numerous applications of SR in fluid dynamics,Ducournau and Fablet (2016); Vandal et al. (2017); Onishi, Sugiyama, and Matsuda (2019); Fukami, Fukagata, and Taira (2019); Wang et al. (2020, 2021); Yasuda et al. (2022); Yasuda, Onishi, and Matsuda (2023); McGibbon et al. (2024) as reviewed in a recent survey.Fukami, Fukagata, and Taira (2023) An important characteristic of SR is its non-local nature, where SR neural networks map low-resolution (LR) patterns with a certain spatial extent to high-resolution (HR) patterns.Dong et al. (2014); Gu and Dong (2021) Neural networks can focus on input LR patterns that are similar or spatially correlated to the target HR patterns, generating high-quality HR images using the non-local information in the input.Gu and Dong (2021)
Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA).Barthélémy et al. (2022); Yasuda and Onishi (2023) One reason for this integration is the increasing availability of HR observations.Imaoka et al. (2010); Durand et al. (2010); Cifelli et al. (2018) In typical DA methods, HR observations are assimilated into HR background states from HR fluid models,Li, Wang, and Fu (2019); Honda et al. (2022) which require a large amount of computational resources for time integration. In contrast, in SRDA frameworks, background states are computed with LR fluid models. The LR results are then super-resolved to HR using neural networks, and HR observations are assimilated into these HR backgrounds. SRDA methods are computationally efficient because they do not require numerical integration of HR fluid models. Since SRDA has recently been developed, several research questions remain, such as those regarding its theoretical background. For instance, the reasons for combining SR with DA are not clear, apart from the aforementioned reduction in computational costs.
This study proposes a theory of unsupervised SRDA using CVAEs and demonstrates that SR operators can act as background error covariance matrices. This result provides a theoretical foundation for combining SR with DA. We first review an evidence lower bound (ELBO) for DA (Section II), which serves as the loss function for unsupervised learning. We develop a new DA theory using the ELBO and show that this framework is an extension of the traditional 3D-Var method (Section III). This DA theory requires a non-diagonal background error covariance matrix, like the 3D-Var. We then extend the new DA framework by incorporating SR for non-local inference, without estimating such a non-diagonal matrix (Section IV). To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section V). The results are analyzed through comparison with an ensemble Kalman filter (Section VI). Conclusions are presented in Section VII. The findings of this study provide a theoretical basis for the fusion of SR and DA, potentially stimulating further research on their integration.
II Formulation of ELBO for DA
We reformulate an evidence lower bound,Ormerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al. (2017) which is the objective function in the proposed framework. ELBO is also known as the variational lower bound.
The present study considers assimilation at a single time step. In this context, $\bm{x}$, $\bm{y}$, and $\bm{z}$ represent a background state, observations, and the true state, respectively, where $\bm{x}\in\mathbb{R}^{n_{x}}$, $\bm{y}\in\mathbb{R}^{n_{y}}$, and $\bm{z}\in\mathbb{R}^{n_{z}}$: $\bm{x}$ is a prior forecast from an underlying CFD model, $\bm{y}$ represents spatially sparse observations, and $\bm{z}$ is the true state vector in the physical space. The objective of DA is to infer $\bm{z}$ from $\bm{x}$ and $\bm{y}$.
Consider the probabilistic model illustrated in Fig. 1, which is mathematically expressed as
[…]
The probability $p\left(\bm{y}\mid\bm{x}\right)$ is decomposed into $p\left(\bm{y}\mid\bm{z}\right)$ and $p\left(\bm{z}\mid\bm{x}\right)$: the former describes the probability of observing $\bm{y}$ when the true state is $\bm{z}$ and the latter is the prior distribution for $\bm{z}$ given the model forecast $\bm{x}$.
Figure 1: Probabilistic model expressed by Eq. (1). All variables are in the physical space.
The ELBO is derivedOrmerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al. (2017) by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$:
[…]
$=$ $\underbrace{\mathbb{E}_{q}\left[\ln p\left(\bm{y}\mid\bm{z}\right)\right]}_{\text{reconstruction error}}-\underbrace{\mathbb{E}_{q}\left[\ln\frac{q\left(\bm{z}\mid\bm{x},\bm{y}\right)}{p\left(\bm{z}\mid\bm{x}\right)}\right]}_{\text{KL divergence}}=:-l,$ (2c)
where $\mathbb{E}_{q}$ is the expectation operator with respect to $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$. Jensen’s inequality is applied in Eq. (2b), where the equality holds only when $q(\bm{z}\mid\bm{x},\bm{y})=p(\bm{z}\mid\bm{x},\bm{y})$, that is, when the approximate posterior is equal to the exact posterior. In Eq. (2c), the first term is referred to as the reconstruction error for $\bm{y}$, and the second term is known as the Kullback-Leibler (KL) divergence. The final line gives the ELBO ($=-l$), which is the lower bound for the log-likelihood, $\ln p\left(\bm{y}\mid\bm{x}\right)$.
DA is usually performed by maximum likelihood estimation.Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], the proposed framework maximizes the ELBO, or equivalently minimizes $l$ in Eq. (2c), to obtain $q(\bm{z}\mid\bm{x},\bm{y})$. The true state $\bm{z}$ is then inferred from $\bm{x}$ and $\bm{y}$ through $q(\bm{z}\mid\bm{x},\bm{y})$.
III DA using the ELBO under Gaussianity
We propose a theory of DA using the ELBO and show that this framework is an extension of the 3D-Var formalism. This result explains the necessity of non-diagonal background error covariance matrices, leading to the introduction of SR in Section IV.
III.1 Transformation of the ELBO under Gaussianity
The following Gaussian distributions are introduced to transform the ELBO:
$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\;\exp\left(-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}(\bm{x},\bm{y})\right\rVert_{V(\bm{x},\bm{y})^{-1}}^{2}\right),$ (3a)
$p\left(\bm{z}\mid\bm{x}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det B}}\;\exp\left(-\frac{1}{2}\left\lVert\bm{z}-\bm{x}\right\rVert_{B^{-1}}^{2}\right),$ (3b)
$p\left(\bm{y}\mid\bm{z}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{y}}\det R}}\exp\left(-\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{z}))\right\rVert_{R^{-1}}^{2}\right),$ (3c)
[…]
The ELBO is transformed by substituting Eqs. (3a)–(3c) into (2c). The first term, the reconstruction error in Eq. (2c), becomes
$\mathbb{E}_{q}\left[\ln p\left(\bm{y}\mid\bm{z}\right)\right]$ $=$ $\mathbb{E}_{q}\left[-\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{z})\right\rVert_{R^{-1}}^{2}-\frac{n_{y}}{2}\ln(2\pi)-\frac{1}{2}\ln\left(\det R\right)\right],$ (4a)
$\approx$ $-\frac{1}{n_{\text{sample}}}\frac{1}{2}\sum_{\text{sample}}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\text{const},$ (4b)
where
$\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\bm{a}+\bm{\varepsilon}^{\text{T}}V^{1/2}.$ (5)
Here, the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trickKingma and Welling (2014)). In the actual implementation, the sample size $n_{\text{sample}}$ can be set to $1$ if the batch size is sufficiently large.Kingma and Welling (2014) Each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$, where $\bm{a}$ and $V$ denote the vector and the matrix given by $\mathsf{A}(\bm{x},\bm{y})$ and $V(\bm{x},\bm{y})$, respectively, $\bm{\varepsilon}$ is an $n_{z}$-dimensional standard normal variable, and $V^{1/2}$ is obtained by Cholesky decomposition. Equation (4b) indicates that the reconstruction error measures the difference between $\bm{y}$ and $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
The KL divergence, the second term in Eq. (2c), is transformed as follows:
[…]
$=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}}}\int d\widetilde{\bm{z}}\;\exp\left(-\frac{\lVert\widetilde{\bm{z}}\rVert^{2}}{2}\right)\Bigg{\{}-\frac{1}{2}\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]$
$\quad+\frac{1}{2}\left\lVert\left[\widetilde{\bm{z}}^{\text{T}}V(\bm{x},\bm{y})^{1/2}+\mathsf{A}(\bm{x},\bm{y})\right]-\bm{x}\right\rVert_{B^{-1}}^{2}\Bigg{\}}+\text{const},$ (6b)
$=$ $\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\}+\text{const}.$ (6c)
To perform the integral, we use an auxiliary variable $\widetilde{\bm{z}}=V(\bm{x},\bm{y})^{-1/2}\left[\bm{z}-\mathsf{A}(\bm{x},\bm{y})\right]$. The first term in Eq. (6c) represents the difference between the analysis $\mathsf{A}(\bm{x},\bm{y})$ and the background state $\bm{x}$, while the second term is a convex function of $V(\bm{x},\bm{y})$, which takes the minimum when $V(\bm{x},\bm{y})=B$.
Using Eqs. (4b) and (6c), the ELBO in Eq. (2c) is transformed into
$l_{\text{Gauss}}=\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$ (7)
where constant terms are omitted and the sample size is set to 1 in Eq. (4b).
III.2 Reduction of the ELBO to the 3D-Var objective function
We show that $l_{\text{Gauss}}$ in Eq. (7) can be reduced to the 3D-Var objective function. This result suggests that DA can be performed by minimizing $l_{\text{Gauss}}$.
To show this result, we set the covariance $V(\bm{x},\bm{y})$ to a constant and quite small matrix. In this case, $\bm{\hat{z}}$ in Eq. (5) becomes $\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})=\bm{a}$ and the third term in Eq. (7) becomes constant. Therefore, $l_{\text{Gauss}}$ is reduced to $l_{\text{3D-Var}}$:
$l_{\text{3D-Var}}=\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{a})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\bm{a}-\bm{x}\right\rVert_{B^{-1}}^{2},$ (8)
where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from the inputs $\bm{x}$ and $\bm{y}$. The function $l_{\text{3D-Var}}$ is equivalent to the 3D-Var objective function,Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018); Lorenc (1986) and the vector $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$. The first term of $l_{\text{3D-Var}}$, which comes from the reconstruction error in Eq. (2c), compares the observation vector $\bm{y}$ and the analysis $\bm{a}$ through the observation operator $\mathsf{H}$. The second term of $l_{\text{3D-Var}}$, which comes from the KL divergence in Eq. (2c), measures the difference between the analysis $\bm{a}$ and the background state $\bm{x}$. By minimizing $l_{\text{3D-Var}}$, $\bm{y}$ and $\bm{x}$ are integrated, and the true state is estimated as $\bm{a}$. These results suggest that the ELBO describes the balance between observations and background states in estimating the true state.
The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is linear (i.e., $\mathsf{H}=H$, where $H$ is a matrix).Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018)
$\bm{a}$ $=$ $\bm{x}+K\left(\bm{y}-H\bm{x}\right),$ (9a)
$K$ $=$ $BH^{\text{T}}\left(R+HBH^{\text{T}}\right)^{-1},$ (9b)
where $K$ is known as the Kalman gain matrix. The second term in Eq. (9a) describes the non-local nature of DA. The background state $\bm{x}$ is corrected through the spatially sparse observations $\bm{y}$ by the non-diagonal matrix $K$. Typically, this non-diagonal property originates from the non-zero off-diagonal elements of the background error covariance matrix $B$. In other words, the non-local correction is realized through the spatial correlation patterns of background values. The necessity of spatial correlations leads to a difficulty for DA by neural networks (Section III.3).
III.3 Implementation of DA by CVAE
DA can be performed by solving the minimization problem of $l_{\text{Gauss}}$. However, this minimization at every assimilation time step would be computationally intensive. We propose to efficiently conduct this minimization using a trained CVAE.
Generally, the loss function for VAEs, including CVAEs, is an ELBO.Kingma and Welling (2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al. (2014); Sohn, Lee, and Yan (2015); Kingma and Welling (2019) This can be understood from the ELBO formula, Eq. (7). To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from a given input of $\bm{x}$ and $\bm{y}$; then, the observation vector $\bm{\hat{y}}$ ($=\mathsf{H}(\bm{\hat{z}})$) is reconstructed from $\bm{a}$ and $V$ via Eq. (5). The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$. The structure that reconstructs the input is common to autoencoders.Dong et al. (2018) The second and third terms in Eq. (7) are considered as the conditions for the estimated true state $\bm{\hat{z}}$. For instance, $l_{\text{Gauss}}$ is decreased as $\bm{a}$ becomes close to $\bm{x}$. Thus, a CVAE is necessary to incorporate these conditions, and this CVAE is trained by minimizing $l_{\text{Gauss}}$.
The CVAE consists of an encoder and a decoder (Fig. 2). The encoder estimates $\bm{a}$ and $V$ from the input $\bm{x}$ and $\bm{y}$, while the decoder reconstructs $\bm{\hat{y}}$. In other words, the encoder performs assimilation and the decoder acts as the nonlinear observation operator. The encoder learns how to integrate $\bm{x}$ and $\bm{y}$, yielding the analysis $\bm{a}$, while the decoder learns how to convert $\bm{a}$ into the observation space. Once the CVAE is trained, the encoder performs DA without minimizing $l_{\text{Gauss}}$ as long as the statistics of the background $\bm{x}$ and the observations $\bm{y}$ do not change significantly.
Figure 2: Architecture of the CVAE for DA. The encoder estimates the analysis vector $\bm{a}$ and its covariance $V$, both of which give a sample of the true state $\bm{\hat{z}}$ in Eq. (5). The decoder acts as the observation operator and reconstructs the observation vector $\bm{\hat{y}}=\mathsf{H}(\bm{\hat{z}})$.
The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters). Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices: $R=rI_{n_{y}}$ and $B=bI_{n_{z}}$, where $I_{n}$ is an identity matrix of size $n$, and $r$ and $b$ are positive real parameters. However, off-diagonal elements of $B$ describe the spatial correlation patterns of background states and are essential for the non-local estimation in DA (Section III.2).
There are several methods for implementing non-local estimation. For instance, latent variables can be transformed to have covariance or even follow non-Gaussian distributions by combining VAEs and normalizing flows,Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) where this transformation is learnable from data. An alternative approach is to divide observed values used as input to CVAEs and those targeted for reconstruction. CVAEs reconstruct observations from the other observations at different grid points, which enables CVAEs to learn non-locality. A similar technique is employed in MetNet-3, a surrogate neural network for weather predictions developed by Google.Andrychowicz et al. (2023) As a final method, we adopt super-resolution (SR). The non-locality is incorporated into CVAEs through SR even when $B$ and $R$ are diagonal.
IV SRDA using the ELBO
We incorporate SR into the DA theory proposed in Section III. A key characteristic of SR models is their non-locality:Dong et al. (2014); Gu and Dong (2021) SR models map LR patterns with a certain spatial extent to HR patterns. These SR mappings (i.e., SR operators) are integrated into DA, resulting in SRDA, which facilitates the non-local correction of background values by observations across distant grid points.
IV.1 Integration of SR into DA
To develop a theory of SRDA, we specify the resolutions of a background state $\bm{x}$, observations $\bm{y}$, and the true state $\bm{z}$: $\bm{x}$ is defined at LR grid points, while $\bm{y}$ and $\bm{z}$ are defined at HR grid points. The vector forms are obtained by arranging three-dimensional grid point values into columns. Observations are not available at all locations due to various factors, such as measurement errors. These missing grid points are excluded when arranging observations into the column vector $\bm{y}$.
We modify the probability distributions in Eqs. (3a)–(3c) for SRDA:
$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\exp\left\{-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\right\},$ (10a)
$p\left(\bm{z}\mid\bm{x}\right)$ $=$ $\frac{1}{(\sqrt{2\pi b})^{n_{z}}}\exp\left\{-\frac{1}{2b}\left\lVert\bm{z}-\mathsf{F}(\bm{x})\right\rVert^{2}\right\},$ (10b)
$p\left(\bm{y}\mid\bm{z}\right)$ $=$ $\frac{1}{(\sqrt{2\pi r})^{n_{y}}}\exp\left\{-\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{z})\right\rVert^{2}\right\},$ (10c)
where
$\mathsf{A}_{\text{HR}}=\mathsf{F}\circ\mathsf{A}_{\text{LR}}.$ (11)
[…]
$V(\bm{x},\bm{y})=\text{diag}\left[v_{1}(\bm{x},\bm{y}),\ldots,v_{n_{z}}(\bm{x},\bm{y})\right],$ (12)
where each variance $v_{i}(\bm{x},\bm{y})$ ($i=1,\ldots,n_{z}$) is a positive value determined by the vectors $\bm{x}$ and $\bm{y}$.
There are two essential changes from Eqs. (3) to (10). First, an SR operator $\mathsf{F}$ is introduced in Eq. (10b), where $\mathsf{F}$ maps an $n_{x}$-dimensional vector into an $n_{z}$-dimensional vector. Since $\bm{x}$ and $\bm{z}$ are defined at LR and HR grid points, respectively (i.e., $n_{x}<n_{z}$), this mapping performs SR. Comparing Eqs. (3b) and (10b), the mean of the prior distribution $p\left(\bm{z}\mid\bm{x}\right)$ is replaced with the super-resolved background state $\mathsf{F}(\bm{x})$. Second, the analysis state is also super-resolved using $\mathsf{F}$. The function $\mathsf{A}$ in Eq. (3a) is replaced with a composition $\mathsf{A}_{\text{HR}}=\mathsf{F}\circ\mathsf{A}_{\text{LR}}$ in Eq. (10a): the LR analysis is first calculated as $\bm{a}_{\text{LR}}=\mathsf{A}_{\text{LR}}(\bm{x},\bm{y})$ and then it is super-resolved as $\bm{a}_{\text{HR}}=\mathsf{F}(\bm{a}_{\text{LR}})$. The influences of these two changes are further discussed in Sections IV.2 and IV.3.
The ELBO is transformed by substituting Eqs. (10a)–(10c) into (2c), as in Section III.1:
$l_{\text{SRDA}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y}))-\mathsf{F}(\bm{x})\right\rVert^{2}+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$ (13)
where
$\bm{a}_{\text{HR}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y}))=\mathsf{F}(\mathsf{A}_{\text{LR}}(\bm{x},\bm{y})),$ (14a)
$\bm{a}_{\text{LR}}$ $=$ $\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}),$ (14b)
$\bm{\hat{z}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\mathsf{F}(\bm{a}_{\text{LR}})+\sum_{i=1}^{n_{z}}\varepsilon_{i}\sqrt{v_{i}}.$ (14c)
Eqs. (13) and (14c) corresponds to Eqs. (7) and (5) in Section III.1, respectively. Like in Eq. (7), the Monte Carlo method (i.e., the reparameterization trickKingma and Welling (2014)) is applied in Eq. (13) with the sample size being 1. In the right-hand side of Eq. (14c), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definitions of $\bm{a}_{\text{HR}}$ and $\bm{a}_{\text{LR}}$ are used. The first term of $l_{\text{SRDA}}$ represents the difference between the observation vector $\bm{y}$ and the estimated true state $\mathsf{H}(\bm{\hat{z}})$ in the observation space. The second term is the difference between the super-resolved analysis $\mathsf{F}(\bm{a}_{\text{LR}})$ and the super-resolved background $\mathsf{F}(\bm{x})$. The third term is a convex function of $v_{i}$, which takes the minimum when $v_{i}(\bm{x},\bm{y})=b$. These points indicate that the ELBO describes the balance between the observations and the super-resolved forecast in estimating the HR true state.
IV.2 Case of linear SR
We interpret $l_{\text{SRDA}}$ in Eq. (13) using a linear SR operator and show that SR operators act as background error covariance matrices. Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning.Park, Park, and Kang (2003) When generation processes of LR images from HR images are linear (e.g., subsampling and blurring), the inverse processes, namely the SR processes, are also linear.Park, Park, and Kang (2003) Here, an SR operator is denoted as a matrix $F$, which has the size of $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) and maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors. Since SR is generally non-local, the matrix $F$ has non-zero off-diagonal elements.
To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section III.2. The estimated true state $\bm{\hat{z}}$ in Eq. (14c) is then equal to $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$. Further, the third term in Eq. (13) is ignored because it becomes constant.
Thus, $l_{\text{SRDA}}$ is reduced to
$l_{\text{linear SRDA}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(F\bm{a}_{\text{LR}})\right\rVert^{2}+\frac{1}{2b}\left(\bm{a}_{\text{LR}}-\bm{x}\right)^{\text{T}}\left(F^{\text{T}}F\right)\left(\bm{a}_{\text{LR}}-\bm{x}\right).$ (15)
The matrix $F^{\text{T}}F$ is an $n_{x}\times n_{x}$ positive semi-definite matrix. Compared to the ELBO for 3D-Var in Eq. (8), $F^{\text{T}}F$ can be interpreted as the covariance matrix $B$ and describes the spatial correlations among the LR components.
When the observation operator is linear ($\mathsf{H}=H$) as in Section III.2, the minimizer for $l_{\text{linear SRDA}}$ is always obtained due to the convexity of $l_{\text{linear SRDA}}$. We first consider the case where $F^{\text{T}}F$ is positive definite (i.e., invertible). The minimizer is then derived, as in Eq. (9):
$\bm{a}_{\text{LR}}$ $=$ $\bm{x}+\widetilde{K}\left(\bm{y}-\widetilde{H}\bm{x}\right),$ (16a)
$\widetilde{K}$ $=$ $\widetilde{B}\widetilde{H}^{\text{T}}\left(rI_{n_{y}}+\widetilde{H}\widetilde{B}\widetilde{H}^{\text{T}}\right)^{-1},$ (16b)
where $\widetilde{B}=b\left(F^{\text{T}}F\right)^{-1}$ and $\widetilde{H}=HF$. The observation operator for the LR background is given by $\widetilde{H}=HF$ through the SR operator $F$. Generally, the Kalman gain matrix $\widetilde{K}$ has non-zero off-diagonal elements due to $F$ being non-diagonal. Thus, the LR background $\bm{x}$ is non-locally corrected by the observations $\bm{y}$, giving the LR analysis $\bm{a}_{\text{LR}}$ in Eq. (16a).
Even when $F^{\text{T}}F$ is positive semi-definite (i.e., at least one eigenvalue is zero), minimizers for $l_{\text{linear SRDA}}$ exist due to its convexity. In this case, a solution similar to Eq. (16a) is derived using pseudo-inverse matrices (not shown in detail).
IV.3 Case of non-linear SR
In practice, SR models are neural networks, so SR operators are non-linear. Since the nonlinear case includes the linear case, the non-local inference is achieved in both cases. Here, we briefly examine non-linear SR operators from a different perspective: the transformation of probability distributions.
We consider a complete non-linear SR operator $\mathsf{F}$, that is, a bijective SR operator. Rigorously speaking, SR is an inverse problem and SR operations are not bijective; however, a virtual bijective map is sometimes considered in theoretical studies of SR. In this case, any HR true state $\bm{z}$ has the corresponding LR state: $\bm{z}_{\text{LR}}=\mathsf{F}^{-1}(\bm{z})$ and $\bm{z}=\mathsf{F}(\bm{z}_{\text{LR}})$. We can regard the posterior and prior distributions in Eqs. (10a) and (10b) as non-Gaussian distributions for $\bm{z}_{\text{LR}}$:
$q\left(\bm{z}_{\text{LR}}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{|D\mathsf{F}|}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\exp\left\{-\frac{1}{2}\left\lVert\mathsf{F}(\bm{z}_{\text{LR}})-\mathsf{F}(\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}))\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\right\},$ (17a)
$p\left(\bm{z}_{\text{LR}}\mid\bm{x}\right)$ $=$ $\frac{|D\mathsf{F}|}{(\sqrt{2\pi b})^{n_{z}}}\exp\left\{-\frac{1}{2b}\left\lVert\mathsf{F}(\bm{z}_{\text{LR}})-\mathsf{F}(\bm{x})\right\rVert^{2}\right\},$ (17b)
where $|D\mathsf{F}|$ denotes the Jacobian determinant of $\mathsf{F}$.
This result suggests that SR operators can be interpreted as normalizing flows.Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) Normalizing flow is a deep generative model that converts Gaussian variables into non-Gaussian variables, generating new data that obey complicated probability distributions. In normalizing flows, invertible non-linear transformations are implemented as neural-network layers and are stacked to form the whole network to achieve complex transformations. Importantly, normalizing flows can be used to transform latent variables in VAEs.Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) This combined architecture is similar to the CVAE proposed in the present study (Section IV.5), where the SR model transforms $\bm{a}_{\text{LR}}$ into $\bm{a}_{\text{HR}}$ inside the encoder. This similarity implies the potential for using normalizing flows to further develop SRDA.
IV.4 Loss function for SR models
So far, we have employed the SR operator $\mathsf{F}$, namely the trained SR model. This section presents a loss function for training this model. We do not know the true state of the system and only know the LR background $\bm{x}$ and the HR observations $\bm{y}$. The former $\bm{x}$ is spatially complete, while the latter $\bm{y}$ is spatially incomplete. The SR model is trained in this setup.
We propose the following loss function to train SR models:
$l_{\text{SR}}=\sum_{i=1}^{n_{y}}\left|y_{i}-[P\mathsf{F}(\bm{x})]_{i}\right|,$ (18)
where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space. The right-hand side clarifies that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies.Ha et al. (2019); Anwar, Khan, and Barnes (2020); Fukami, Fukagata, and Taira (2023) Since $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$, backpropagation can be performed; that is, the SR model $\mathsf{F}$ can be trained in a supervised manner, where the input is $\bm{x}$ and the target is $\bm{y}$. When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$, as shown in Hastie et al.Hastie, Tibshirani, and Friedman (2009) This fact suggests that the SR model can make unbiased inferences if the median of the observation error distribution is zero.
IV.5 Implementation of SRDA by CVAE
The efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, similar to Section III.3. Thus, SRDA can be implemented using CVAEs.
The CVAE for SRDA consists of an encoder and a decoder (Fig. 3). The difference from the CVAE in Fig. 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously. Here, we employ the SR model described in Section IV.4, which is trained with the loss function in Eq. (18). This encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$ and estimates the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$. The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ by calculating the variances. The decoder acts as the observation operator, as in Fig. 2. The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ as the loss function.
Figure 3: Architecture of the CVAE for SRDA. The encoder estimates the HR analysis vector $\bm{a}_{\text{HR}}$ using the SR model. The encoder also estimates the covariance $V$ for $\bm{a}_{\text{HR}}$. The variables $\bm{a}_{\text{HR}}$ and $V$ give a sample of the true state $\bm{\hat{z}}$ in Eq. (14c). The decoder acts as the observation operator and reconstructs the observation vector $\bm{\hat{y}}=\mathsf{H}(\bm{\hat{z}})$.
Strictly speaking, a different SR model should be used within the encoder because the SR model in Section IV.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$. This difference in input may introduce bias to $\bm{a}_{\text{LR}}$. In other words, the trained SR model assumes $\bm{x}$ as input and may lead to a bias of $\bm{a}_{\text{LR}}$ similar to that of $\bm{x}$. On the other hand, the output $\bm{a}_{\text{HR}}$ ($=\mathsf{F}(\bm{a}_{\text{LR}})$) is expected to be less biased because the SR-model output is unbiased if the observation error distribution has a zero median (Section IV.4). Thus, $\bm{a}_{\text{HR}}$ is used as the initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section IV.6). One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model during the training of the CVAE. For simplicity, however, the present study does not use such an advanced learning method, which remains a future research topic.
IV.6 Sequential SRDA by the CVAE
We introduce a sequential SRDA method applicable to CFD simulations. The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model. The forecast and assimilation are executed alternately, as in other statistical DA methods,Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018) such as 3D-Var. This assimilation is performed using a trained CVAE’s encoder (Fig. 3).
Figure 4 shows a schematic of the sequential SRDA. Here, we assume a constant assimilation interval $\Delta T$. At every assimilation time step, the encoder assimilates the HR observations $\bm{y}$ into the LR forecast state $\bm{x}$, thereby estimating the HR analysis $\bm{a}_{\text{HR}}$. This output is resized to the LR using an algebraic method, such as linear interpolation. In the next assimilation cycle, the resized analysis is used as the initial condition for the LR fluid model, and forecasts are computed by numerical integration over $\Delta T$. The forecast at the final time step is fed to the encoder to infer the subsequent HR analysis.
Figure 4: Schematic of the sequential SRDA. “HR obs” denotes HR observations. A constant assimilation interval $\Delta T$ is assumed, where the output time step is a quarter of $\Delta T$ as an example.
The decoder is not used in the sequential SRDA; rather, it is used only in the training phase to learn the observation operator (Section IV.5). In contrast, the trained SR model can be used in the test phase (i.e., the operational phase) to super-resolve LR forecast states. This SR operation is independent of assimilation, allowing for the generation of HR forecast states at any time step (Section VI.3).
V Methods
We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs).Fletcher (2023a) The ground truth (or the nature run) is generated by an HR fluid model, from which pseudo (or synthetic) observations are computed. The SRDA employs an LR version of the fluid model, and the resultant forecast and analysis are compared with the ground truth. A simple jet stream system was adopted for these experiments, as a proof of concept for the SRDA.
V.1 Fluid simulation
An idealized barotropic ocean jet, proposed by David et al.,David, Marshall, and Zanna (2017) was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi.Yasuda and Onishi (2023) The experimental setup is briefly explained here. See the above references for further details.
The governing equations in the periodic channel $(x,y)$ are as followsDavid, Marshall, and Zanna (2017)
$\frac{\partial\omega}{\partial t}+\bm{u\cdot\nabla}\omega+\beta\frac{\partial\psi}{\partial x}$ $=$ $-\kappa\omega-\nu\Delta^{2}\omega-\frac{d\tau(y)}{dy},$ (19a)
$\Delta\psi$ $=$ $\omega,$ (19b)
$\bm{u}$ $=$ $\left(-\frac{\partial\psi}{\partial y},\frac{\partial\psi}{\partial x}\right),$ (19c)
where $x$ is the channel-wise direction and $y$ is the transverse direction ($x\in[0,2\pi]$ and $y\in[0,\pi]$). Equation (19a) describes the evolution of vorticity $\omega$ over time $t$. The left-hand side represents the advection of $\omega$ and the beta effect,Vallis (2017a) while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$. Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively.
The parameters in Eq. (19a) were set as follows: $\beta=0.1$, $\kappa=1\times 10^{-2}$, and $\nu=1\times 10^{-5}$. The wind stress was given by
$\tau(y)=\tau_{0}\left[{\text{sech}}^{2}\left(\frac{y-y_{0}}{\delta}\right)-c\right],$ (20)
where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero. This setup falls within the parameter regime of mixing barriers with strong eddies,David, Marshall, and Zanna (2017) where coherent vortices persist in a statistically steady state (Section VI.1).
The initial condition was a zonal jet superimposed with random perturbations. The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion,Vallis (2017b) which is a necessary condition for barotropic instability. Perturbations were added to the vorticity field for each wavenumber component, with phase and amplitude drawn from Gaussian distributions. The simulation data for deep learning were generated by varying these random perturbations.
Numerical integration was conducted using the modified Euler and pseudo-spectral methods at two spatial resolutions: LR and HR. Table 1 provides detailed configurations for the LR and HR. All simulations ran from $t=0$ to $20$ with output time intervals of $0.25$. The CFD models using the LR and HR configurations are referred to as the LR and HR fluid models, respectively.
Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR. The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule).Canuto et al. (1988)
Name Grid size ($x\times y$) Cutoff wavenumber Integration time step
LR (low resolution) $32\times 16$ 10 $5\times 10^{-4}$
HR (high resolution) $128\times 64$ 42 $\frac{1}{4}\times 5\times 10^{-4}$
V.2 Synthetic observations
Synthetic observations were generated from the HR simulations. We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere.Lin and Yang (2020); Fletcher (2023b) Observed grid points were randomly subsampled at constant spatial intervals, namely $8\times 8$ grids, which means that 1.56% ($=1/8^{2}$) of grid points are observed. To demonstrate that the SRDA can address changes in observed points, the subsampled points were varied in time by random shifts. For the unobserved points, Not a Numbers (NaNs) were assigned and replaced with missing values in the subsequent preprocessing (Section V.3). To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity. The noise has a mean of 0 and a standard deviation of 0.1. This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
V.3 Neural networks
We outline the network architectures of an SR model and a CVAE for the sequential SRDA (Section IV.6). Hyperparameters are provided in Appendix and the full implementation is available at our GitHub repository (see Data Availability Statement). PyTorch 1.11.0Paszke et al. (2019) was used for the implementation.
The SR model architecture is shown by the top panel in Fig. 5. This model resembles the neural network for SR known as VDSR, proposed by Kim et al.Kim, Lee, and Lee (2016) The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR. First, this input is nonlinearly transformed through a stack of convolution and ReLU layers. Upsampling is then executed by two blocks that include pixel shuffle layers.Shi et al. (2016) Each block doubles the size of each dimension, resulting in a fourfold increase in resolution. Additionally, a skip connectionHe et al. (2016) is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output. This skip connection allows the SR model to capture small-scale patterns, that is, discrepancies between the LR input and the HR output.
Figure 5: Network architectures of the SR model and the CVAE. The latter consists of the encoder and the decoder. The label “Res Block” refers to a residual block, and “Res Block x6” indicates the stacking of six Res Blocks. As in the typical implementation of VAEs,Kingma and Welling (2019) the variance output is regarded as $\ln V$ because neural networks can generally output both positive and negative values. Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp[1/2\ln V]$ provides a realization of the true state $\bm{\hat{z}}$ in Eq. (14c).
The CVAE consists of an encoder and a decoder (Fig. 5). The encoder performs SRDA and integrates the inputs, an LR forecast and an HR observation, into the HR analysis. The HR observation is first downsampled through convolution to align the resolution to the LR (totally, downsampled by a factor of 4). The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$. Then, $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$. The variance for $\bm{a}_{\text{HR}}$ is denoted as $V$ in Eq. (12) and is also estimated by the encoder. Although $V$ and $\bm{a}_{\text{HR}}$ can be treated independently in the encoder (Section IV.5), the present study computes $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of neural network parameters. Moreover, we employed the feature extractor in the trained SR model to compute $V$ (Feature Extractor-1 in Fig. 5).
The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$. This operation is achieved through a stack of convolutional and Leaky ReLUMaas et al. (2013) layers. We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), although this replacement slightly decreases the inference accuracy. This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section V.2); that is, the observation operator is linear. Although the proposed SRDA can theoretically address nonlinear observation operators (Section IV), such a nonlinear case is not examined in this proof of concept, which remains a topic for future research.
In preprocessing, the vorticity $\omega$ at each grid point is transformed by
${\text{clip}}_{[0,1]}\left(\frac{\omega-m_{\omega}}{s_{\omega}}\right).$ (21)
The clipping function, ${\text{clip}}_{[0,1]}(z)=\min\{1,\max\{0,z\}\}$ ($z\in\mathbb{R}$), restricts the value range to the interval $[0,1]$. The parameters $m_{\omega}$ and $s_{\omega}$ are determined such that 99.9% of the vorticity values fall within the $[0,1]$ range. After applying Eq. (21), zero is assigned to the HR grid points without observations, and zero is regarded as a missing value. This clear interpretation of zero facilitates feature extraction from input data.Yasuda, Onishi, and Matsuda (2023)
V.4 Training and testing methods for the neural networks
V.4.1 Training method
We adopted an offline strategy in which training data are pre-generated [Fig. 6(a)]. The training data consist of pairs of HR observations and LR forecasts, as discussed in Section IV.5. The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain. After time integration over $\Delta T=1$, the resultant LR forecasts were used in training. This $\Delta T$ equals the assimilation interval, as explained below, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet. The corresponding observations were obtained by subsampling and adding Gaussian noise (Section V.2). Figure 6(a) denotes pairs of the LR forecast and the HR observations as stars. By varying the random initial perturbations (Section V.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
Figure 6: Schematics of time series for (a) training and (b) testing phases. The assimilation interval is set to $\Delta T=1$, while the output interval (i.e., the forecast interval) is $\Delta T/4=0.25$.
In principle, no HR simulation is required because the training data include LR forecasts rather than HR results (Sections IV.4 and IV.5). These LR forecasts are obtained by numerical integration in which the analysis states are used as initial conditions. In offline training, however, such an analysis is unavailable because the CVAE is not trained yet. In realistic applications, these initial conditions can be given from existing assimilated data, such as reanalysis,Hersbach et al. (2020); KOSAKA et al. (2024) or online training will be effective,Rasp (2020) where neural-network training and data generation are executed alternately.
The SR model was trained, followed by training the CVAE. During the CVAE training, all parameters in the SR model were fixed, as discussed in Section IV.5. The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq. (18) and $l_{\text{SRDA}}$ in Eq. (13), respectively. The CVAE and the SR model were trained using the Adam optimizerKingma and Ba (2015) with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128. Each training was terminated using early stopping with a patience parameter of 50 epochs. The parameters in $l_{\text{SRDA}}$ were set to $r=1.2\times 10^{-5}$ and $b=1.0\times 10^{-3}$. The value of $r$ was given by the variance of the measurement errors (Section V.2), while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO. Without this balancing, the performance of the CVAE would deteriorate. This phenomenon is often observed in VAEs,Higgins et al. (2017); Alemi et al. (2018) and the loss balancing can be optimized during training.Asperti and Trentin (2020) For simplicity, however, such an advanced technique was not employed in the present study.
V.4.2 Testing method
The sequential SRDA was tested using the method described in Section IV.6. Figure 6(b) shows a schematic of the test phase. The only difference between Figs. 4 and 6(b) is that observations are generated by subsampling the HR simulation results. Due to this difference, the HR results are regarded as the ground truth, and the SRDA can be evaluated by comparing its inference to the ground truth. We newly conducted 500 HR fluid simulations. For each case, the LR fluid model was initialized with the low-pass filtered HR vorticity at $t=0$; the SRDA was subsequently conducted until $t=20$ with the assimilation interval $\Delta T=1$. Without SRDA, the vorticity evolution would differ from the ground truth, not only due to the difference in the initial condition (i.e., all high-wavenumber components are removed), but also due to the difference in the fluid model resolution (i.e., LR versus HR).
Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth. The mean absolute error (MAE) ratio quantifies pixel-wise errors.
MAE $\text{MAE ratio}=\frac{\sum_{i}|\omega_{i}-\hat{\omega}_{i}|}{\sum_{i}|\omega_{i}|},$ (22)
where the summation is taken over all grid points, $\omega_{i}$ denotes the ground-truth vorticity at the $i$-th grid point, and $\hat{\omega}_{i}$ is the corresponding inference. The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns:Wang et al. (2004)
MSSIM loss $=$ $1-{\text{MSSIM}},$ (23)
$=$ $1-\sum_{i}\frac{\left(2\mu_{i}\hat{\mu}_{i}+{\text{C}_{1}}^{2}\right)\left(2\gamma_{i}+{\text{C}_{2}}^{2}\right)}{\left(\mu_{i}^{2}+\hat{\mu}_{i}^{2}+{\text{C}_{1}}^{2}\right)\left(\sigma_{i}^{2}+\hat{\sigma}_{i}^{2}+{\text{C}_{2}}^{2}\right)},$
where ${\text{C}_{1}}=0.01$, ${\text{C}_{2}}=0.03$, $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the ground truth, respectively, $\hat{\mu}_{i}$ and $\hat{\sigma}_{i}^{2}$ are the corresponding quantities for the inference, and $\gamma_{i}$ is the covariance between the ground truth and the inference. The variables in the summation are calculated locally in space using a Gaussian filter. The MSSIM loss takes a value greater than or equal to 0; a smaller value indicates spatial patterns more similar to the ground truth. A detailed discussion of MSSIM can be found in Wang et al.Wang et al. (2004)
The MAE ratio and MSSIM loss were calculated at each time step in each test simulation. The values averaged across all test simulations are referred to by the same names. Both MAE ratio and MSSIM loss are collectively called the test errors.
V.5 Baseline model using Ensemble Kalman Filter (EnKF)
As a comparison to the SRDA, we employed an ensemble Kalman filter (EnKF)Evensen (1994) for the LR fluid model, using the perturbed observation method.Burgers, van Leeuwen, and Evensen (1998) The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations. The uncertainty with the EnKF is compared with that with the SRDA.
Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, where the assimilation interval $\Delta T=1$ is the same as in the SRDA. Barthélémy et al.Barthélémy et al. (2022) argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points. Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn.Gaspari and Cohn (1999) These error covariances were inflated by adding Gaussian noise to the analysis, prior to the forecast process.Whitaker et al. (2008) This Gaussian noise incorporates spatial correlation by estimating its covariance from the training data at each time step. The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution. We confirmed that the results are not sensitive to the statistics of the Gaussian noise for inflation. Indeed, similar results were obtained when the Gaussian noise had a constant correlation length.
Hyperparameters for the EnKF were adjusted using the training data to minimize the MAE. Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations. The full implementation of the EnKF is also available at our GitHub repository (see Data Availability Statement).
VI Results and discussion
VI.1 Time evolution of vorticity in the jet system
Before discussing the SRDA results, we describe the time evolution of HR vorticity in the jet system and confirm the sensitivity to initial conditions (i.e., a positive Lyapunov exponent).
Figure 7 shows a typical evolution of HR vorticity. Initially ($t=0$), an unstable jet exists around the $y$ center with small perturbations. The jet starts meandering and breaks down into multiple vortices, accompanied by the development of fine filaments ($1\lesssim t\lesssim 4$). Subsequently, vortex merging occurs, intensifying small-scale structures again ($4\lesssim t\lesssim 12$). Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$). This final state is statistically steady and can be understood by Rossby wave dynamics.David, Marshall, and Zanna (2017)
Figure 7: Typical vorticity evolution from the HR ground-truth data.
We confirmed that the maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$. This suggests that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability of the initial conditions. Indeed, without DA, the deviation from the ground truth continues to increase over time, while the SRDA effectively suppresses this error growth (Section VI.3).
VI.2 Accuracy of SR
We first confirm the accuracy of inference by the SR model, as the SRDA is based on this model. Figure 8 shows three sets of snapshots for the ground truth, observations, LR input, and SR output. The results indicate that the SR model reproduces the fine filaments and small-scale structures within vortices. These patterns are not evident in the LR inputs or observations.
Figure 8: Vorticity snapshots for the ground truth, observations, LR input, and SR output. The SR model was trained using the observation data as target. The input data are not indexed by time because each sample was randomly selected for testing without simulating time evolution.
Flow-field reconstruction has been extensively studied using neural networks.Lluís et al. (2020); Maulik et al. (2020); Gundersen et al. (2021); Wang et al. (2022) In these studies, inputs are incomplete, such as spatially sparse sensor measurements, where only a portion of the entire domain is available. Neural networks are then trained by comparing their output to complete target data. The setup of our problem is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
The success of the SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution. The loss function $l_{\text{SR}}$ in Eq. (18) operates at individual grid points, and only the non-missing observed points contribute to $l_{\text{SR}}$. This property enables backpropagation even when the target data are missing, as discussed in Section IV.4. Furthermore, since convolution shares its kernel across the space (known as weight sharing),Aloysius and Geetha (2017) the efficient learning was realized using the small amount of non-missing observational data.
Although the inference appears to be accurate, we confirmed that the SR model cannot adequately correct forecast errors. One might think that the SR output could be used as the analysis due to its high accuracy (Fig. 8). However, we found that test errors accumulated over time when the downsampled SR output was used as the initial condition for the LR fluid model. This result indicates that observations need to be assimilated into forecasts.
VI.3 Accuracy of LR forecast by SRDA
We demonstrate that the sequential SRDA can make more accurate inferences than the EnKF. The LR forecast is compared in this section, and the HR analysis is examined next (Section VI.4).
Figure 9 shows snapshots of the ground truth and forecasts for the EnKF and the SRDA. To align the resolution, a low-pass filter was applied to the HR ground-truth vorticity. The figure also includes LR simulation results without SR or DA. During the jet collapse ($t=2$), differences in the inferences are not obvious, likely because the LR fluid model cannot well emulate fine filaments. During the vortex merging ($t=9$), the SRDA most accurately describes the vortex-split distribution near the center. The EnKF also captures this pattern, but the SRDA inference is closer to the ground truth. Without SR or DA, the vortex-split structure is not reproduced. In the steady state ($t=16$), both EnKF and SRDA accurately infer the vorticity phase. Without SR or DA, the vortex shapes and locations still differ from the ground truth.
Figure 9: Vorticity snapshots of the ground truth and forecasts for the EnKF and the SRDA at $t=2$, 9, and 16. For comparison, the figure also shows LR simulation results without SR or DA. A low-pass filter was applied to the ground-truth snapshots to align the resolution to the LR.
The quantitative analysis further confirms the accuracy of the SRDA forecast. The top panels in Fig. 10 show the time series of the mean values for the MAE ratio and the MSSIM loss. Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (i.e., the positive Lyapunov exponent in Section VI.1). In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines). The error reduction occurs every time unit since the assimilation interval was set to 1. When averaged over time, both time series have two peaks, corresponding to the two developments of small-scale patterns by the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). The bottom panels in Fig. 10 show the differences in the test errors between the SRDA and the EnKF, where the gray areas represent the 10th to 90th percentile values. The test errors of the SRDA are significantly smaller than those of the EnKF. Thus, we can conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
Figure 10: Time series of the MAE ratio and the MSSIM loss for the forecast states in the 500 test simulations. The top panels show the mean test errors at each time step with an output interval of 0.25. The bottom panels show the differences between the SRDA and the EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
Within the SRDA framework, the LR forecast can be super-resolved to the HR at an arbitrary time by the SR model (Section IV.6). Figure 11 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast. The merging of two vortices is not obvious in the LR, but is well captured in the super-resolved forecast.
Figure 11: Typical vortex merging in the ground truth, the LR forecasts, and the super-resolved forecast.
VI.4 Accuracy of HR analysis by SRDA
We next demonstrate the accuracy of the HR analysis provided by the SRDA. Figure 12 shows snapshots of the ground truth and analyses. During the jet collapse ($t=2$), the filament structure appears much more distinct in the SRDA analysis than in the EnKF. Furthermore, during the vortex merging ($t=9$ and 11), the SRDA snapshots exhibit sharper vortex shapes and clearer internal structures compared to the EnKF.
Figure 12: Vorticity snapshots of the ground truth and analyses for the EnKF and the SRDA at $t=2$, 9, and 11.
The quantitative analysis also confirms the accuracy of the SRDA analysis. Figure 13 shows the time series of the MAE ratio and the MSSIM loss, as in Fig. 10. The dots represent the assimilation time steps, with assimilation intervals at 1. The SRDA and EnKF analyses are inferred only at these times. The EnKF time series (orange dashed lines) show two peaks, corresponding to the two developments of small-scale patterns linked to the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). These peaks are less apparent in the SRDA time series (blue solid lines), especially for the first peak. This result suggests that the SRDA effectively mitigates the error increase during the jet collapse. The bottom panels in Fig. 13 indicate that the MAE ratio and the MSSIM loss for the SRDA are significantly smaller than those of the EnKF. Thus, we can conclude that the SRDA analysis is more accurate than the EnKF analysis, in line with the forecast results (Section VI.3).
Figure 13: Time series of the MAE ratio and the MSSIM loss for the analysis states in the 500 test simulations. These states were inferred only at the analysis time steps, denoted by dots. The top panels show the mean test errors. The bottom panels show the differences between the SRDA and the EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
We compare uncertainties estimated by the EnKF and the SRDA. Here, standard deviation (SD) is used as an indicator of uncertainty. For the SRDA, SD was directly calculated by taking the square root of the estimated variances in Eq. (12), while for the EnKF, SD at each grid point was calculated from all ensemble members (i.e., 300 members).
Figure 14 compares distributions of the analysis values and their SDs for the EnKF and the SRDA. The SRDA SD tends to be large around vortex filaments or vortex edges. This tendency appears to be most evident in the jet collapse ($t=2$). The EnKF SD also exhibits a similar tendency; however, the SD distribution is blurred, likely reflecting the blurred vortex patterns. The SD value ranges also differ: the EnKF SD is between 0.3 and 0.9, while the SRDA SD is between 0.8 and 0.9. Related to this range difference, the SRDA tends to overestimate SD in regions where vorticity is close to zero, which can be found at $t=9$ and 11 in Fig. 14. In contrast, the EnKF SD tends to be small in small-vorticity regions. This overestimation of the SRDA may be due to a simplified variance estimation in the present CVAE (Section V.3), where the variances are computed using the SR model from the LR analysis. Therefore, although the SRDA demonstrated the potential for uncertainty quantification without relying on ensemble calculation, further model development is necessary. Recent studies demonstrate that VAEs can learn error distributions of model forecasts,Grooms (2021); Yang and Grooms (2021); Grooms et al. (2023); Xiao et al. (2024) and their model architectures may be effective for the SRDA.
Figure 14: Vorticity snapshots of the analyses and their standard deviations (SDs) for the EnKF and the SRDA at $t=2$, 9 and 11. The analyses shown here are the same as in Fig. 12. To compare them with their SDs, these panels are displayed again.
VI.5 Computation time of SRDA
Lastly, we confirm that the computational time of the SRDA is much less than that of the EnKF. The elapsed times for the simulations from $t=0$ to 20 were measured using a single process on a CPU (Intel Xeon Gold 6326). For the SRDA, in addition to the assimilation by the encoder, the LR forecast was super-resolved by the SR model at every output time step ($\Delta t=0.25$). The average wall time was 21.3 s for the SRDA, considerably less than 205.7 s for the EnKF. This difference is mainly due to the calculation of the ensemble evolution, which is not required in the SRDA.
VII Conclusions
This study has proposed a theory of unsupervised SRDA using the CVAE. We have shown that a new DA framework using the ELBO encompasses the 3D-Var formalism (Section III). Since ELBO serves as the loss function for CVAEs, the assimilation can be executed using the CVAE. In this process, it is essential to use a non-diagonal background error covariance matrix to incorporate the spatial correlations of forecast states into the CVAE, similar to the 3D-Var. Instead of using such non-diagonal matrices, we have further extended the framework by using the non-local nature of SR operations to achieve DA (Section IV). In other words, SR operators can serve as background error covariance matrices. This result provides a theoretical foundation for combining SR and DA.
The combination of SR models and CVAEs enables the assimilation of HR observations into LR forecasts that are computed with LR fluid models. This approach eliminates the need for numerical integration using HR fluid models, thereby reducing computational burden. Specifically, the CVAE’s encoder estimates the HR analysis from the input of LR forecasts and HR observations, while the decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis. The resulting HR analysis is then downsampled to the LR using algebraic methods, such as linear interpolation, serving as the initial condition for the LR fluid model in the next assimilation cycle.
The effectiveness of the proposed method was evaluated through numerical experiments using an idealized barotropic ocean jet system (Sections V and VI). Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times. These results suggest the potential of SRDA as an efficient approach for integrating observations into CFD simulations.
There are several future research topics, as SRDA has been recently developed. Although the proposed theory enables learning of nonlinear observation operators, the proof of concept in the present study has examined only the linear observation operator. Furthermore, advanced training methods, such as online trainingRasp (2020) and loss balancing for VAEs,Asperti and Trentin (2020) will be important to develop SRDA, whereas a simple offline training method was employed in our study (Section V.4). These points will be addressed in a more realistic experimental setup. In practical applications, data partitioning will be important because the data size for three-dimensional flows is larger, which would require the division of snapshots into small patches.Ouala et al. (2018); Yang and Grooms (2021) Non-Gaussianity is also important in more realistic DA, which has recently been addressed by VAEsXiao et al. (2024) and normalizing flows.Xie et al. (2023) The theory proposed in our study will provide a foundation for using such advanced techniques, leading to further development of SRDA.
Acknowledgements.
This work used computational resources of the TSUBAME3.0 supercomputer provided by the Institute of Science Tokyo through the HPCI System Research Project (Project ID: hp220102) to conduct the CFD simulations. The deep learning was performed on the Earth Simulator system (Project IDs: 1-23007 and 1-24009) at the Japan Agency for Marine-Earth Science and Technology (JAMSTEC). This paper is based on results obtained from a project, JPNP22002, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
Data Availability Statement
The data and source code that support the findings of this study are preserved and openly available at https://github.com/YukiYasuda2718/srda-cvae.
Appendix A Hyperparameters of the SR model and the CVAE
We list here the hyperparameters of the SR model and the CVAE shown in Fig. 5. Table 2 lists the hyperparameters for the convolution layers in the SR model and the CVAE. The kernel size of all convolutions is 3. The slope of all leaky ReLUMaas et al. (2013) is $-0.01$. The upscale factor of all pixel shuffle operationsShi et al. (2016) is 2. The resizing in the SR model is performed by bicubic interpolation. The weight parameters of all convolution layers are randomly initialized using uniform distributions.He et al. (2015)
Table 2: Hyperparameters for the convolution layers in the SR model and the CVAE. The layer names are defined in Fig. 5.
Layer name Input channels Output channels Stride
Conv2D-0 1 128 1
Conv2D-1 128 128 1
Conv2D-2 128 512 1
Conv2D-3 128 256 1
Conv2D-4 256 1024 1
Conv2D-5 256 256 1
Conv2D-6 256 1 1
Conv2D-7 1 32 1
Conv2D-8 256 1 1
Conv2D-9 256 1 1
Conv2D-10 1 8 1
Conv2D-11 8 8 1
Conv2D-12 8 1 1
Conv2D-13 1 112 2
Conv2D-14 112 112 1
Conv2D-15 112 224 2
Conv2D-16 224 224 1
Conv2D-17 256 256 1

### v4

@@
-Theory of Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Using Super-Resolution Operators as Background Error Covariance Matrices
+Theory of Unsupervised Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Estimating Background Covariances via Super-Resolution
 Abstract
 This study proposes a theory of unsupervised super-resolution data assimilation (SRDA) using conditional variational autoencoders (CVAEs).
-This framework employs an evidence lower bound as the objective function and encompasses the three-dimensional variational (3D-Var) formalism.
+The theory is based on an evidence lower bound that leads to an objective function for unsupervised learning.
+This function can be reduced to the objective function of a traditional data assimilation method, namely the three-dimensional variational (3D-Var) formalism.
 Similar to 3D-Var, it is essential to use a non-diagonal background error covariance matrix to assimilate distant observations.
-Instead of using such a non-diagonal matrix, we extend the proposed framework to employ the non-local nature of super-resolution operations.
-In the extended framework, super-resolution operators can serve as background error covariance matrices.
-Since the loss function for CVAEs is generally an evidence lower bound, SRDA can be performed using CVAEs.
-By incorporating the SR model into the CVAE, the encoder estimates the high-resolution (HR) analysis from the input of HR observations and low-resolution (LR) forecasts, while the decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis.
-The effectiveness of the proposed method was evaluated through numerical experiments using an idealized barotropic ocean jet system.
-Compared to inference by an ensemble Kalman filter, the proposed SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times because the SRDA does not require HR numerical integration or ensemble calculations.
-The findings of this study provide a theoretical basis for the fusion of SR and DA, which will stimulate further research on their integration.
-preprint: AIP/123-QED
-I Introduction
-Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD).Asch, Bocquet, and Nodet (2016); Carrassi et al.
-(2018) Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time.
-Uncertainties in model parameters can also lead to discrepancies from reality.
-DA reduces these errors by correcting forecast values or model parameters based on observational data, thereby being essential for accurate predictions.
+Instead of using such a non-diagonal matrix, we extend the proposed theory by using the non-locality of super-resolution (SR).
+For linear SR, SR operators serve as background error covariance matrices.
+For nonlinear SR, error backpropagation through SR neural networks induces covariance structures in inference.
+SRDA can be performed using CVAEs because the loss function for CVAEs is generally an evidence lower bound.
+By incorporating the SR neural network into the CVAE, the encoder estimates the high-resolution (HR) analysis from HR observations and low-resolution (LR) forecasts.
+The decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis.
+The effectiveness of SRDA was evaluated through numerical experiments using an idealized barotropic ocean jet system.
+Compared to inference with an ensemble Kalman filter, SRDA demonstrated superior accuracy in HR inference.
+SRDA was also computationally efficient because it does not require HR numerical integration or ensemble calculations.
+The findings of this study provide a theoretical basis for integrating SR and DA, which will stimulate further research in this direction.
+keywords:
+super-resolution , data assimilation , unsupervised learning , evidence lower bound , background error covariance , variational autoencoder
+journal: Comput.
+Methods Appl.
+Mech.
+Engrg.corresponding_author_infocorresponding_author_infofootnotetext: Corresponding author.
+2-12-1-i9-111 Ookayama, Meguro-ku, Tokyo, 1528550, Japan.
+Tel.: +81 3 57343173 (office).
+Email address: yasuda.yuki@scrc.iir.isct.ac.jp
+[…]
+[inst1] organization=Supercomputing Research Center, Institute of Integrated Research, Institute of Science Tokyo, addressline=2-12-1 Ookayama, Meguro-ku, city=Tokyo, postcode=1528550, country=Japan
+1 Introduction
+Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD) [1, 2].
+Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time.
+Uncertainties in model parameters may also lead to discrepancies from reality.
+DA reduces these errors by correcting forecast values or model parameters based on observational data, thereby being essential for accurate predictions, such as numerical weather prediction.
 Recent advances in deep learning are impacting various scientific fields and are also found in DA.
-One reason is that both DA and deep learning use backpropagation based on adjoint equations.Abarbanel, Rozdeba, and Shirman (2018); Chen et al.
-(2018); Geer (2021) Another reason stems from the use of the Bayesian approach.Carrassi et al.
-(2018); Bocquet et al.
-(2020); Geer (2021) In this approach, computational processes for target variables are modeled by probability distributions based on known data.
-For instance, in image generation,Kingma and Welling (2019); Croitoru et al.
-(2023) image data distributions are modeled and learned from existing images.
-Once trained, the neural network can generate a new image by sampling from the learned distribution.
-In DA, the true states of fluid systems are estimated through Bayesian modeling, namely by integrating known data, such as observations and fluid model outputs.
-A typical neural network based on the Bayesian approach is the variational autoencoder (VAE).Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014) Like standard autoencoders,Dong et al.
-(2018) a VAE consists of an encoder and a decoder, both of which are implemented by neural networks.
+One reason is that both DA and deep learning use backpropagation based on adjoint equations [3, 4, 5].
+Another reason is the use of the Bayesian approach [2, 5, 6].
+This approach models the true states of fluid systems by probability distributions based on known data, such as observations and fluid model outputs.
+A typical neural network using the Bayesian approach is the variational autoencoder (VAE) [7, 8, 9].
+Like standard autoencoders [10], a VAE consists of an encoder and a decoder, both of which are implemented by neural networks.
 Typically, the encoder transforms the input data into a compressed representation in the latent space, while the decoder reconstructs the original data from these latent variables.
-The integration of Bayesian inference introduces two key characteristics into VAEs.Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014) First, a regularization term is added to the loss function, which structures the latent space.
-For instance, in a problem of generating handwritten digits, a specific dimension in the latent space may govern the digit styles.Kingma and Welling (2014); Rezende, Mohamed, and Wierstra (2014) Second, the VAE acts as a generative model and can produce new data by sampling a latent variable and decoding it.
-The Conditional VAE (CVAE) is an extension of VAEs that is based on conditional probability modeling.Kingma et al.
-(2014); Sohn, Lee, and Yan (2015) In the problem of handwritten digits, for example, CVAEs make it feasible to designate handwritten digits to be generated through this conditioning.Kingma et al.
-(2014); Sohn, Lee, and Yan (2015)
+A major extension of VAEs is the conditional VAE (CVAE), which is based on conditional probability modeling [11, 12].
+This conditioning allows us to model target variables, such as the true states of systems, conditioned on known data, such as observations.
 Applications of VAEs in DA can be categorized into two groups.
-The first group employs VAEs for dimensionality reduction.Liu, Grana, and de Figueiredo (2021); Mohd Razak et al.
-(2022); Melinc, Boštjan and Zaplotnik, Žiga (2024) In this approach, the encoder maps fluid model states into a low-dimensional latent space.
-DA methods, such as Kalman filters, are then applied in this space.
-Due to the reduced dimensionality, the computational cost is significantly lowered.
-This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks.Mack et al.
-(2020); Cheng et al.
-(2022, 2024) The second group employs VAEs as generators that learn the probability distribution of flow fields.Grooms (2021); Yang and Grooms (2021); Grooms et al.
-(2023) In this scheme, a fluid model state (i.e., a background state) is first encoded.
-Noise is then added in the latent space, creating various potential latent variables.
-Ensemble members are finally obtained by decoding these latent variables back into the physical space.
-This process facilitates efficient computation of covariance matrices used in DA methods, such as Kalman filters.
-A common characteristic in both groups is that VAEs do not directly perform DA; rather, VAEs are used to make DA methods more efficient.
-The present study proposes a new method, categorized into the third group, where VAEs directly perform DA.
-In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state.Asch, Bocquet, and Nodet (2016); Carrassi et al.
-(2018) Typically, observations represent a state close to reality, but they are spatially sparse.
+The first group employs VAEs for dimensionality reduction [13, 14, 15].
+In this approach, the encoder maps fluid model states into a low-dimensional latent space.
+Statistical DA methods, such as Kalman filters, are then applied in this space.
+Due to the reduced dimensionality, the computational cost for DA is significantly lowered.
+This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks [16, 17, 18, 19, 20, 21].
+The second group employs VAEs as generators that learn the probability distributions of flow fields [22, 23, 24].
+In this approach, a fluid model state is encoded, and noise is then added in the latent space, creating various potential latent variables.
+Ensemble members are obtained by decoding these latent variables back into the physical space.
+This process facilitates fast computation of covariance matrices used in statistical DA methods.
+A common characteristic in both groups is that VAEs do not directly perform DA; rather, VAEs are used to make statistical DA methods more efficient.
+We proposes a new method, categorized into the third group, where VAEs directly perform DA.
+In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state [1, 2].
+Generally, observations represent a state close to reality, but they are spatially sparse.
 In contrast, fluid model outputs (i.e., background states) include larger errors, but they are spatially dense.
-In DA, these background states are corrected by observations in a non-local manner using covariance matrices.
-For instance, when an observed value exists at a grid point, this observation is used to correct surrounding background values that are correlated with the background at that point.
+DA methods correct these background states with observations in a non-local manner using covariance matrices.
+For instance, when an observed value exists at a grid point, this observation is used to correct surrounding (i.e., slightly distant) background values that are correlated with the background at that point.
+Hereafter, the term “non-local” means that an inferred value at a grid point depends on input values in the neighborhood (i.e., at slightly distant grid points).
+Conversely, “local” means that an inferred value is computed only from the input values at the same point.
 The estimation of covariance matrices is also important for DA using autoencoders (including VAEs).
-In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown.Cheng et al.
-(2022); Melinc, Boštjan and Zaplotnik, Žiga (2024) More generally, the background error distributions are non-Gaussian, and such statistics can be estimated by VAEs.Xiao et al.
-(2024) The distributions learned by VAEs can be incorporated into a traditional DA method,Xiao et al.
-(2024) namely the three-dimensional variational (3D-Var)Lorenc (1986) formalism.
-In the present study, we focus on the non-local nature of super-resolution (SR) and utilize SR operators as covariance matrices.
-Originally, SR is a technique to enhance the resolution of images and has been studied in computer vision as an application of neural networks.Ha et al.
-(2019); Anwar, Khan, and Barnes (2020) The success of such neural networks has led to numerous applications of SR in fluid dynamics,Ducournau and Fablet (2016); Vandal et al.
-(2017); Onishi, Sugiyama, and Matsuda (2019); Fukami, Fukagata, and Taira (2019); Wang et al.
-(2020, 2021); Yasuda et al.
-(2022); Yasuda, Onishi, and Matsuda (2023); McGibbon et al.
-(2024) as reviewed in a recent survey.Fukami, Fukagata, and Taira (2023) An important characteristic of SR is its non-local nature, where SR neural networks map low-resolution (LR) patterns with a certain spatial extent to high-resolution (HR) patterns.Dong et al.
-(2014); Gu and Dong (2021) Neural networks can focus on input LR patterns that are similar or spatially correlated to the target HR patterns, generating high-quality HR images using the non-local information in the input.Gu and Dong (2021)
-Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA).Barthélémy et al.
-(2022); Yasuda and Onishi (2023) One reason for this integration is the increasing availability of HR observations.Imaoka et al.
-(2010); Durand et al.
-(2010); Cifelli et al.
-(2018) In typical DA methods, HR observations are assimilated into HR background states from HR fluid models,Li, Wang, and Fu (2019); Honda et al.
-(2022) which require a large amount of computational resources for time integration.
+In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown [15, 19].
+More generally, the background error distributions are non-Gaussian and can be estimated by VAEs [25].
+The non-Gaussian distributions learned by VAEs are then incorporated into a traditional DA method [25], namely, the three-dimensional variational (3D-Var [26]) formalism.
+In the present study, we focus on the non-locality of super-resolution (SR) and utilize SR operators for estimating covariance structures.
+SR is originally a technique to enhance the resolution of images.
+In computer vision, deep learning-based SR has been actively studied [27, 28, 29].
+A key characteristic of SR is its non-local nature [30, 31, 32, 33]: neural networks focus on input pixels at low resolution (LR) that are distant from a target pixel at high resolution (HR), but correlated to this target, generating high-quality HR images.
+Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA) [34, 35].
+One reason for this integration is the increasing availability of HR observations [36, 37, 38].
+In typical DA methods, HR observations are assimilated using HR fluid models [39, 40], which require a large amount of computational resources for time integration.
 In contrast, in SRDA frameworks, background states are computed with LR fluid models.
-The LR results are then super-resolved to HR using neural networks, and HR observations are assimilated into these HR backgrounds.
+The LR results are then super-resolved to HR using neural networks while assimilating HR observations into these super-resolved backgrounds.
 SRDA methods are computationally efficient because they do not require numerical integration of HR fluid models.
-Since SRDA has recently been developed, several research questions remain, such as those regarding its theoretical background.
-For instance, the reasons for combining SR with DA are not clear, apart from the aforementioned reduction in computational costs.
-This study proposes a theory of unsupervised SRDA using CVAEs and demonstrates that SR operators can act as background error covariance matrices.
+One advantage of combining SR with DA is that advanced techniques developed in SR can be applied in DA.
+Numerous SR studies have been conducted not only in computer vision [27, 28, 29] but also in fluid dynamics [41].
+In contrast, DA using deep learning has only recently become active [42].
+As for SRDA, several research questions remain, such as those regarding its theoretical background.
+Indeed, previous studies on SRDA [34, 35] did not utilize the non-locality of SR.
+If theoretical backgrounds for combining SR with DA are clarified, further integration of both will be promoted.
+This study proposes a theory of unsupervised SRDA using CVAEs and shows that SR operators can be used to estimate background covariances.
 This result provides a theoretical foundation for combining SR with DA.
-We first review an evidence lower bound (ELBO) for DA (Section II), which serves as the loss function for unsupervised learning.
-We develop a new DA theory using the ELBO and show that this framework is an extension of the traditional 3D-Var method (Section III).
-This DA theory requires a non-diagonal background error covariance matrix, like the 3D-Var.
-We then extend the new DA framework by incorporating SR for non-local inference, without estimating such a non-diagonal matrix (Section IV).
-To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section V).
-The results are analyzed through comparison with an ensemble Kalman filter (Section VI).
-Conclusions are presented in Section VII.
-The findings of this study provide a theoretical basis for the fusion of SR and DA, potentially stimulating further research on their integration.
-II Formulation of ELBO for DA
-We reformulate an evidence lower bound,Ormerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al.
-(2017) which is the objective function in the proposed framework.
-ELBO is also known as the variational lower bound.
+We first review an evidence lower bound (ELBO) for DA (Section 2), which serves as the loss function for unsupervised learning.
+We then develop a new theory of DA using the ELBO and show that this theory is an extension of the 3D-Var (Section 3).
+Similar to the 3D-Var, our theory requires non-diagonal background error covariance matrices.
+We further extend our theory by incorporating SR for non-local inference, which eliminates the need to estimate non-diagonal covariance matrices (Section 4).
+To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section 5).
+The results are analyzed through comparison with an ensemble Kalman filter (Section 6).
+Conclusions are presented in Section 7.
+2 Review of ELBO for DA
+We reformulate an evidence lower bound (ELBO) [43, 44, 45], which is the objective function in the proposed theory. ELBO is also known as the variational lower bound.
 The present study considers assimilation at a single time step.
-In this context, $\bm{x}$, $\bm{y}$, and $\bm{z}$ represent a background state, observations, and the true state, respectively, where $\bm{x}\in\mathbb{R}^{n_{x}}$, $\bm{y}\in\mathbb{R}^{n_{y}}$, and $\bm{z}\in\mathbb{R}^{n_{z}}$: $\bm{x}$ is a prior forecast from an underlying CFD model, $\bm{y}$ represents spatially sparse observations, and $\bm{z}$ is the true state vector in the physical space.
+In this context, $\bm{x}$, $\bm{y}$, and $\bm{z}$ represent a background state, observations, and the true state, respectively, where $\bm{x}\in\mathbb{R}^{n_{x}}$, $\bm{y}\in\mathbb{R}^{n_{y}}$, and $\bm{z}\in\mathbb{R}^{n_{z}}$.
+Specifically, $\bm{x}$ is a prior forecast from an underlying CFD model, $\bm{y}$ represents spatially sparse observations, and $\bm{z}$ is the true state vector in the physical space.
 The objective of DA is to infer $\bm{z}$ from $\bm{x}$ and $\bm{y}$.
@@
 […]
-The probability $p\left(\bm{y}\mid\bm{x}\right)$ is decomposed into $p\left(\bm{y}\mid\bm{z}\right)$ and $p\left(\bm{z}\mid\bm{x}\right)$: the former describes the probability of observing $\bm{y}$ when the true state is $\bm{z}$ and the latter is the prior distribution for $\bm{z}$ given the model forecast $\bm{x}$.
+The probability $p\left(\bm{y}\mid\bm{x}\right)$ is decomposed into two probability distributions: $p\left(\bm{y}\mid\bm{z}\right)$ describes the probability of observing $\bm{y}$ when the true state is $\bm{z}$, and $p\left(\bm{z}\mid\bm{x}\right)$ is the prior distribution for $\bm{z}$ given the model forecast $\bm{x}$.
 Figure 1: Probabilistic model expressed by Eq. (1). All variables are in the physical space.
-The ELBO is derivedOrmerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al. (2017) by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$:
+The ELBO is derived by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$ [43, 44, 45]:
 […]
@@
 (2c), the first term is referred to as the reconstruction error for $\bm{y}$, and the second term is known as the Kullback-Leibler (KL) divergence.
-The final line gives the ELBO ($=-l$), which is the lower bound for the log-likelihood, $\ln p\left(\bm{y}\mid\bm{x}\right)$.
-DA is usually performed by maximum likelihood estimation.Asch, Bocquet, and Nodet (2016); Carrassi et al.
-(2018) Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], the proposed framework maximizes the ELBO, or equivalently minimizes $l$ in Eq.
-(2c), to obtain $q(\bm{z}\mid\bm{x},\bm{y})$.
+The final line gives the ELBO ($=-l$), which is a lower bound for the log-likelihood, $\ln p\left(\bm{y}\mid\bm{x}\right)$.
+DA is usually performed via maximum likelihood estimation [1, 2].
+Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], we maximize the ELBO, or equivalently minimize $l$ in Eq.
+(2c) to obtain $q(\bm{z}\mid\bm{x},\bm{y})$.
 The true state $\bm{z}$ is then inferred from $\bm{x}$ and $\bm{y}$ through $q(\bm{z}\mid\bm{x},\bm{y})$.
-III DA using the ELBO under Gaussianity
-We propose a theory of DA using the ELBO and show that this framework is an extension of the 3D-Var formalism.
-This result explains the necessity of non-diagonal background error covariance matrices, leading to the introduction of SR in Section IV.
-III.1 Transformation of the ELBO under Gaussianity
-The following Gaussian distributions are introduced to transform the ELBO:
+3 DA using the ELBO under Gaussianity
+Starting from the ELBO, we derive a new objective function that can be used to train CVAEs in an unsupervised manner.
+This loss function requires non-diagonal background error covariance matrices, similar to those in 3D-Var.
+3.1 Transformation of the ELBO under Gaussianity
+To transform the ELBO, the following Gaussian distributions are introduced:
 $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\;\exp\left(-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}(\bm{x},\bm{y})\right\rVert_{V(\bm{x},\bm{y})^{-1}}^{2}\right),$ (3a)
 $p\left(\bm{z}\mid\bm{x}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det B}}\;\exp\left(-\frac{1}{2}\left\lVert\bm{z}-\bm{x}\right\rVert_{B^{-1}}^{2}\right),$ (3b)
-$p\left(\bm{y}\mid\bm{z}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{y}}\det R}}\exp\left(-\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{z}))\right\rVert_{R^{-1}}^{2}\right),$ (3c)
+$p\left(\bm{y}\mid\bm{z}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{y}}\det R}}\exp\left(-\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{z})\right\rVert_{R^{-1}}^{2}\right),$ (3c)
 […]
-The ELBO is transformed by substituting Eqs. (3a)–(3c) into (2c). The first term, the reconstruction error in Eq. (2c), becomes
+The ELBO is transformed by substituting Eqs. (3a)–(3c) into Eq. (2c). The first term, the reconstruction error in Eq. (2c), becomes
 $\mathbb{E}_{q}\left[\ln p\left(\bm{y}\mid\bm{z}\right)\right]$ $=$ $\mathbb{E}_{q}\left[-\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{z})\right\rVert_{R^{-1}}^{2}-\frac{n_{y}}{2}\ln(2\pi)-\frac{1}{2}\ln\left(\det R\right)\right],$ (4a)
@@
 $\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\bm{a}+\bm{\varepsilon}^{\text{T}}V^{1/2}.$ (5)
-Here, the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trickKingma and Welling (2014)).
-In the actual implementation, the sample size $n_{\text{sample}}$ can be set to $1$ if the batch size is sufficiently large.Kingma and Welling (2014) Each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$, where $\bm{a}$ and $V$ denote the vector and the matrix given by $\mathsf{A}(\bm{x},\bm{y})$ and $V(\bm{x},\bm{y})$, respectively, $\bm{\varepsilon}$ is an $n_{z}$-dimensional standard normal variable, and $V^{1/2}$ is obtained by Cholesky decomposition.
+Here, $\bm{a}$ and $V$ denote the vector and matrix given by $\mathsf{A}(\bm{x},\bm{y})$ and $V(\bm{x},\bm{y})$, respectively, $\bm{\varepsilon}$ is an $n_{z}$-dimensional standard normal variable, and $V^{1/2}$ is obtained by Cholesky decomposition.
+In Eq.
+(4b), the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trick [8]), where each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ as in Eq.
+(5).
+In the actual implementation, the sample size $n_{\text{sample}}$ can be set to 1 if the batch size is sufficiently large [8].
 Equation (4b) indicates that the reconstruction error measures the difference between $\bm{y}$ and $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
 The KL divergence, the second term in Eq. (2c), is transformed as follows:
+$\mathbb{E}_{q}$ $\left[\ln\frac{q\left(\bm{z}\mid\bm{x},\bm{y}\right)}{p\left(\bm{z}\mid\bm{x}\right)}\right]$
 […]
@@
 $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}}}\int d\widetilde{\bm{z}}\;\exp\left(-\frac{\lVert\widetilde{\bm{z}}\rVert^{2}}{2}\right)\Bigg{\{}-\frac{1}{2}\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]$
-$\quad+\frac{1}{2}\left\lVert\left[\widetilde{\bm{z}}^{\text{T}}V(\bm{x},\bm{y})^{1/2}+\mathsf{A}(\bm{x},\bm{y})\right]-\bm{x}\right\rVert_{B^{-1}}^{2}\Bigg{\}}+\text{const},$ (6b)
-$=$ $\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\}+\text{const}.$ (6c)
-To perform the integral, we use an auxiliary variable $\widetilde{\bm{z}}=V(\bm{x},\bm{y})^{-1/2}\left[\bm{z}-\mathsf{A}(\bm{x},\bm{y})\right]$.
+$\quad+\frac{1}{2}\left\lVert\left[V(\bm{x},\bm{y})^{1/2}\widetilde{\bm{z}}+\mathsf{A}(\bm{x},\bm{y})\right]-\bm{x}\right\rVert_{B^{-1}}^{2}\Bigg{\}},$ (6b)
+$=$ $\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\}.$ (6c)
+We introduce an auxiliary variable $\widetilde{\bm{z}}=V(\bm{x},\bm{y})^{-1/2}\left[\bm{z}-\mathsf{A}(\bm{x},\bm{y})\right]$ to perform the integral.
+For brevity, constant terms are omitted in Eqs.
+(6b) and (6c).
 The first term in Eq.
-(6c) represents the difference between the analysis $\mathsf{A}(\bm{x},\bm{y})$ and the background state $\bm{x}$, while the second term is a convex function of $V(\bm{x},\bm{y})$, which takes the minimum when $V(\bm{x},\bm{y})=B$.
-Using Eqs. (4b) and (6c), the ELBO in Eq. (2c) is transformed into
-$l_{\text{Gauss}}=\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$ (7)
-where constant terms are omitted and the sample size is set to 1 in Eq. (4b).
-III.2 Reduction of the ELBO to the 3D-Var objective function
-We show that $l_{\text{Gauss}}$ in Eq. (7) can be reduced to the 3D-Var objective function. This result suggests that DA can be performed by minimizing $l_{\text{Gauss}}$.
-To show this result, we set the covariance $V(\bm{x},\bm{y})$ to a constant and quite small matrix.
-In this case, $\bm{\hat{z}}$ in Eq.
-(5) becomes $\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})=\bm{a}$ and the third term in Eq.
+(6c) represents the difference between the analysis $\mathsf{A}(\bm{x},\bm{y})$ and the background state $\bm{x}$, while the second term is a convex function of $V(\bm{x},\bm{y})$ that takes the minimum when $V(\bm{x},\bm{y})=B$.
+By substituting Eqs. (4b) and (6c) into Eq. (2c), we obtain the ELBO under Gaussianity:
+$l_{\text{Gauss}}$ $=$ $\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}$ (7)
+$\quad+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$
+where constant terms are omitted and the sample size $n_{\rm sample}$ in Eq. (4b) is set to 1.
+3.2 Reduction of the ELBO to the 3D-Var objective function
+We show that $l_{\text{Gauss}}$ can be reduced to the 3D-Var objective function. This suggests that DA can be performed by minimizing $l_{\text{Gauss}}$.
+We start by setting the covariance $V(\bm{x},\bm{y})$ to a constant matrix, where each component is quite small.
+The variable $\bm{\hat{z}}$ in Eq.
+(5) is then approximated as $\bm{\hat{z}}=\mathsf{A}(\bm{x},\bm{y})=\bm{a}$, and the third term in Eq.
 (7) becomes constant.
-Therefore, $l_{\text{Gauss}}$ is reduced to $l_{\text{3D-Var}}$:
+Consequently, $l_{\text{Gauss}}$ is reduced to the 3D-Var objective function $l_{\text{3D-Var}}$:
 $l_{\text{3D-Var}}=\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{a})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\bm{a}-\bm{x}\right\rVert_{B^{-1}}^{2},$ (8)
-where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from the inputs $\bm{x}$ and $\bm{y}$.
-The function $l_{\text{3D-Var}}$ is equivalent to the 3D-Var objective function,Asch, Bocquet, and Nodet (2016); Carrassi et al.
-(2018); Lorenc (1986) and the vector $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$.
+where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from $\bm{x}$ and $\bm{y}$.
+This $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$.
 The first term of $l_{\text{3D-Var}}$, which comes from the reconstruction error in Eq.
-(2c), compares the observation vector $\bm{y}$ and the analysis $\bm{a}$ through the observation operator $\mathsf{H}$.
+(2c), represents the difference between the observation vector $\bm{y}$ and the analysis $\bm{a}$ via the observation operator $\mathsf{H}$.
 The second term of $l_{\text{3D-Var}}$, which comes from the KL divergence in Eq.
-(2c), measures the difference between the analysis $\bm{a}$ and the background state $\bm{x}$.
-By minimizing $l_{\text{3D-Var}}$, $\bm{y}$ and $\bm{x}$ are integrated, and the true state is estimated as $\bm{a}$.
-These results suggest that the ELBO describes the balance between observations and background states in estimating the true state.
-The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is linear (i.e., $\mathsf{H}=H$, where $H$ is a matrix).Asch, Bocquet, and Nodet (2016); Carrassi et al. (2018)
+(2c), represents the difference between the analysis $\bm{a}$ and the background state $\bm{x}$.
+By minimizing $l_{\text{3D-Var}}$, $\bm{x}$ and $\bm{y}$ are integrated to estimate the true state as $\bm{a}$.
+Since $l_{\text{3D-Var}}$ is derived from the ELBO [i.e., $l_{\text{Gauss}}$ in Eq.
+(7)], the ELBO describes the balance between $\bm{x}$ and $\bm{y}$ in estimating the true state.
+The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is a matrix $H$ [1].
 $\bm{a}$ $=$ $\bm{x}+K\left(\bm{y}-H\bm{x}\right),$ (9a)
@@
 where $K$ is known as the Kalman gain matrix.
-The second term in Eq.
-(9a) describes the non-local nature of DA.
-The background state $\bm{x}$ is corrected through the spatially sparse observations $\bm{y}$ by the non-diagonal matrix $K$.
-Typically, this non-diagonal property originates from the non-zero off-diagonal elements of the background error covariance matrix $B$.
-In other words, the non-local correction is realized through the spatial correlation patterns of background values.
-The necessity of spatial correlations leads to a difficulty for DA by neural networks (Section III.3).
-III.3 Implementation of DA by CVAE
-DA can be performed by solving the minimization problem of $l_{\text{Gauss}}$.
-However, this minimization at every assimilation time step would be computationally intensive.
-We propose to efficiently conduct this minimization using a trained CVAE.
-Generally, the loss function for VAEs, including CVAEs, is an ELBO.Kingma and Welling (2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al.
-(2014); Sohn, Lee, and Yan (2015); Kingma and Welling (2019) This can be understood from the ELBO formula, Eq.
+These equations describe the non-locality of DA: an observation at one grid point is incorporated into the surrounding inferred values.
+This non-locality typically originates from the non-zero off-diagonal elements of the covariance matrix $B$.
+If $B$ is diagonal (i.e., all off-diagonal elements are zero), DA becomes local inference, and spatially sparse observations are incorporated into the inference only at their located grid points.
+3.3 Implementation of DA using CVAE
+DA can be performed by minimizing $l_{\text{Gauss}}$ in Eq.
 (7).
-To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from a given input of $\bm{x}$ and $\bm{y}$; then, the observation vector $\bm{\hat{y}}$ ($=\mathsf{H}(\bm{\hat{z}})$) is reconstructed from $\bm{a}$ and $V$ via Eq.
+However, this minimization at each assimilation time step would be computationally expensive.
+We propose an efficient approach to this minimization using a trained CVAE.
+The loss function for VAEs, including CVAEs, is generally an ELBO [7, 8, 9, 11, 12].
+This can be understood from the ELBO formula [Eq.
+(7)].
+To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from given $\bm{x}$ and $\bm{y}$; the observation vector $\bm{\hat{y}}$ [$=\mathsf{H}(\bm{\hat{z}})$] is then reconstructed from $\bm{a}$ and $V$ via Eq.
 (5).
 The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$.
-The structure that reconstructs the input is common to autoencoders.Dong et al.
-(2018) The second and third terms in Eq.
-(7) are considered as the conditions for the estimated true state $\bm{\hat{z}}$.
-For instance, $l_{\text{Gauss}}$ is decreased as $\bm{a}$ becomes close to $\bm{x}$.
-Thus, a CVAE is necessary to incorporate these conditions, and this CVAE is trained by minimizing $l_{\text{Gauss}}$.
+The architecture that reconstructs the input is a characteristic of autoencoders [10].
+The second and third terms in Eq.
+(7) are the conditions for the estimated true state $\bm{\hat{z}}$.
+To incorporate these conditions, we use a CVAE, which is trained by minimizing the loss function $l_{\text{Gauss}}$.
+This training is unsupervised because $l_{\text{Gauss}}$ does not require the ground truth for $\bm{z}$.
+This unsupervised learning is achieved through the use of the ELBO because $l_{\text{Gauss}}$ is derived from the ELBO in Eq.
+(2c).
 The CVAE consists of an encoder and a decoder (Fig.
@@
 The encoder estimates $\bm{a}$ and $V$ from the input $\bm{x}$ and $\bm{y}$, while the decoder reconstructs $\bm{\hat{y}}$.
-In other words, the encoder performs assimilation and the decoder acts as the nonlinear observation operator.
+In other words, the encoder performs assimilation, while the decoder acts as the observation operator.
 The encoder learns how to integrate $\bm{x}$ and $\bm{y}$, yielding the analysis $\bm{a}$, while the decoder learns how to convert $\bm{a}$ into the observation space.
-Once the CVAE is trained, the encoder performs DA without minimizing $l_{\text{Gauss}}$ as long as the statistics of the background $\bm{x}$ and the observations $\bm{y}$ do not change significantly.
+Once the CVAE is trained, the encoder performs DA without minimizing $l_{\text{Gauss}}$ as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
 Figure 2: Architecture of the CVAE for DA.
@@
 The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters).
-Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices: $R=rI_{n_{y}}$ and $B=bI_{n_{z}}$, where $I_{n}$ is an identity matrix of size $n$, and $r$ and $b$ are positive real parameters.
-However, off-diagonal elements of $B$ describe the spatial correlation patterns of background states and are essential for the non-local estimation in DA (Section III.2).
-There are several methods for implementing non-local estimation.
-For instance, latent variables can be transformed to have covariance or even follow non-Gaussian distributions by combining VAEs and normalizing flows,Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) where this transformation is learnable from data.
-An alternative approach is to divide observed values used as input to CVAEs and those targeted for reconstruction.
-CVAEs reconstruct observations from the other observations at different grid points, which enables CVAEs to learn non-locality.
-A similar technique is employed in MetNet-3, a surrogate neural network for weather predictions developed by Google.Andrychowicz et al.
-(2023) As a final method, we adopt super-resolution (SR).
-The non-locality is incorporated into CVAEs through SR even when $B$ and $R$ are diagonal.
-IV SRDA using the ELBO
-We incorporate SR into the DA theory proposed in Section III.
-A key characteristic of SR models is their non-locality:Dong et al.
-(2014); Gu and Dong (2021) SR models map LR patterns with a certain spatial extent to HR patterns.
-These SR mappings (i.e., SR operators) are integrated into DA, resulting in SRDA, which facilitates the non-local correction of background values by observations across distant grid points.
-IV.1 Integration of SR into DA
-To develop a theory of SRDA, we specify the resolutions of a background state $\bm{x}$, observations $\bm{y}$, and the true state $\bm{z}$: $\bm{x}$ is defined at LR grid points, while $\bm{y}$ and $\bm{z}$ are defined at HR grid points.
+Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices (i.e., having zero off-diagonal elements) [7, 8, 9, 11, 12].
+However, the off-diagonal elements of $B$ are essential for the non-local inference in DA (Section 3.2).
+We adopt super-resolution (SR) to realize non-local inference using CVAEs, leading to a new theory of SRDA.
+4 SRDA using the ELBO
+4.1 Integration of SR into DA
+To develop a theory of SRDA, we first specify the resolutions of a background state $\bm{x}$, observations $\bm{y}$, and the true state $\bm{z}$: $\bm{x}$ is defined at LR grid points, while $\bm{y}$ and $\bm{z}$ are defined at HR grid points.
 The vector forms are obtained by arranging three-dimensional grid point values into columns.
@@
 We modify the probability distributions in Eqs. (3a)–(3c) for SRDA:
-$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\exp\left\{-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\right\},$ (10a)
+$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}$ (10a)
+$\times$ $\exp\Bigg{\{}-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\Bigg{\}},$
 $p\left(\bm{z}\mid\bm{x}\right)$ $=$ $\frac{1}{(\sqrt{2\pi b})^{n_{z}}}\exp\left\{-\frac{1}{2b}\left\lVert\bm{z}-\mathsf{F}(\bm{x})\right\rVert^{2}\right\},$ (10b)
@@
 $V(\bm{x},\bm{y})=\text{diag}\left[v_{1}(\bm{x},\bm{y}),\ldots,v_{n_{z}}(\bm{x},\bm{y})\right],$ (12)
-where each variance $v_{i}(\bm{x},\bm{y})$ ($i=1,\ldots,n_{z}$) is a positive value determined by the vectors $\bm{x}$ and $\bm{y}$.
+where each variance $v_{i}(\bm{x},\bm{y})$ ($i=1,\ldots,n_{z}$) is positive and determined by $\bm{x}$ and $\bm{y}$.
 There are two essential changes from Eqs.
@@
 First, an SR operator $\mathsf{F}$ is introduced in Eq.
-(10b), where $\mathsf{F}$ maps an $n_{x}$-dimensional vector into an $n_{z}$-dimensional vector.
+(10b), mapping an $n_{x}$-dimensional vector into an $n_{z}$-dimensional vector.
 Since $\bm{x}$ and $\bm{z}$ are defined at LR and HR grid points, respectively (i.e., $n_{x}<n_{z}$), this mapping performs SR.
@@
 (3a) is replaced with a composition $\mathsf{A}_{\text{HR}}=\mathsf{F}\circ\mathsf{A}_{\text{LR}}$ in Eq.
-(10a): the LR analysis is first calculated as $\bm{a}_{\text{LR}}=\mathsf{A}_{\text{LR}}(\bm{x},\bm{y})$ and then it is super-resolved as $\bm{a}_{\text{HR}}=\mathsf{F}(\bm{a}_{\text{LR}})$.
-The influences of these two changes are further discussed in Sections IV.2 and IV.3.
-The ELBO is transformed by substituting Eqs. (10a)–(10c) into (2c), as in Section III.1:
-$l_{\text{SRDA}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y}))-\mathsf{F}(\bm{x})\right\rVert^{2}+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$ (13)
+(10a): the LR analysis is first calculated and then super-resolved.
+The ELBO for SRDA is obtained by substituting Eqs. (10a)–(10c) into Eq. (2c), as in Section 3.1:
+$l_{\text{SRDA}}$ $=$ $\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})-\mathsf{F}(\bm{x})\right\rVert^{2}$ (13)
+$\quad+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$
 where
-$\bm{a}_{\text{HR}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y}))=\mathsf{F}(\mathsf{A}_{\text{LR}}(\bm{x},\bm{y})),$ (14a)
-$\bm{a}_{\text{LR}}$ $=$ $\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}),$ (14b)
-$\bm{\hat{z}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\mathsf{F}(\bm{a}_{\text{LR}})+\sum_{i=1}^{n_{z}}\varepsilon_{i}\sqrt{v_{i}}.$ (14c)
+$\bm{\hat{z}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})+\bm{\varepsilon}^{\text{T}}V(\bm{x},\bm{y})^{1/2}=\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}V^{1/2},$ (14a)
+$\bm{a}_{\text{HR}}$ $=$ $\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})=\mathsf{F}(\bm{a}_{\text{LR}}),$ (14b)
+$\bm{a}_{\text{LR}}$ $=$ $\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}),$ (14c)
 Eqs.
-(13) and (14c) corresponds to Eqs.
-(7) and (5) in Section III.1, respectively.
-Like in Eq.
-(7), the Monte Carlo method (i.e., the reparameterization trickKingma and Welling (2014)) is applied in Eq.
+(13) and (14a) correspond to Eqs.
+(7) and (5), respectively, where $V$ in Eq.
+(14a) is defined by Eq.
+(12).
+As in Eq.
+(7), the Monte Carlo method (i.e., the reparameterization trick [8]) is applied in Eq.
 (13) with the sample size being 1.
 In the right-hand side of Eq.
-(14c), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definitions of $\bm{a}_{\text{HR}}$ and $\bm{a}_{\text{LR}}$ are used.
+(14a), the arguments of $\bm{x}$ and $\bm{y}$ are omitted, and the definition of $\bm{a}_{\text{HR}}$ [Eq.
+(14b)] is used.
+This $\bm{a}_{\text{HR}}$ is obtained by super-resolving $\bm{a}_{\text{LR}}$.
 The first term of $l_{\text{SRDA}}$ represents the difference between the observation vector $\bm{y}$ and the estimated true state $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
-The second term is the difference between the super-resolved analysis $\mathsf{F}(\bm{a}_{\text{LR}})$ and the super-resolved background $\mathsf{F}(\bm{x})$.
-The third term is a convex function of $v_{i}$, which takes the minimum when $v_{i}(\bm{x},\bm{y})=b$.
-These points indicate that the ELBO describes the balance between the observations and the super-resolved forecast in estimating the HR true state.
-IV.2 Case of linear SR
-We interpret $l_{\text{SRDA}}$ in Eq.
-(13) using a linear SR operator and show that SR operators act as background error covariance matrices.
-Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning.Park, Park, and Kang (2003) When generation processes of LR images from HR images are linear (e.g., subsampling and blurring), the inverse processes, namely the SR processes, are also linear.Park, Park, and Kang (2003) Here, an SR operator is denoted as a matrix $F$, which has the size of $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) and maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
-Since SR is generally non-local, the matrix $F$ has non-zero off-diagonal elements.
-To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section III.2.
+The second term is the difference between the super-resolved analysis $\bm{a}_{\text{HR}}$ [$=\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})$] and the super-resolved background $\mathsf{F}(\bm{x})$.
+The third term is a convex function of $v_{i}$ that takes the minimum when $v_{i}(\bm{x},\bm{y})=b$.
+The ELBO (i.e., $l_{\text{SRDA}}$) describes the balance between the observations and the super-resolved background state in estimating the HR true state.
+4.2 Case of linear SR
+The effects of non-locality of SR are examined separately for linear and nonlinear SR. For linear SR, we show that SR operators act as background error covariance matrices.
+Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning [47].
+When generation processes of LR images from HR images are linear (e.g., subsampling or blurring), the inverse processes, namely the SR processes, are also linear [47].
+An SR operator is denoted by a matrix $F$ of size $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) that maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
+To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section 3.2.
 The estimated true state $\bm{\hat{z}}$ in Eq.
-(14c) is then equal to $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$.
+(14a) is then approximated as $\bm{\hat{z}}=F\bm{a}_{\text{LR}}$.
 Further, the third term in Eq.
 (13) is ignored because it becomes constant.
-Thus, $l_{\text{SRDA}}$ is reduced to
-$l_{\text{linear SRDA}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(F\bm{a}_{\text{LR}})\right\rVert^{2}+\frac{1}{2b}\left(\bm{a}_{\text{LR}}-\bm{x}\right)^{\text{T}}\left(F^{\text{T}}F\right)\left(\bm{a}_{\text{LR}}-\bm{x}\right).$ (15)
+Consequently, $l_{\text{SRDA}}$ is reduced to
+$l_{\text{SRDA-L}}=\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(F\bm{a}_{\text{LR}})\right\rVert^{2}+\frac{1}{2b}\left(\bm{a}_{\text{LR}}-\bm{x}\right)^{\text{T}}\left(F^{\text{T}}F\right)\left(\bm{a}_{\text{LR}}-\bm{x}\right),$ (15)
+where SRDA-L stands for SRDA with linear SR.
 The matrix $F^{\text{T}}F$ is an $n_{x}\times n_{x}$ positive semi-definite matrix.
 Compared to the ELBO for 3D-Var in Eq.
-(8), $F^{\text{T}}F$ can be interpreted as the covariance matrix $B$ and describes the spatial correlations among the LR components.
-When the observation operator is linear ($\mathsf{H}=H$) as in Section III.2, the minimizer for $l_{\text{linear SRDA}}$ is always obtained due to the convexity of $l_{\text{linear SRDA}}$.
-We first consider the case where $F^{\text{T}}F$ is positive definite (i.e., invertible).
-The minimizer is then derived, as in Eq.
-(9):
-$\bm{a}_{\text{LR}}$ $=$ $\bm{x}+\widetilde{K}\left(\bm{y}-\widetilde{H}\bm{x}\right),$ (16a)
-$\widetilde{K}$ $=$ $\widetilde{B}\widetilde{H}^{\text{T}}\left(rI_{n_{y}}+\widetilde{H}\widetilde{B}\widetilde{H}^{\text{T}}\right)^{-1},$ (16b)
-where $\widetilde{B}=b\left(F^{\text{T}}F\right)^{-1}$ and $\widetilde{H}=HF$.
-The observation operator for the LR background is given by $\widetilde{H}=HF$ through the SR operator $F$.
-Generally, the Kalman gain matrix $\widetilde{K}$ has non-zero off-diagonal elements due to $F$ being non-diagonal.
-Thus, the LR background $\bm{x}$ is non-locally corrected by the observations $\bm{y}$, giving the LR analysis $\bm{a}_{\text{LR}}$ in Eq.
-(16a).
-Even when $F^{\text{T}}F$ is positive semi-definite (i.e., at least one eigenvalue is zero), minimizers for $l_{\text{linear SRDA}}$ exist due to its convexity.
-In this case, a solution similar to Eq.
-(16a) is derived using pseudo-inverse matrices (not shown in detail).
-IV.3 Case of non-linear SR
-In practice, SR models are neural networks, so SR operators are non-linear.
-Since the nonlinear case includes the linear case, the non-local inference is achieved in both cases.
-Here, we briefly examine non-linear SR operators from a different perspective: the transformation of probability distributions.
-We consider a complete non-linear SR operator $\mathsf{F}$, that is, a bijective SR operator.
-Rigorously speaking, SR is an inverse problem and SR operations are not bijective; however, a virtual bijective map is sometimes considered in theoretical studies of SR.
-In this case, any HR true state $\bm{z}$ has the corresponding LR state: $\bm{z}_{\text{LR}}=\mathsf{F}^{-1}(\bm{z})$ and $\bm{z}=\mathsf{F}(\bm{z}_{\text{LR}})$.
-We can regard the posterior and prior distributions in Eqs.
-(10a) and (10b) as non-Gaussian distributions for $\bm{z}_{\text{LR}}$:
-$q\left(\bm{z}_{\text{LR}}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{|D\mathsf{F}|}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\exp\left\{-\frac{1}{2}\left\lVert\mathsf{F}(\bm{z}_{\text{LR}})-\mathsf{F}(\mathsf{A}_{\text{LR}}(\bm{x},\bm{y}))\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\right\},$ (17a)
-$p\left(\bm{z}_{\text{LR}}\mid\bm{x}\right)$ $=$ $\frac{|D\mathsf{F}|}{(\sqrt{2\pi b})^{n_{z}}}\exp\left\{-\frac{1}{2b}\left\lVert\mathsf{F}(\bm{z}_{\text{LR}})-\mathsf{F}(\bm{x})\right\rVert^{2}\right\},$ (17b)
-where $|D\mathsf{F}|$ denotes the Jacobian determinant of $\mathsf{F}$.
-This result suggests that SR operators can be interpreted as normalizing flows.Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) Normalizing flow is a deep generative model that converts Gaussian variables into non-Gaussian variables, generating new data that obey complicated probability distributions.
-In normalizing flows, invertible non-linear transformations are implemented as neural-network layers and are stacked to form the whole network to achieve complex transformations.
-Importantly, normalizing flows can be used to transform latent variables in VAEs.Rezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) This combined architecture is similar to the CVAE proposed in the present study (Section IV.5), where the SR model transforms $\bm{a}_{\text{LR}}$ into $\bm{a}_{\text{HR}}$ inside the encoder.
-This similarity implies the potential for using normalizing flows to further develop SRDA.
-IV.4 Loss function for SR models
-So far, we have employed the SR operator $\mathsf{F}$, namely the trained SR model.
-This section presents a loss function for training this model.
-We do not know the true state of the system and only know the LR background $\bm{x}$ and the HR observations $\bm{y}$.
-The former $\bm{x}$ is spatially complete, while the latter $\bm{y}$ is spatially incomplete.
+(8), $F^{\text{T}}F$ acts as the covariance matrix $B$.
+Due to the non-locality of SR, $F^{\text{T}}F$ is non-diagonal and describes the spatial correlations among the LR components (Fig.
+3).
+Figure 3: Schematic of spatial correlation through a linear SR operator $F$.
+This $F$ super-resolves an LR image $\bm{w}$ to the HR image $\bm{\hat{w}}$ ($=F\bm{w}$).
+Here, $\bm{w}$ corresponds to $\bm{a}_{\text{LR}}-\bm{x}$ in Eq.
+(15).
+The two blue pixels in $\bm{w}$ contribute to the red pixel in $\bm{\hat{w}}$.
+By taking an inner product $\bm{\hat{w}}^{\text{T}}\bm{\cdot}\bm{\hat{w}}$, the two blue pixels are correlated with each other.
+4.3 Case of nonlinear SR
+In practice, SR operators are implemented as neural networks and are thus nonlinear.
+Since nonlinear SR encompasses linear SR, non-local inference is also achieved with nonlinear SR.
+In this section, non-locality is examined from a perspective of backpropagation through an SR neural network.
+We consider the nonlinear version of Eq. (15):
+$l_{\text{SRDA-NL}}=\frac{1}{2r}\left\lVert H\mathsf{F}(\bm{a})-\bm{y}\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{F}(\bm{a})-\mathsf{F}(\bm{x})\right\rVert^{2},$ (16)
+where SRDA-NL stands for SRDA with nonlinear SR.
+For simplicity, we set the observation operator as a matrix ($\mathsf{H}=H$) and denote $\bm{a}_{\text{LR}}$ as $\bm{a}$.
+The minimizer $\bm{a}$ can be estimated by gradient descent in an iterative manner:
+$a^{(n+1)}_{i}$ $=$ $a^{(n)}_{i}-\eta^{(n)}\frac{\partial}{\partial a^{(n)}_{i}}l_{\text{SRDA-NL}},$ (17a)
+$\frac{\partial}{\partial a^{(n)}_{i}}l_{\text{SRDA-NL}}$ $=$ $\frac{1}{r}\sum_{j,k,l}\left[H_{jk}\mathsf{F}_{k}(\bm{a}^{(n)})-y_{j}\right]H_{jl}\frac{\partial\mathsf{F}_{l}(\bm{a}^{(n)})}{\partial a^{(n)}_{i}}$ (17b)
+$+$ $\frac{1}{b}\sum_{j}\left[\mathsf{F}_{j}(\bm{a}^{(n)})-\mathsf{F}_{j}(\bm{x})\right]\frac{\partial\mathsf{F}_{j}(\bm{a}^{(n)})}{\partial a^{(n)}_{i}},$
+where a superscript, such as $(n)$, denotes the number of iterations, and $\eta^{(n)}$ is a small positive value for the $n$-th iteration.
+Eqs.
+(17a) and (17b) are similar to backpropagation formulas using mean squared errors as a loss function [48].
+The differences from the observations (i.e., errors), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagate backward through the SR neural network $\mathsf{F}$.
+This backpropagation updates the current solution $\bm{a}^{(n)}$ (Fig.
+4).
+Note that all parameters of $\mathsf{F}$ remain fixed because $\mathsf{F}$ is already trained.
+Since trained SR neural networks generally exhibit non-locality [30, 31, 32, 33], the backpropagation updates the analysis vector $\bm{a}$ from observations over distant grid points (Fig.
+4).
+We demonstrate that a trained SR model actually exhibits this non-locality in Section 6.2.
+Equation (17b) indicates that the differences from the super-resolved background state, $\mathsf{F}(\bm{a}^{(n)})-\mathsf{F}(\bm{x})$, also induce updates through backpropagation.
+The ELBO (16) determines the balance between the updates from observation and background values.
+Figure 4: Schematic of the non-local update by backpropagation through an SR neural network (i.e., a non-linear SR model).
+A large error (red) is propagated backward through yellow nodes in the hidden layers and updates the LR input values.
+Minimizing Eq.
+(16) through backpropagation may be computationally expensive, although automatic differentiation in deep learning frameworks facilitates the calculation of $\partial\mathsf{F}/\partial\bm{a}^{(n)}$ in Eq.
+(17b).
+As in Section 3.3, we propose to solve the minimization problem using CVAEs.
+4.4 Loss function for SR models
+So far, we have used the trained SR model $\mathsf{F}$.
+This section introduces a loss function for training $\mathsf{F}$.
+We do not know the true state of the system and only know the LR background $\bm{x}$ and the HR observations $\bm{y}$, where $\bm{x}$ is spatially complete and $\bm{y}$ is spatially incomplete.
 The SR model is trained in this setup.
-We propose the following loss function to train SR models:
+Assuming that the observation operator is a projection matrix, we propose the following loss function to train SR models:
 $l_{\text{SR}}=\sum_{i=1}^{n_{y}}\left|y_{i}-[P\mathsf{F}(\bm{x})]_{i}\right|,$ (18)
-where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space.
-The right-hand side clarifies that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies.Ha et al.
-(2019); Anwar, Khan, and Barnes (2020); Fukami, Fukagata, and Taira (2023) Since $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$, backpropagation can be performed; that is, the SR model $\mathsf{F}$ can be trained in a supervised manner, where the input is $\bm{x}$ and the target is $\bm{y}$.
-When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$, as shown in Hastie et al.Hastie, Tibshirani, and Friedman (2009) This fact suggests that the SR model can make unbiased inferences if the median of the observation error distribution is zero.
-IV.5 Implementation of SRDA by CVAE
-The efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, similar to Section III.3. Thus, SRDA can be implemented using CVAEs.
+where ||$\cdot$|| ⋅ | denotes the absolute value, $i$ specifies vector components, and $P$ is the projection matrix into the observation space ($\mathsf{H}=P$).
+Typically, $P$ consists of interpolation and/or subsampling operations.
+The right-hand side shows that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies [27, 28, 29, 41].
+Backpropagation is feasible because $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$.
+The SR model $\mathsf{F}$ is trained using $\bm{x}$ as input and $\bm{y}$ as target.
+When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$ [49].
+This result suggests that the SR model becomes unbiased if the median of the observation error distribution is zero.
+4.5 Implementation of SRDA using CVAE
+An efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, as in Section 3.3, as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
 The CVAE for SRDA consists of an encoder and a decoder (Fig.
-3).
+5).
 The difference from the CVAE in Fig.
 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously.
-Here, we employ the SR model described in Section IV.4, which is trained with the loss function in Eq.
+We use the SR model in Section 4.4, which is trained with the loss function in Eq.
 (18).
-This encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$ and estimates the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$.
-The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ by calculating the variances.
+The encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$, estimating the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$.
+This computation of $\bm{a}_{\text{LR}}$ [$=A_{\text{LR}}(\bm{x},\bm{y})$] emulates the result of iterative backpropagation through $\mathsf{F}$ (i.e., the result of minimizing $l_{\text{SRDA-NL}}$ in Section 4.3).
+The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] by calculating the variance $V$ in Eq.
+(12).
 The decoder acts as the observation operator, as in Fig.
@@
 The encoder estimates the HR analysis vector $\bm{a}_{\text{HR}}$ using the SR model.
-The encoder also estimates the covariance $V$ for $\bm{a}_{\text{HR}}$.
+The encoder also estimates the variance $V$ for $\bm{a}_{\text{HR}}$.
 The variables $\bm{a}_{\text{HR}}$ and $V$ give a sample of the true state $\bm{\hat{z}}$ in Eq.
-(14c).
+(14a).
 The decoder acts as the observation operator and reconstructs the observation vector $\bm{\hat{y}}=\mathsf{H}(\bm{\hat{z}})$.
-Strictly speaking, a different SR model should be used within the encoder because the SR model in Section IV.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$.
-This difference in input may introduce bias to $\bm{a}_{\text{LR}}$.
-In other words, the trained SR model assumes $\bm{x}$ as input and may lead to a bias of $\bm{a}_{\text{LR}}$ similar to that of $\bm{x}$.
-On the other hand, the output $\bm{a}_{\text{HR}}$ ($=\mathsf{F}(\bm{a}_{\text{LR}})$) is expected to be less biased because the SR-model output is unbiased if the observation error distribution has a zero median (Section IV.4).
-Thus, $\bm{a}_{\text{HR}}$ is used as the initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section IV.6).
-One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model during the training of the CVAE.
-For simplicity, however, the present study does not use such an advanced learning method, which remains a future research topic.
-IV.6 Sequential SRDA by the CVAE
+Rigorously speaking, a different SR model should be used within the encoder because the SR model in Section 4.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$.
+This difference in input may introduce bias in $\bm{a}_{\text{LR}}$.
+In contrast, the output $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] is expected to be less biased because the SR model is unbiased if the observation error distribution has a zero median (Section 4.4).
+Thus, $\bm{a}_{\text{HR}}$ is used as initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section 4.6).
+One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model $\mathsf{F}$ during the training of the CVAE.
+For simplicity, this study does not use such an advanced learning method, which remains a future research topic; instead, the parameters of $\mathsf{F}$ are fixed during the CVAE training.
+4.6 Sequential SRDA using the CVAE
 We introduce a sequential SRDA method applicable to CFD simulations.
 The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model.
-The forecast and assimilation are executed alternately, as in other statistical DA methods,Asch, Bocquet, and Nodet (2016); Carrassi et al.
-(2018) such as 3D-Var.
-This assimilation is performed using a trained CVAE’s encoder (Fig.
-3).
+The forecast and assimilation are executed alternately, as in other statistical DA methods [1, 2].
+This assimilation is performed using a trained CVAE’s encoder.
 Figure 6 shows a schematic of the sequential SRDA.
-Here, we assume a constant assimilation interval $\Delta T$.
-At every assimilation time step, the encoder assimilates the HR observations $\bm{y}$ into the LR forecast state $\bm{x}$, thereby estimating the HR analysis $\bm{a}_{\text{HR}}$.
-This output is resized to the LR using an algebraic method, such as linear interpolation.
+We assume a constant assimilation interval $\Delta T$.
+At each assimilation time step, the encoder assimilates the HR observations $\bm{y}$ into the LR forecast state $\bm{x}$, thereby estimating the HR analysis $\bm{a}_{\text{HR}}$.
+This $\bm{a}_{\text{HR}}$ is resized to the LR using an algebraic method, such as linear interpolation.
 In the next assimilation cycle, the resized analysis is used as the initial condition for the LR fluid model, and forecasts are computed by numerical integration over $\Delta T$.
@@
 A constant assimilation interval $\Delta T$ is assumed, where the output time step is a quarter of $\Delta T$ as an example.
-The decoder is not used in the sequential SRDA; rather, it is used only in the training phase to learn the observation operator (Section IV.5).
+The decoder is not used in the sequential SRDA; rather, it is used only in the training phase to learn the observation operator.
 In contrast, the trained SR model can be used in the test phase (i.e., the operational phase) to super-resolve LR forecast states.
-This SR operation is independent of assimilation, allowing for the generation of HR forecast states at any time step (Section VI.3).
-V Methods
-We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs).Fletcher (2023a) The ground truth (or the nature run) is generated by an HR fluid model, from which pseudo (or synthetic) observations are computed.
+This SR operation is independent of assimilation, enabling the inference of HR forecasts at any time step.
+5 Methods
+We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs) [50], which are commonly used in testing DA methods for the atmosphere and ocean.
+The ground truth (or the nature run) is generated by an HR fluid model, from which synthetic observations are computed.
 The SRDA employs an LR version of the fluid model, and the resultant forecast and analysis are compared with the ground truth.
-A simple jet stream system was adopted for these experiments, as a proof of concept for the SRDA.
-V.1 Fluid simulation
-An idealized barotropic ocean jet, proposed by David et al.,David, Marshall, and Zanna (2017) was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi.Yasuda and Onishi (2023) The experimental setup is briefly explained here.
+A simple oceanic jet system was adopted for these experiments as a proof of concept for the SRDA.
+This jet exhibits the development of small-scale flow patterns and the sensitivity to initial conditions.
+SR can infer such small-scale patterns, while DA can suppress error growth due to the initial condition sensitivity.
+This jet system is suitable for assessing both aspects of SR and DA.
+5.1 Fluid simulation
+An idealized barotropic ocean jet, proposed by David et al.
+[51], was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi [35].
+The experimental setup is briefly explained here.
 See the above references for further details.
-The governing equations in the periodic channel $(x,y)$ are as followsDavid, Marshall, and Zanna (2017)
+The governing equations in the periodic channel $(x,y)$ are as follows [51]:
 $\frac{\partial\omega}{\partial t}+\bm{u\cdot\nabla}\omega+\beta\frac{\partial\psi}{\partial x}$ $=$ $-\kappa\omega-\nu\Delta^{2}\omega-\frac{d\tau(y)}{dy},$ (19a)
@@
 Equation (19a) describes the evolution of vorticity $\omega$ over time $t$.
-The left-hand side represents the advection of $\omega$ and the beta effect,Vallis (2017a) while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$.
+The left-hand side represents the advection of $\omega$ and the beta effect [52], while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$.
 Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively.
@@
 where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero.
-This setup falls within the parameter regime of mixing barriers with strong eddies,David, Marshall, and Zanna (2017) where coherent vortices persist in a statistically steady state (Section VI.1).
+This setup falls within the parameter regime of mixing barriers with strong eddies [51], where coherent vortices persist in a statistically steady state (Section 6.1).
 The initial condition was a zonal jet superimposed with random perturbations.
-The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion,Vallis (2017b) which is a necessary condition for barotropic instability.
-Perturbations were added to the vorticity field for each wavenumber component, with phase and amplitude drawn from Gaussian distributions.
-The simulation data for deep learning were generated by varying these random perturbations.
+The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion [53], which is a necessary condition for barotropic instability.
+Perturbations were added to the vorticity field for each wavenumber component, with phases and amplitudes drawn from Gaussian distributions.
+The data for deep learning were generated by varying these random perturbations.
 Numerical integration was conducted using the modified Euler and pseudo-spectral methods at two spatial resolutions: LR and HR.
-Table 1 provides detailed configurations for the LR and HR.
+Table 1 provides detailed configurations for the LR and HR simulations.
 All simulations ran from $t=0$ to $20$ with output time intervals of $0.25$.
@@
 Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR.
-The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule).Canuto et al.
-(1988)
+The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule) [54].
 Name Grid size ($x\times y$) Cutoff wavenumber Integration time step
@@
 HR (high resolution) $128\times 64$ 42 $\frac{1}{4}\times 5\times 10^{-4}$
-V.2 Synthetic observations
+5.2 Synthetic observations
 Synthetic observations were generated from the HR simulations.
-We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere.Lin and Yang (2020); Fletcher (2023b) Observed grid points were randomly subsampled at constant spatial intervals, namely $8\times 8$ grids, which means that 1.56% ($=1/8^{2}$) of grid points are observed.
-To demonstrate that the SRDA can address changes in observed points, the subsampled points were varied in time by random shifts.
-For the unobserved points, Not a Numbers (NaNs) were assigned and replaced with missing values in the subsequent preprocessing (Section V.3).
+We consider an idealized point-wise observation network that imitates moored buoys in the ocean or balloon-borne radiosondes in the atmosphere.
+Observation points were subsampled at constant spatial intervals of $8\times 8$ grid points, which means that 1.56% ($=1/8^{2}$) of grid points were observed.
+To demonstrate that the SRDA can address changes in observation points, the subsampled points were varied in time by random shifts.
+For the unobserved points, Not a Numbers (NaNs) were assigned and later replaced with missing values in the preprocessing (Section 5.3).
 To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity.
@@
 This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
-V.3 Neural networks
-We outline the network architectures of an SR model and a CVAE for the sequential SRDA (Section IV.6).
-Hyperparameters are provided in Appendix and the full implementation is available at our GitHub repository (see Data Availability Statement).
-PyTorch 1.11.0Paszke et al.
-(2019) was used for the implementation.
-The SR model architecture is shown by the top panel in Fig.
-5.
-This model resembles the neural network for SR known as VDSR, proposed by Kim et al.Kim, Lee, and Lee (2016) The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR.
+5.3 Neural networks
+We outline the network architectures of an SR model and a CVAE for the sequential SRDA.
+Hyperparameters are provided in the Appendix, and the full implementation is available at our Zenodo repository (see Data Availability Statement).
+PyTorch 1.11.0 [55] was used for the implementation.
+The SR model architecture is shown in the top panel of Fig.
+7.
+This model resembles the neural network for SR known as VDSR, proposed by Kim et al.
+[56].
+The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR.
 First, this input is nonlinearly transformed through a stack of convolution and ReLU layers.
-Upsampling is then executed by two blocks that include pixel shuffle layers.Shi et al.
-(2016) Each block doubles the size of each dimension, resulting in a fourfold increase in resolution.
-Additionally, a skip connectionHe et al.
-(2016) is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output.
-This skip connection allows the SR model to capture small-scale patterns, that is, discrepancies between the LR input and the HR output.
+Upsampling is then executed by two blocks that include pixel shuffle layers [57].
+Each block doubles the size of each dimension, resulting in a fourfold increase in resolution.
+Additionally, a skip connection [58] is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output.
+This skip connection allows the SR model to capture the discrepancies between the LR input and the HR output.
 Figure 7: Network architectures of the SR model and the CVAE.
-The latter consists of the encoder and the decoder.
-The label “Res Block” refers to a residual block, and “Res Block x6” indicates the stacking of six Res Blocks.
-As in the typical implementation of VAEs,Kingma and Welling (2019) the variance output is regarded as $\ln V$ because neural networks can generally output both positive and negative values.
-Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp[1/2\ln V]$ provides a realization of the true state $\bm{\hat{z}}$ in Eq.
-(14c).
+The latter consists of an encoder and a decoder.
+The label “Res Block” refers to a residual block, and “Res Block x6” refers to the stacking of six Res Blocks.
+As in the typical implementation of VAEs [7], the output of variance is regarded as $\ln V$ because neural networks can generally output both positive and negative values.
+Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp(1/2\ln V)$ provides a realization of the true state $\bm{\hat{z}}$ in Eq.
+(14a).
+All inputs and outputs consist only of vorticity (i.e., 1 channel).
 The CVAE consists of an encoder and a decoder (Fig.
-5).
+7).
 The encoder performs SRDA and integrates the inputs, an LR forecast and an HR observation, into the HR analysis.
-The HR observation is first downsampled through convolution to align the resolution to the LR (totally, downsampled by a factor of 4).
+The HR observation is first downsampled through convolution to match its resolution with the LR (downsampled by a factor of 4).
 The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$.
-Then, $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$.
-The variance for $\bm{a}_{\text{HR}}$ is denoted as $V$ in Eq.
-(12) and is also estimated by the encoder.
-Although $V$ and $\bm{a}_{\text{HR}}$ can be treated independently in the encoder (Section IV.5), the present study computes $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of neural network parameters.
-Moreover, we employed the feature extractor in the trained SR model to compute $V$ (Feature Extractor-1 in Fig.
-5).
+This $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$.
+The variance for $\bm{a}_{\text{HR}}$ [$V$ in Eq.
+(12)] is also estimated by the encoder.
+Although $\bm{a}_{\text{HR}}$ and $V$ can be treated independently in the encoder (Section 4.5), we compute $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of parameters, using the feature extractor in the trained SR model (i.e., Feature Extractor-1 in Fig.
+7).
 The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$.
-This operation is achieved through a stack of convolutional and Leaky ReLUMaas et al.
-(2013) layers.
-We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), although this replacement slightly decreases the inference accuracy.
-This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section V.2); that is, the observation operator is linear.
-Although the proposed SRDA can theoretically address nonlinear observation operators (Section IV), such a nonlinear case is not examined in this proof of concept, which remains a topic for future research.
+This operation comprises convolution and Leaky ReLU [59].
+We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), though this replacement slightly decreases the inference accuracy.
+This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section 5.2).
 In preprocessing, the vorticity $\omega$ at each grid point is transformed by
@@
 After applying Eq.
-(21), zero is assigned to the HR grid points without observations, and zero is regarded as a missing value.
-This clear interpretation of zero facilitates feature extraction from input data.Yasuda, Onishi, and Matsuda (2023)
-V.4 Training and testing methods for the neural networks
-V.4.1 Training method
-We adopted an offline strategy in which training data are pre-generated [Fig.
-6(a)].
-The training data consist of pairs of HR observations and LR forecasts, as discussed in Section IV.5.
+(21), zero is regarded as a missing value and is assigned to the HR grid points without observations.
+This clear interpretation of zero facilitates feature extraction from the input data [60].
+5.4 Training and testing methods for the neural networks
+5.4.1 Training method
+We adopted an offline strategy in which training data were pre-generated (Fig.
+8a).
+The training data consist of pairs of HR observations and LR forecasts, as discussed in Sections 4.4 and 4.5.
 The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain.
-After time integration over $\Delta T=1$, the resultant LR forecasts were used in training.
-This $\Delta T$ equals the assimilation interval, as explained below, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet.
-The corresponding observations were obtained by subsampling and adding Gaussian noise (Section V.2).
-Figure 6(a) denotes pairs of the LR forecast and the HR observations as stars.
-By varying the random initial perturbations (Section V.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
+This filtering completely removes all high-wavenumber components using discrete Fourier transform.
+After time integration over $\Delta T$ ($=1$), the resultant LR forecasts were used in training.
+This $\Delta T$ equals the assimilation interval, which is a reasonable timescale because $\Delta T=1$ is approximately half of the advection time of the zonal jet.
+The corresponding observations were obtained by subsampling.
+Figure 8a denotes pairs of the LR forecasts and HR observations as stars.
+By varying the random initial perturbations (Section 5.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
 Figure 8: Schematics of time series for (a) training and (b) testing phases.
 The assimilation interval is set to $\Delta T=1$, while the output interval (i.e., the forecast interval) is $\Delta T/4=0.25$.
-In principle, no HR simulation is required because the training data include LR forecasts rather than HR results (Sections IV.4 and IV.5).
+In principle, HR simulations are not required because the training data include LR forecasts rather than HR results.
 These LR forecasts are obtained by numerical integration in which the analysis states are used as initial conditions.
 In offline training, however, such an analysis is unavailable because the CVAE is not trained yet.
-In realistic applications, these initial conditions can be given from existing assimilated data, such as reanalysis,Hersbach et al.
-(2020); KOSAKA et al.
-(2024) or online training will be effective,Rasp (2020) where neural-network training and data generation are executed alternately.
+Thus, the filtered HR results were used as the initial conditions.
+In realistic applications, these initial conditions can be obtained from existing assimilated data [61, 62], or online training would be effective [63, 64], where training and data generation are performed alternately or in parallel.
 The SR model was trained, followed by training the CVAE.
-During the CVAE training, all parameters in the SR model were fixed, as discussed in Section IV.5.
+For both trainings, the Adam optimizer [65] was used with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128.
+Each training was terminated using early stopping with a patience parameter of 50 epochs.
+During the CVAE training, we fixed all parameters in the SR model.
 The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq.
@@
 (13), respectively.
-The CVAE and the SR model were trained using the Adam optimizerKingma and Ba (2015) with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128.
-Each training was terminated using early stopping with a patience parameter of 50 epochs.
 The parameters in $l_{\text{SRDA}}$ were set to $r=1.2\times 10^{-5}$ and $b=1.0\times 10^{-3}$.
-The value of $r$ was given by the variance of the measurement errors (Section V.2), while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO.
+The value of $r$ was determined by the variance of the measurement errors, while $b$ was determined through grid search to balance the reconstruction error and the KL divergence in the ELBO.
 Without this balancing, the performance of the CVAE would deteriorate.
-This phenomenon is often observed in VAEs,Higgins et al.
-(2017); Alemi et al.
-(2018) and the loss balancing can be optimized during training.Asperti and Trentin (2020) For simplicity, however, such an advanced technique was not employed in the present study.
-V.4.2 Testing method
-The sequential SRDA was tested using the method described in Section IV.6.
-Figure 6(b) shows a schematic of the test phase.
+This phenomenon frequently occurs in VAEs [66, 67], and the loss balancing can be optimized during training [68].
+For simplicity, however, this advanced technique was not employed in our study.
+5.4.2 Testing method
+The sequential SRDA was tested using the method described in Section 4.6.
+Figure 8b shows a schematic of the test phase.
 The only difference between Figs.
-4 and 6(b) is that observations are generated by subsampling the HR simulation results.
-Due to this difference, the HR results are regarded as the ground truth, and the SRDA can be evaluated by comparing its inference to the ground truth.
+6 and 8b is that observations are generated by subsampling the HR simulation results.
+Due to this difference, the HR results are regarded as the ground truth, and the SRDA is evaluated by comparing its inferences to the ground truth.
 We newly conducted 500 HR fluid simulations.
 For each case, the LR fluid model was initialized with the low-pass filtered HR vorticity at $t=0$; the SRDA was subsequently conducted until $t=20$ with the assimilation interval $\Delta T=1$.
-Without SRDA, the vorticity evolution would differ from the ground truth, not only due to the difference in the initial condition (i.e., all high-wavenumber components are removed), but also due to the difference in the fluid model resolution (i.e., LR versus HR).
+Without SRDA, the vorticity evolution would significantly differ from the ground truth, not only due to differences in the initial condition (i.e., all high-wavenumber components are completely removed) but also due to the difference in the fluid model resolution (i.e., LR versus HR).
 Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth.
+These two types of metrics are frequently used in SR studies [29] because they complement each other.
+The pixel-wise accuracy quantifies localized differences but may overestimate discrepancies when spatial patterns are correctly inferred with a slight shift.
+The pattern consistency mitigates this effect by evaluating structural similarity between inferred and ground-truth patterns, although it may overlook significant pixel-wise deviations.
 The mean absolute error (MAE) ratio quantifies pixel-wise errors.
 MAE $\text{MAE ratio}=\frac{\sum_{i}|\omega_{i}-\hat{\omega}_{i}|}{\sum_{i}|\omega_{i}|},$ (22)
-where the summation is taken over all grid points, $\omega_{i}$ denotes the ground-truth vorticity at the $i$-th grid point, and $\hat{\omega}_{i}$ is the corresponding inference.
-The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns:Wang et al.
-(2004)
+where $\omega_{i}$ denotes the ground-truth vorticity at the $i$-th grid point, $\hat{\omega}_{i}$ is the corresponding inference, and the summation is taken over all grid points.
+The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns [69]:
 MSSIM loss $=$ $1-{\text{MSSIM}},$ (23)
@@
 where ${\text{C}_{1}}=0.01$, ${\text{C}_{2}}=0.03$, $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the ground truth, respectively, $\hat{\mu}_{i}$ and $\hat{\sigma}_{i}^{2}$ are the corresponding quantities for the inference, and $\gamma_{i}$ is the covariance between the ground truth and the inference.
-The variables in the summation are calculated locally in space using a Gaussian filter.
-The MSSIM loss takes a value greater than or equal to 0; a smaller value indicates spatial patterns more similar to the ground truth.
-A detailed discussion of MSSIM can be found in Wang et al.Wang et al.
-(2004)
+These variables in the summation are calculated locally in space using a Gaussian filter.
+The MSSIM loss takes a value greater than or equal to zero; a smaller value indicates spatial patterns more similar to the ground truth.
+A detailed discussion of MSSIM can be found in Wang et al.
+[69].
 The MAE ratio and MSSIM loss were calculated at each time step in each test simulation.
@@
 Both MAE ratio and MSSIM loss are collectively called the test errors.
-V.5 Baseline model using Ensemble Kalman Filter (EnKF)
-As a comparison to the SRDA, we employed an ensemble Kalman filter (EnKF)Evensen (1994) for the LR fluid model, using the perturbed observation method.Burgers, van Leeuwen, and Evensen (1998) The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations.
-The uncertainty with the EnKF is compared with that with the SRDA.
-Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, where the assimilation interval $\Delta T=1$ is the same as in the SRDA.
-Barthélémy et al.Barthélémy et al.
-(2022) argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points.
+5.5 Baseline model using Ensemble Kalman Filter (EnKF)
+As a baseline for comparison with the SRDA, we employed an ensemble Kalman filter (EnKF) [70] for the LR fluid model, using the perturbed observation method [71].
+The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations.
+The uncertainties estimated by the EnKF and SRDA are compared in Section 6.4.
+Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, with the assimilation interval $\Delta T$ ($=1$) being the same as in the SRDA.
+Barthélémy et al.
+[34] argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points.
 Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
-The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn.Gaspari and Cohn (1999) These error covariances were inflated by adding Gaussian noise to the analysis, prior to the forecast process.Whitaker et al.
-(2008) This Gaussian noise incorporates spatial correlation by estimating its covariance from the training data at each time step.
+The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn [72].
+These error covariances were inflated by adding Gaussian noise to the analysis before the forecast process [73].
+This Gaussian noise incorporates spatial correlation by estimating the covariance from the training data at each time step.
 The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution.
-We confirmed that the results are not sensitive to the statistics of the Gaussian noise for inflation.
-Indeed, similar results were obtained when the Gaussian noise had a constant correlation length.
 Hyperparameters for the EnKF were adjusted using the training data to minimize the MAE.
 Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations.
-The full implementation of the EnKF is also available at our GitHub repository (see Data Availability Statement).
-VI Results and discussion
-VI.1 Time evolution of vorticity in the jet system
-Before discussing the SRDA results, we describe the time evolution of HR vorticity in the jet system and confirm the sensitivity to initial conditions (i.e., a positive Lyapunov exponent).
+The full implementation of the EnKF is also available at our Zenodo repository (see Data Availability Statement).
+6 Results and discussion
+6.1 Time evolution of vorticity in the jet system
+We first describe the time evolution of HR vorticity in the jet system and confirm the sensitivity to initial conditions (i.e., a positive Lyapunov exponent).
 Figure 9 shows a typical evolution of HR vorticity.
@@
 Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$).
-This final state is statistically steady and can be understood by Rossby wave dynamics.David, Marshall, and Zanna (2017)
+This final state is statistically steady and can be understood by Rossby wave dynamics [51].
 Figure 9: Typical vorticity evolution from the HR ground-truth data.
-We confirmed that the maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$.
-This suggests that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability of the initial conditions.
-Indeed, without DA, the deviation from the ground truth continues to increase over time, while the SRDA effectively suppresses this error growth (Section VI.3).
-VI.2 Accuracy of SR
-We first confirm the accuracy of inference by the SR model, as the SRDA is based on this model.
+The maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$.
+This indicates that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability.
+Without DA, the deviation from the ground truth continues to increase over time, as shown in Section 6.3.
+6.2 Accuracy of SR inference
+We show the accuracy of inference by the SR model, as the SRDA is based on this model.
 Figure 10 shows three sets of snapshots for the ground truth, observations, LR input, and SR output.
-The results indicate that the SR model reproduces the fine filaments and small-scale structures within vortices.
+The SR model reproduces the fine filaments and the small-scale structures within vortices.
 These patterns are not evident in the LR inputs or observations.
@@
 The input data are not indexed by time because each sample was randomly selected for testing without simulating time evolution.
-Flow-field reconstruction has been extensively studied using neural networks.Lluís et al.
-(2020); Maulik et al.
-(2020); Gundersen et al.
-(2021); Wang et al.
-(2022) In these studies, inputs are incomplete, such as spatially sparse sensor measurements, where only a portion of the entire domain is available.
-Neural networks are then trained by comparing their output to complete target data.
-The setup of our problem is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
-The success of the SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution.
+Flow-field reconstruction has been extensively studied using neural networks [74, 75, 76, 77].
+In these studies, inputs are incomplete, such as spatially sparse sensor measurements.
+Neural networks are then trained by comparing their outputs to complete target data.
+Our problem setup is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
+The success of SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution.
 The loss function $l_{\text{SR}}$ in Eq.
 (18) operates at individual grid points, and only the non-missing observed points contribute to $l_{\text{SR}}$.
-This property enables backpropagation even when the target data are missing, as discussed in Section IV.4.
-Furthermore, since convolution shares its kernel across the space (known as weight sharing),Aloysius and Geetha (2017) the efficient learning was realized using the small amount of non-missing observational data.
-Although the inference appears to be accurate, we confirmed that the SR model cannot adequately correct forecast errors.
+This point-wise nature enables backpropagation even when the target data are missing.
+Furthermore, since convolution shares its kernel across space (known as weight sharing) [78], the efficient learning was realized using the small amount of non-missing observational data.
+Although the inference appears to be accurate, the SR model cannot adequately correct forecast errors.
 One might think that the SR output could be used as the analysis due to its high accuracy (Fig.
-8).
-However, we found that test errors accumulated over time when the downsampled SR output was used as the initial condition for the LR fluid model.
+10).
+However, test errors accumulated over time when the downsampled SR output was used as the initial condition for the LR fluid model.
 This result indicates that observations need to be assimilated into forecasts.
-VI.3 Accuracy of LR forecast by SRDA
-We demonstrate that the sequential SRDA can make more accurate inferences than the EnKF. The LR forecast is compared in this section, and the HR analysis is examined next (Section VI.4).
-Figure 9 shows snapshots of the ground truth and forecasts for the EnKF and the SRDA.
-To align the resolution, a low-pass filter was applied to the HR ground-truth vorticity.
+The trained SR model makes non-local inference.
+This non-locality is essential for assimilating observations, as discussed in Section 4.3.
+We use two indicators for interpreting the SR model: gradients and local attribution maps (LAMs) [31].
+Gradients are expressed as $\partial\mathsf{F}(\bm{x})/\partial\bm{x}$, where $\mathsf{F}$ is the trained SR model.
+Such a simple gradient is sometimes not a good indicator for interpreting how a model output depends on its input due to nonlinearity [31, 79].
+LAM, proposed by Gu and Dong [31], combines gradient methods with path integral techniques for interpreting behaviors of nonlinear SR models.
+Figure 11 shows gradient and LAM fields for the three SR samples previously discussed in Fig.
+10.
+We focus on the square regions (black boxes) in the figures.
+The gradients and LAMs indicate that colored points in the LR inputs contribute to the HR outputs in these boxes.
+If there are observed values in these boxes, the colored points in the inputs are updated by minimizing the loss function (Section 4.3).
+Thus, observations have a non-local influence through the trained SR model.
+This is an essential reason for the success of the proposed SRDA.
+Interestingly, the LAMs tend to take larger values over wider regions, compared to the gradients (Fig.
+11), suggesting the nonlinearity of the SR model.
+For instance, in the first row of the figure, the LAM has a larger value at one grid point below the black box, matching the position of the vortex filament.
+Figure 11: Attribution maps for SR.
+The left two columns, SR output and LR input, are the same as in Fig.
+10, but shown here for comparison.
+Local attribution maps (LAMs) [31] were calculated for the small square regions of $11\times 1$ 11 (i.e., black boxes in the figure).
+These boxes were specified on the HR grids, so they are not necessarily aligned with the LR grids.
+The gradients and LAMs were normalized by their maximum values, with ranges between 0 and 1, following Gu and Dong [31].
+6.3 Accuracy of LR forecast
+We demonstrate that the sequential SRDA can make more accurate inferences than the EnKF. The LR forecast is discussed here, and the HR analysis is examined next.
+Figure 12 shows snapshots of the ground truth and forecasts for the EnKF and SRDA.
+To match the resolution, a low-pass filter was applied to the HR ground-truth vorticity.
 The figure also includes LR simulation results without SR or DA.
@@
 Without SR or DA, the vortex shapes and locations still differ from the ground truth.
-Figure 9: Vorticity snapshots of the ground truth and forecasts for the EnKF and the SRDA at $t=2$, 9, and 16.
+Figure 12: Vorticity snapshots of the ground truth and forecasts for the EnKF and SRDA at $t=2$, 9, and 16.
 For comparison, the figure also shows LR simulation results without SR or DA.
-A low-pass filter was applied to the ground-truth snapshots to align the resolution to the LR.
-The quantitative analysis further confirms the accuracy of the SRDA forecast.
-The top panels in Fig.
-10 show the time series of the mean values for the MAE ratio and the MSSIM loss.
-Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (i.e., the positive Lyapunov exponent in Section VI.1).
+A low-pass filter was applied to the ground-truth snapshots to match their resolution to the LR.
+Figure 13 shows the time series of the mean values for the MAE ratio and MSSIM loss.
+Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (Section 6.1).
 In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines).
 The error reduction occurs every time unit since the assimilation interval was set to 1.
-When averaged over time, both time series have two peaks, corresponding to the two developments of small-scale patterns by the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$).
+When smoothed over time, the time series of the EnKF and SRDA have two peaks, corresponding to the two developments of small-scale patterns during the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$).
 The bottom panels in Fig.
-10 show the differences in the test errors between the SRDA and the EnKF, where the gray areas represent the 10th to 90th percentile values.
+13 show the differences in the test errors between the EnKF and SRDA, where the gray areas represent the 10th to 90th percentile values.
 The test errors of the SRDA are significantly smaller than those of the EnKF.
-Thus, we can conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
-Figure 10: Time series of the MAE ratio and the MSSIM loss for the forecast states in the 500 test simulations.
-The top panels show the mean test errors at each time step with an output interval of 0.25.
-The bottom panels show the differences between the SRDA and the EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
-Within the SRDA framework, the LR forecast can be super-resolved to the HR at an arbitrary time by the SR model (Section IV.6).
+Thus, we conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
+Figure 13: Time series of the MAE ratio and MSSIM loss for the forecast states in the 500 test simulations.
+The top panels show the mean test errors.
+The bottom panels show the differences between the SRDA and EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
+In the proposed SRDA, the LR forecast can be super-resolved to the HR at an arbitrary time using the SR model (Section 4.6).
 Figure 14 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast.
 The merging of two vortices is not obvious in the LR, but is well captured in the super-resolved forecast.
-Figure 11: Typical vortex merging in the ground truth, the LR forecasts, and the super-resolved forecast.
-VI.4 Accuracy of HR analysis by SRDA
-We next demonstrate the accuracy of the HR analysis provided by the SRDA.
+Figure 14: Typical vortex merging in the ground truth, the LR forecasts, and the super-resolved forecasts.
+6.4 Accuracy of HR analysis
 Figure 15 shows snapshots of the ground truth and analyses.
 During the jet collapse ($t=2$), the filament structure appears much more distinct in the SRDA analysis than in the EnKF.
-Furthermore, during the vortex merging ($t=9$ and 11), the SRDA snapshots exhibit sharper vortex shapes and clearer internal structures compared to the EnKF.
-Figure 12: Vorticity snapshots of the ground truth and analyses for the EnKF and the SRDA at $t=2$, 9, and 11.
-The quantitative analysis also confirms the accuracy of the SRDA analysis.
-Figure 13 shows the time series of the MAE ratio and the MSSIM loss, as in Fig.
-10.
+Furthermore, during the vortex merging ($t=9$ and 11), the SRDA infers sharper vortex shapes and clearer internal structures than the EnKF.
+Figure 15: Vorticity snapshots of the ground truth and analyses for the EnKF and SRDA at $t=2$, 9, and 11.
+Figure 16 shows the time series of the MAE ratio and MSSIM loss, as in Fig.
+13.
 The dots represent the assimilation time steps, with assimilation intervals at 1.
-The SRDA and EnKF analyses are inferred only at these times.
-The EnKF time series (orange dashed lines) show two peaks, corresponding to the two developments of small-scale patterns linked to the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$).
-These peaks are less apparent in the SRDA time series (blue solid lines), especially for the first peak.
-This result suggests that the SRDA effectively mitigates the error increase during the jet collapse.
+The EnKF and SRDA analyses are inferred only at these times.
+The EnKF time series (orange dashed lines) show two peaks, corresponding to the two developments of small-scale patterns during the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$).
+These peaks are less apparent in the SRDA time series (blue solid lines).
 The bottom panels in Fig.
-13 indicate that the MAE ratio and the MSSIM loss for the SRDA are significantly smaller than those of the EnKF.
-Thus, we can conclude that the SRDA analysis is more accurate than the EnKF analysis, in line with the forecast results (Section VI.3).
-Figure 13: Time series of the MAE ratio and the MSSIM loss for the analysis states in the 500 test simulations.
+16 indicate that the test errors of the SRDA are significantly smaller than those of the EnKF.
+Thus, we conclude that the SRDA analysis is more accurate than the EnKF analysis.
+Figure 16: Time series of the MAE ratio and MSSIM loss for the analysis states in the 500 test simulations.
 These states were inferred only at the analysis time steps, denoted by dots.
 The top panels show the mean test errors.
-The bottom panels show the differences between the SRDA and the EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
-We compare uncertainties estimated by the EnKF and the SRDA.
-Here, standard deviation (SD) is used as an indicator of uncertainty.
+The bottom panels show the differences between the SRDA and EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
+We compare uncertainties estimated by the EnKF and SRDA, using standard deviation (SD) as an indicator.
+For the EnKF, SD at each grid point was calculated from all ensemble members (i.e., 300 members).
 For the SRDA, SD was directly calculated by taking the square root of the estimated variances in Eq.
-(12), while for the EnKF, SD at each grid point was calculated from all ensemble members (i.e., 300 members).
-Figure 14 compares distributions of the analysis values and their SDs for the EnKF and the SRDA.
+(12).
+Figure 17 compares the analyses and their SDs for the EnKF and SRDA.
 The SRDA SD tends to be large around vortex filaments or vortex edges.
-This tendency appears to be most evident in the jet collapse ($t=2$).
+This tendency appears to be most evident during the jet collapse ($t=2$).
+This result is reasonable because vortex locations are generally uncertain, and small changes in their locations lead to large differences in vorticity.
 The EnKF SD also exhibits a similar tendency; however, the SD distribution is blurred, likely reflecting the blurred vortex patterns.
 The SD value ranges also differ: the EnKF SD is between 0.3 and 0.9, while the SRDA SD is between 0.8 and 0.9.
-Related to this range difference, the SRDA tends to overestimate SD in regions where vorticity is close to zero, which can be found at $t=9$ and 11 in Fig.
-14.
-In contrast, the EnKF SD tends to be small in small-vorticity regions.
-This overestimation of the SRDA may be due to a simplified variance estimation in the present CVAE (Section V.3), where the variances are computed using the SR model from the LR analysis.
-Therefore, although the SRDA demonstrated the potential for uncertainty quantification without relying on ensemble calculation, further model development is necessary.
-Recent studies demonstrate that VAEs can learn error distributions of model forecasts,Grooms (2021); Yang and Grooms (2021); Grooms et al.
-(2023); Xiao et al.
-(2024) and their model architectures may be effective for the SRDA.
-Figure 14: Vorticity snapshots of the analyses and their standard deviations (SDs) for the EnKF and the SRDA at $t=2$, 9 and 11.
-The analyses shown here are the same as in Fig.
-12.
-To compare them with their SDs, these panels are displayed again.
-VI.5 Computation time of SRDA
-Lastly, we confirm that the computational time of the SRDA is much less than that of the EnKF.
+Compared to the EnKF, the SRDA tends to estimate larger SD in regions where vorticity is close to zero, especially at $t=9$ and 11 in Fig.
+17.
+Although the true value of SD is unknown, our results suggest that the SRDA tends to estimate larger SD in low-vorticity regions.
+This tendency may be due to a simplified variance estimation in the present CVAE (Section 5.3), where the variances are computed using the SR model from the LR analysis.
+The SRDA demonstrated a potential for uncertainty quantification without relying on ensemble calculation; however, further model development is necessary.
+Recent studies demonstrate that VAEs can learn error distributions of model forecasts [22, 23, 24, 25].
+Their model architectures may be effective for the SRDA.
+Figure 17: Vorticity snapshots of the analyses and their standard deviations (SDs) for the EnKF and SRDA at $t=2$, 9 and 11. The analyses shown here are the same as in Fig. 15.
+6.5 Computation time of SRDA
+The computational time of the SRDA is much less than that of the EnKF.
 The elapsed times for the simulations from $t=0$ to 20 were measured using a single process on a CPU (Intel Xeon Gold 6326).
@@
 This difference is mainly due to the calculation of the ensemble evolution, which is not required in the SRDA.
-VII Conclusions
-This study has proposed a theory of unsupervised SRDA using the CVAE.
-We have shown that a new DA framework using the ELBO encompasses the 3D-Var formalism (Section III).
-Since ELBO serves as the loss function for CVAEs, the assimilation can be executed using the CVAE.
-In this process, it is essential to use a non-diagonal background error covariance matrix to incorporate the spatial correlations of forecast states into the CVAE, similar to the 3D-Var.
-Instead of using such non-diagonal matrices, we have further extended the framework by using the non-local nature of SR operations to achieve DA (Section IV).
-In other words, SR operators can serve as background error covariance matrices.
-This result provides a theoretical foundation for combining SR and DA.
-The combination of SR models and CVAEs enables the assimilation of HR observations into LR forecasts that are computed with LR fluid models.
-This approach eliminates the need for numerical integration using HR fluid models, thereby reducing computational burden.
-Specifically, the CVAE’s encoder estimates the HR analysis from the input of LR forecasts and HR observations, while the decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis.
-The resulting HR analysis is then downsampled to the LR using algebraic methods, such as linear interpolation, serving as the initial condition for the LR fluid model in the next assimilation cycle.
-The effectiveness of the proposed method was evaluated through numerical experiments using an idealized barotropic ocean jet system (Sections V and VI).
+7 Conclusions
+This study has proposed a theory of unsupervised SRDA with the CVAE.
+Using the ELBO, we have developed an unsupervised theory of DA, which is an extension of the 3D-Var (Section 3).
+Since ELBOs generally serve as the loss function for CVAEs, the assimilation is feasible using the CVAE.
+In this theory, it is essential to use a non-diagonal background error covariance matrix, as in the 3D-Var.
+Instead of using such non-diagonal matrices, we have extended the theory by using the non-locality of SR to perform DA (Section 4).
+This extension provides a theoretical foundation for combining SR and DA.
+The combination of SR models and CVAEs enables us to assimilate HR observations into LR forecasts computed with LR fluid models.
+This approach eliminates the need for time integration of HR fluid models, thereby reducing computational burden.
+The effectiveness of the proposed method was evaluated via numerical experiments using an idealized barotropic ocean jet system (Sections 5 and 6).
 Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times.
-These results suggest the potential of SRDA as an efficient approach for integrating observations into CFD simulations.
-There are several future research topics, as SRDA has been recently developed.
-Although the proposed theory enables learning of nonlinear observation operators, the proof of concept in the present study has examined only the linear observation operator.
-Furthermore, advanced training methods, such as online trainingRasp (2020) and loss balancing for VAEs,Asperti and Trentin (2020) will be important to develop SRDA, whereas a simple offline training method was employed in our study (Section V.4).
-These points will be addressed in a more realistic experimental setup.
-In practical applications, data partitioning will be important because the data size for three-dimensional flows is larger, which would require the division of snapshots into small patches.Ouala et al.
-(2018); Yang and Grooms (2021) Non-Gaussianity is also important in more realistic DA, which has recently been addressed by VAEsXiao et al.
-(2024) and normalizing flows.Xie et al.
-(2023) The theory proposed in our study will provide a foundation for using such advanced techniques, leading to further development of SRDA.
-Acknowledgements.
+Our results suggest the potential of SRDA as a promising approach for integrating observations into CFD simulations while obtaining HR inferences with a low computational cost.
+There are several future research topics for SRDA.
+The proof of concept here has only examined the linear observation operator, although the proposed SRDA theoretically allows learning of nonlinear observation operators.
+Furthermore, advanced training methods, such as online training [63, 64] and loss balancing for VAEs [68], will be important to develop SRDA.
+As for generalizability, we have not evaluated to what extent the CVAE remains effective when the statistical properties of the input data differ.
+These aspects should be addressed in a more realistic experimental setup.
+Indeed, the fluid system used here is simple compared to actual atmospheric and oceanic systems.
+It is necessary to evaluate the proposed SRDA for more complex systems.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgements
 This work used computational resources of the TSUBAME3.0 supercomputer provided by the Institute of Science Tokyo through the HPCI System Research Project (Project ID: hp220102) to conduct the CFD simulations.
@@
 This paper is based on results obtained from a project, JPNP22002, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
-Data Availability Statement
-The data and source code that support the findings of this study are preserved and openly available at https://github.com/YukiYasuda2718/srda-cvae.
 Appendix A Hyperparameters of the SR model and the CVAE
-We list here the hyperparameters of the SR model and the CVAE shown in Fig.
-5.
-Table 2 lists the hyperparameters for the convolution layers in the SR model and the CVAE.
+We list the hyperparameters of the SR model and the CVAE shown in Fig.
+7.
+Table 2 shows the numbers of channels and strides in all convolution layers.
 The kernel size of all convolutions is 3.
-The slope of all leaky ReLUMaas et al.
-(2013) is $-0.01$.
-The upscale factor of all pixel shuffle operationsShi et al.
-(2016) is 2.
+The slope of all leaky ReLU [59] is $-0.01$.
+The upscale factor of all pixel shuffle operations [57] is 2.
 The resizing in the SR model is performed by bicubic interpolation.
-The weight parameters of all convolution layers are randomly initialized using uniform distributions.He et al.
-(2015)
+The weight parameters of all convolution layers are initialized randomly using uniform distributions [80].
 Table 2: Hyperparameters for the convolution layers in the SR model and the CVAE. The layer names are defined in Fig. 7.
@@
 Conv2D-17 256 256 1
+Data availability
+The data and source code that support the findings of this study are preserved and openly available at our Zenodo repository (https://doi.org/10.5281/zenodo.13958042).
+These files are also available at our GitHub repository (https://github.com/YukiYasuda2718/srda-cvae/releases/tag/v0.2.1).
+Declaration of Generative AI and AI-assisted technologies in the writing process
+During the preparation of this work, the authors used Claude 3.5 only for English editing.
+After using this service, the authors reviewed and edited the content as needed.
+The authors take full responsibility for the content of the publication.

### v5

@@
-Theory of Unsupervised Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Estimating Background Covariances via Super-Resolution
+Unsupervised Super-Resolution Data Assimilation Using Conditional Variational Autoencoders with Estimating Background Covariances via Super-Resolution
 Abstract
 This study proposes a theory of unsupervised super-resolution data assimilation (SRDA) using conditional variational autoencoders (CVAEs).
-The theory is based on an evidence lower bound that leads to an objective function for unsupervised learning.
-This function can be reduced to the objective function of a traditional data assimilation method, namely the three-dimensional variational (3D-Var) formalism.
-Similar to 3D-Var, it is essential to use a non-diagonal background error covariance matrix to assimilate distant observations.
-Instead of using such a non-diagonal matrix, we extend the proposed theory by using the non-locality of super-resolution (SR).
-For linear SR, SR operators serve as background error covariance matrices.
-For nonlinear SR, error backpropagation through SR neural networks induces covariance structures in inference.
-SRDA can be performed using CVAEs because the loss function for CVAEs is generally an evidence lower bound.
-By incorporating the SR neural network into the CVAE, the encoder estimates the high-resolution (HR) analysis from HR observations and low-resolution (LR) forecasts.
-The decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis.
+We derive an evidence lower bound for unsupervised learning, showing that our theory is an extension of a traditional data assimilation (DA) method, namely the three-dimensional variational (3D-Var) formalism.
+In contrast to 3D-Var, our theory exploits the non-locality of super-resolution (SR) to learn background covariances without explicitly imposing them for assimilating distant observations.
+For linear SR, SR operators serve as background error covariance matrices,whereas for nonlinear SR, error backpropagation through SR neural networks induces covariance structures in inference.
+SRDA can naturally be realized with CVAEs because the loss function for CVAEs is generally an evidence lower bound.
+By incorporating the SR neural network into the CVAE, the encoder estimates the high-resolution (HR) analysis from HR observations and low-resolution forecasts.
+The decoder acts as the observation operator by reconstructing the HR observations from the estimated HR analysis.
 The effectiveness of SRDA was evaluated through numerical experiments using an idealized barotropic ocean jet system.
@@
 The findings of this study provide a theoretical basis for integrating SR and DA, which will stimulate further research in this direction.
-keywords:
-super-resolution , data assimilation , unsupervised learning , evidence lower bound , background error covariance , variational autoencoder
-journal: Comput.
-Methods Appl.
-Mech.
-Engrg.corresponding_author_infocorresponding_author_infofootnotetext: Corresponding author.
-2-12-1-i9-111 Ookayama, Meguro-ku, Tokyo, 1528550, Japan.
-Tel.: +81 3 57343173 (office).
-Email address: yasuda.yuki@scrc.iir.isct.ac.jp
-[…]
-[inst1] organization=Supercomputing Research Center, Institute of Integrated Research, Institute of Science Tokyo, addressline=2-12-1 Ookayama, Meguro-ku, city=Tokyo, postcode=1528550, country=Japan
-1 Introduction
-Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD) [1, 2].
-Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time.
+preprint: AIP/123-QED
+I Introduction
+Data assimilation (DA) is an essential technique for various modeling in computational fluid dynamics (CFD).Asch, Bocquet, and Nodet (2016); Carrassi et al.
+(2018) Fluid systems generally exhibit chaotic behavior, where errors in initial conditions grow over time.
 Uncertainties in model parameters may also lead to discrepancies from reality.
@@
 Recent advances in deep learning are impacting various scientific fields and are also found in DA.
-One reason is that both DA and deep learning use backpropagation based on adjoint equations [3, 4, 5].
-Another reason is the use of the Bayesian approach [2, 5, 6].
-This approach models the true states of fluid systems by probability distributions based on known data, such as observations and fluid model outputs.
-A typical neural network using the Bayesian approach is the variational autoencoder (VAE) [7, 8, 9].
-Like standard autoencoders [10], a VAE consists of an encoder and a decoder, both of which are implemented by neural networks.
+One reason is that DA and deep learning use backpropagation based on adjoint equations.Abarbanel, Rozdeba, and Shirman (2018); Chen et al.
+(2018); Geer (2021) Another reason is the use of the Bayesian approach.Carrassi et al.
+(2018); Geer (2021); Bocquet et al.
+(2020) This approach models the true states of fluid systems by probability distributions based on known data, such as observations and fluid model outputs.
+A typical neural network using the Bayesian approach is the variational autoencoder (VAE).Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014) Like standard autoencoders,Dong et al.
+(2018) a VAE consists of an encoder and a decoder, both of which are implemented by neural networks.
 Typically, the encoder transforms the input data into a compressed representation in the latent space, while the decoder reconstructs the original data from these latent variables.
-A major extension of VAEs is the conditional VAE (CVAE), which is based on conditional probability modeling [11, 12].
-This conditioning allows us to model target variables, such as the true states of systems, conditioned on known data, such as observations.
+A major extension of VAEs is the conditional VAE (CVAE), which is based on conditional probability modeling.Kingma et al.
+(2014); Sohn, Yan, and Lee (2015) This conditioning allows us to model target variables, such as the true states of systems, conditioned on known data, such as observations.
 Applications of VAEs in DA can be categorized into two groups.
-The first group employs VAEs for dimensionality reduction [13, 14, 15].
-In this approach, the encoder maps fluid model states into a low-dimensional latent space.
+The first group employs VAEs for dimensionality reduction.Liu, Grana, and de Figueiredo (2021); Mohd Razak et al.
+(2022); Melinc, Boštjan and Zaplotnik, Žiga (2024) In this approach, the encoder maps fluid model states into a low-dimensional latent space.
 Statistical DA methods, such as Kalman filters, are then applied in this space.
 Due to the reduced dimensionality, the computational cost for DA is significantly lowered.
-This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks [16, 17, 18, 19, 20, 21].
-The second group employs VAEs as generators that learn the probability distributions of flow fields [22, 23, 24].
-In this approach, a fluid model state is encoded, and noise is then added in the latent space, creating various potential latent variables.
+This scheme is referred to as latent DA and has been extensively studied using various encoder-decoder networks.Mack et al.
+(2020); Amendola et al.
+(2021); Peyron et al.
+(2021); Cheng et al.
+(2022); Liu et al.
+(2022); Cheng et al.
+(2024) The second group employs VAEs as generators that learn the probability distributions of flow fields.Grooms (2021); Yang and Grooms (2021); Grooms et al.
+(2023) In this approach, a fluid model state is encoded, and noise is then added in the latent space, creating various potential latent variables.
 Ensemble members are obtained by decoding these latent variables back into the physical space.
@@
 We proposes a new method, categorized into the third group, where VAEs directly perform DA.
-In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state [1, 2].
-Generally, observations represent a state close to reality, but they are spatially sparse.
+In statistical DA methods, such as Kalman filters, the estimation of background error covariance matrices is essential for inferring the true state.Asch, Bocquet, and Nodet (2016); Carrassi et al.
+(2018) Generally, observations represent a state close to reality, but they are spatially sparse.
 In contrast, fluid model outputs (i.e., background states) include larger errors, but they are spatially dense.
@@
 Conversely, “local” means that an inferred value is computed only from the input values at the same point.
-The estimation of covariance matrices is also important for DA using autoencoders (including VAEs).
-In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown [15, 19].
-More generally, the background error distributions are non-Gaussian and can be estimated by VAEs [25].
-The non-Gaussian distributions learned by VAEs are then incorporated into a traditional DA method [25], namely, the three-dimensional variational (3D-Var [26]) formalism.
+The estimation of covariance matrices is also important for DA using autoencoders, including VAEs.
+In latent DA, the covariance structure needs to be estimated because the statistics of the latent space are typically unknown.Melinc, Boštjan and Zaplotnik, Žiga (2024); Cheng et al.
+(2022) More generally, the background error distributions are non-Gaussian and can be estimated by VAEs.Xiao et al.
+(2025) The non-Gaussian distributions learned by VAEs are then incorporated into a traditional DA method,Xiao et al.
+(2025) namely, the three-dimensional variational (3D-Var)Lorenc (1986) formalism.
 In the present study, we focus on the non-locality of super-resolution (SR) and utilize SR operators for estimating covariance structures.
 SR is originally a technique to enhance the resolution of images.
-In computer vision, deep learning-based SR has been actively studied [27, 28, 29].
-A key characteristic of SR is its non-local nature [30, 31, 32, 33]: neural networks focus on input pixels at low resolution (LR) that are distant from a target pixel at high resolution (HR), but correlated to this target, generating high-quality HR images.
-Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA) [34, 35].
-One reason for this integration is the increasing availability of HR observations [36, 37, 38].
-In typical DA methods, HR observations are assimilated using HR fluid models [39, 40], which require a large amount of computational resources for time integration.
+In computer vision, deep learning-based SR has been actively studied.Ha et al.
+(2019); Anwar, Khan, and Barnes (2020); Lepcha et al.
+(2023) A key characteristic of SR is its non-local nature:Dong et al.
+(2014); Gu and Dong (2021); Chen et al.
+(2023); Zhou et al.
+(2023) neural networks focus on input pixels at low resolution (LR) that are distant from a target pixel at high resolution (HR), but correlated to this target, generating high-quality HR images.
+Recently, SR has been integrated into DA, referred to as super-resolution data assimilation (SRDA).Barthélémy et al.
+(2022); Yasuda and Onishi (2023) One reason for this integration is the increasing availability of HR observations.Imaoka et al.
+(2010); Durand et al.
+(2010); Cifelli et al.
+(2018) In typical DA methods, HR observations are assimilated using HR fluid models,Li, Wang, and Fu (2019); Honda et al.
+(2022) which require a large amount of computational resources for time integration.
 In contrast, in SRDA frameworks, background states are computed with LR fluid models.
@@
 One advantage of combining SR with DA is that advanced techniques developed in SR can be applied in DA.
-Numerous SR studies have been conducted not only in computer vision [27, 28, 29] but also in fluid dynamics [41].
-In contrast, DA using deep learning has only recently become active [42].
-As for SRDA, several research questions remain, such as those regarding its theoretical background.
-Indeed, previous studies on SRDA [34, 35] did not utilize the non-locality of SR.
+Numerous SR studies have been conducted not only in computer visionHa et al.
+(2019); Anwar, Khan, and Barnes (2020); Lepcha et al.
+(2023) but also in fluid dynamics.Fukami, Fukagata, and Taira (2023) In contrast, DA using deep learning has only recently become active.Cheng et al.
+(2023) As for SRDA, several research questions remain, such as those regarding its theoretical background.
+Indeed, previous studies on SRDABarthélémy et al.
+(2022); Yasuda and Onishi (2023) did not utilize the non-locality of SR.
 If theoretical backgrounds for combining SR with DA are clarified, further integration of both will be promoted.
@@
 This result provides a theoretical foundation for combining SR with DA.
-We first review an evidence lower bound (ELBO) for DA (Section 2), which serves as the loss function for unsupervised learning.
-We then develop a new theory of DA using the ELBO and show that this theory is an extension of the 3D-Var (Section 3).
+We first review an evidence lower bound (ELBO) for DA (Section II), which serves as the loss function for unsupervised learning.
+We then develop a new theory of DA using the ELBO and show that this theory is an extension of the 3D-Var (Section III).
 Similar to the 3D-Var, our theory requires non-diagonal background error covariance matrices.
-We further extend our theory by incorporating SR for non-local inference, which eliminates the need to estimate non-diagonal covariance matrices (Section 4).
-To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section 5).
-The results are analyzed through comparison with an ensemble Kalman filter (Section 6).
-Conclusions are presented in Section 7.
-2 Review of ELBO for DA
-We reformulate an evidence lower bound (ELBO) [43, 44, 45], which is the objective function in the proposed theory. ELBO is also known as the variational lower bound.
+We further extend our theory by incorporating SR for non-local inference, which eliminates the need to estimate non-diagonal covariance matrices (Section IV).
+To evaluate the proposed SRDA, numerical experiments are conducted using an idealized ocean jet system (Section V).
+The results are analyzed through comparison with an ensemble Kalman filter (Section VI).
+Conclusions are presented in Section VII.
+II Review of ELBO for DA
+We reformulate an evidence lower bound (ELBO),Ormerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al.
+(2017) which is the objective function in the proposed theory.
+ELBO is also known as the variational lower bound.
 The present study considers assimilation at a single time step.
@@
 Figure 1: Probabilistic model expressed by Eq. (1). All variables are in the physical space.
-The ELBO is derived by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$ [43, 44, 45]:
+The ELBO is derived by introducing an approximate posterior distribution $q(\bm{z}\mid\bm{x},\bm{y})$:Ormerod and Wand (2010); Zhu, Chen, and Xing (2014); Ghimire et al. (2017)
 […]
@@
 The final line gives the ELBO ($=-l$), which is a lower bound for the log-likelihood, $\ln p\left(\bm{y}\mid\bm{x}\right)$.
-DA is usually performed via maximum likelihood estimation [1, 2].
-Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], we maximize the ELBO, or equivalently minimize $l$ in Eq.
+DA is usually performed via maximum likelihood estimation.Asch, Bocquet, and Nodet (2016); Carrassi et al.
+(2018) Instead of maximizing the log-likelihood [i.e., $\ln p\left(\bm{y}\mid\bm{x}\right)$], we maximize the ELBO, or equivalently minimize $l$ in Eq.
 (2c) to obtain $q(\bm{z}\mid\bm{x},\bm{y})$.
 The true state $\bm{z}$ is then inferred from $\bm{x}$ and $\bm{y}$ through $q(\bm{z}\mid\bm{x},\bm{y})$.
-3 DA using the ELBO under Gaussianity
+III DA using the ELBO under Gaussianity
 Starting from the ELBO, we derive a new objective function that can be used to train CVAEs in an unsupervised manner.
 This loss function requires non-diagonal background error covariance matrices, similar to those in 3D-Var.
-3.1 Transformation of the ELBO under Gaussianity
-To transform the ELBO, the following Gaussian distributions are introduced:
+III.1 Transformation of the ELBO under Gaussianity
+We make two assumptions to derive the loss function for DA: (A) the ELBO in Eq. (2c) and (B) the following Gaussian distributions:
 $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\;\exp\left(-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}(\bm{x},\bm{y})\right\rVert_{V(\bm{x},\bm{y})^{-1}}^{2}\right),$ (3a)
@@
 In Eq.
-(4b), the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trick [8]), where each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ as in Eq.
+(4b), the expected value is approximated by the sample mean using a Monte Carlo method (i.e., the reparameterization trick),Kingma and Welling (2014) where each realization $\bm{\hat{z}}$ is sampled from $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ as in Eq.
 (5).
-In the actual implementation, the sample size $n_{\text{sample}}$ can be set to 1 if the batch size is sufficiently large [8].
-Equation (4b) indicates that the reconstruction error measures the difference between $\bm{y}$ and $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
+In the actual implementation, the sample size $n_{\text{sample}}$ can be set to 1 if the batch size is sufficiently large.Kingma and Welling (2014) Equation (4b) indicates that the reconstruction error measures the difference between $\bm{y}$ and $\mathsf{H}(\bm{\hat{z}})$ in the observation space.
 The KL divergence, the second term in Eq. (2c), is transformed as follows:
-$\mathbb{E}_{q}$ $\left[\ln\frac{q\left(\bm{z}\mid\bm{x},\bm{y}\right)}{p\left(\bm{z}\mid\bm{x}\right)}\right]$
 […]
@@
 By substituting Eqs. (4b) and (6c) into Eq. (2c), we obtain the ELBO under Gaussianity:
-$l_{\text{Gauss}}$ $=$ $\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}$ (7)
-$\quad+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$
-where constant terms are omitted and the sample size $n_{\rm sample}$ in Eq. (4b) is set to 1.
-3.2 Reduction of the ELBO to the 3D-Var objective function
+$l_{\text{Gauss}}$ $=$ $\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\mathsf{A}(\bm{x},\bm{y})-\bm{x}\right\rVert_{B^{-1}}^{2}+\frac{1}{2}\left\{\operatorname{tr}\left[B^{-1}V(\bm{x},\bm{y})\right]-\ln\left[\frac{\det V(\bm{x},\bm{y})}{\det B}\right]\right\},$ (7)
+where constant terms are omitted and the sample size $n_{\text{sample}}$ in Eq. (4b) is set to 1.
+III.2 Reduction of the ELBO to the 3D-Var objective function
 We show that $l_{\text{Gauss}}$ can be reduced to the 3D-Var objective function. This suggests that DA can be performed by minimizing $l_{\text{Gauss}}$.
-We start by setting the covariance $V(\bm{x},\bm{y})$ to a constant matrix, where each component is quite small.
+To show this fact, we make further one assumption: the covariance $V(\bm{x},\bm{y})$ is a constant matrix, where each component is quite small.
 The variable $\bm{\hat{z}}$ in Eq.
@@
 $l_{\text{3D-Var}}=\frac{1}{2}\left\lVert\bm{y}-\mathsf{H}(\bm{a})\right\rVert_{R^{-1}}^{2}+\frac{1}{2}\left\lVert\bm{a}-\bm{x}\right\rVert_{B^{-1}}^{2},$ (8)
-where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from $\bm{x}$ and $\bm{y}$.
+where $\bm{a}$ denotes the analysis vector given by the function $\mathsf{A}$ from $\bm{x}$ and $\bm{y}$ [i.e., $\bm{a}=\mathsf{A}(\bm{x},\bm{y})$].
 This $\bm{a}$ is obtained by minimizing $l_{\text{3D-Var}}$.
@@
 (7)], the ELBO describes the balance between $\bm{x}$ and $\bm{y}$ in estimating the true state.
-The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is a matrix $H$ [1].
+The minimizer of $l_{\text{3D-Var}}$ is easily derived when the observation operator is a matrix $H$:
 $\bm{a}$ $=$ $\bm{x}+K\left(\bm{y}-H\bm{x}\right),$ (9a)
 $K$ $=$ $BH^{\text{T}}\left(R+HBH^{\text{T}}\right)^{-1},$ (9b)
-where $K$ is known as the Kalman gain matrix.
-These equations describe the non-locality of DA: an observation at one grid point is incorporated into the surrounding inferred values.
+where $K$ is known as the Kalman gain matrix.Asch, Bocquet, and Nodet (2016) These equations describe the non-locality of DA: an observation at one grid point is incorporated into the surrounding inferred values.
 This non-locality typically originates from the non-zero off-diagonal elements of the covariance matrix $B$.
 If $B$ is diagonal (i.e., all off-diagonal elements are zero), DA becomes local inference, and spatially sparse observations are incorporated into the inference only at their located grid points.
-3.3 Implementation of DA using CVAE
+III.3 Implementation of DA using CVAE
 DA can be performed by minimizing $l_{\text{Gauss}}$ in Eq.
@@
 We propose an efficient approach to this minimization using a trained CVAE.
-The loss function for VAEs, including CVAEs, is generally an ELBO [7, 8, 9, 11, 12].
-This can be understood from the ELBO formula [Eq.
-(7)].
-To evaluate $l_{\text{Gauss}}$, the analysis $\bm{a}$ and its covariance $V$ are first computed from given $\bm{x}$ and $\bm{y}$; the observation vector $\bm{\hat{y}}$ [$=\mathsf{H}(\bm{\hat{z}})$] is then reconstructed from $\bm{a}$ and $V$ via Eq.
+The loss function for VAEs, including CVAEs, is generally an ELBO.Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al.
+(2014); Sohn, Yan, and Lee (2015) This can be understood from the ELBO formula [Eq.
+(7)] and naturally leads to the use of CVAEs rather than other models (e.g., generative adversarial networksCreswell et al.
+(2018)) that do not have ELBOs as their loss functions.
+To evaluate $l_{\text{Gauss}}$ in Eq.
+(7), the analysis $\bm{a}$ and its covariance $V$ are first computed from given $\bm{x}$ and $\bm{y}$; the observation vector $\bm{\hat{y}}$ [$=\mathsf{H}(\bm{\hat{z}})$] is then reconstructed from $\bm{a}$ and $V$ via Eq.
 (5).
 The value of $l_{\text{Gauss}}$ decreases as $\bm{\hat{y}}$ becomes closer to the input $\bm{y}$.
-The architecture that reconstructs the input is a characteristic of autoencoders [10].
-The second and third terms in Eq.
+The architecture that reconstructs the input is a characteristic of autoencoders.Dong et al.
+(2018) The second and third terms in Eq.
 (7) are the conditions for the estimated true state $\bm{\hat{z}}$.
@@
 The covariance matrices $R$ and $B$ in $l_{\text{Gauss}}$ are treated as prescribed parameters (i.e., hyperparameters).
-Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices (i.e., having zero off-diagonal elements) [7, 8, 9, 11, 12].
-However, the off-diagonal elements of $B$ are essential for the non-local inference in DA (Section 3.2).
+Since it is difficult to determine many hyperparameters, these matrices are usually assumed to be proportional to identity matrices (i.e., having zero off-diagonal elements).Kingma and Welling (2019, 2014); Rezende, Mohamed, and Wierstra (2014); Kingma et al.
+(2014); Sohn, Yan, and Lee (2015) However, the off-diagonal elements of $B$ are essential for the non-local inference in DA (Section III.2).
 We adopt super-resolution (SR) to realize non-local inference using CVAEs, leading to a new theory of SRDA.
-4 SRDA using the ELBO
-4.1 Integration of SR into DA
+IV SRDA using the ELBO
+IV.1 Integration of SR into DA
 To develop a theory of SRDA, we first specify the resolutions of a background state $\bm{x}$, observations $\bm{y}$, and the true state $\bm{z}$: $\bm{x}$ is defined at LR grid points, while $\bm{y}$ and $\bm{z}$ are defined at HR grid points.
@@
 These missing grid points are excluded when arranging observations into the column vector $\bm{y}$.
-We modify the probability distributions in Eqs. (3a)–(3c) for SRDA:
-$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}$ (10a)
-$\times$ $\exp\Bigg{\{}-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\Bigg{\}},$
+We make two assumptions to derive the loss function for SRDA: (A) the ELBO in Eq.
+(2c) and (B’) the following Gaussian distributions.
+The first assumption is the same as in the theory for DA in Section III, while the second assumption is a modification from Eqs.
+(3a)–(3c):
+$q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ $=$ $\frac{1}{\sqrt{(2\pi)^{n_{z}}\det V(\bm{x},\bm{y})}}\;\exp\Bigg{\{}-\frac{1}{2}\left\lVert\bm{z}-\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})\right\rVert^{2}_{V(\bm{x},\bm{y})^{-1}}\Bigg{\}},$ (10a)
 $p\left(\bm{z}\mid\bm{x}\right)$ $=$ $\frac{1}{(\sqrt{2\pi b})^{n_{z}}}\;\exp\left\{-\frac{1}{2b}\left\lVert\bm{z}-\mathsf{F}(\bm{x})\right\rVert^{2}\right\},$ (10b)
@@
 (10a): the LR analysis is first calculated and then super-resolved.
-The ELBO for SRDA is obtained by substituting Eqs. (10a)–(10c) into Eq. (2c), as in Section 3.1:
-$l_{\text{SRDA}}$ $=$ $\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})-\mathsf{F}(\bm{x})\right\rVert^{2}$ (13)
-$\quad+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$
+The ELBO for SRDA is obtained by substituting Eqs. (10a)–(10c) into Eq. (2c), as in Section III.1:
+$l_{\text{SRDA}}$ $=$ $\frac{1}{2r}\left\lVert\bm{y}-\mathsf{H}(\bm{\hat{z}})\right\rVert^{2}+\frac{1}{2b}\left\lVert\mathsf{A}_{\text{HR}}(\bm{x},\bm{y})-\mathsf{F}(\bm{x})\right\rVert^{2}+\frac{1}{2}\sum_{i=1}^{n_{z}}\left\{\frac{v_{i}(\bm{x},\bm{y})}{b}-\ln\left[\frac{v_{i}(\bm{x},\bm{y})}{b}\right]\right\},$ (13)
 where
@@
 As in Eq.
-(7), the Monte Carlo method (i.e., the reparameterization trick [8]) is applied in Eq.
+(7), the Monte Carlo method (i.e., the reparameterization trick)Kingma and Welling (2014) is applied in Eq.
 (13) with the sample size being 1.
@@
 The ELBO (i.e., $l_{\text{SRDA}}$) describes the balance between the observations and the super-resolved background state in estimating the HR true state.
-4.2 Case of linear SR
-The effects of non-locality of SR are examined separately for linear and nonlinear SR. For linear SR, we show that SR operators act as background error covariance matrices.
-Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning [47].
-When generation processes of LR images from HR images are linear (e.g., subsampling or blurring), the inverse processes, namely the SR processes, are also linear [47].
-An SR operator is denoted by a matrix $F$ of size $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) that maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
-To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section 3.2.
+IV.2 Case of linear SR
+The loss function in Eq.
+(13) is applicable to both linear and nonlinear SR.
+This section and the next (Section IV.3) are for interpreting this loss function through examining the effects of non-locality of SR in the linear and nonlinear regimes, respectively.
+Here, we show that linear SR operators act as background error covariance matrices.
+Linear SR operators were frequently used in SR studies, especially in theoretical studies, before the emergence of deep learning.Park, Park, and Kang (2003) When generation processes of LR images from HR images are linear (e.g., subsampling or blurring), the inverse processes, namely the SR processes, are also linear.Park, Park, and Kang (2003) An SR operator is denoted by a matrix $F$ of size $n_{z}\times n_{x}$ ($n_{z}>n_{x}$) that maps $n_{x}$-dimensional LR vectors into $n_{z}$-dimensional HR vectors.
+To simplify the discussion, we take the variances $v_{i}$ ($i=1,\ldots,n_{z}$) to be constant and small values, as in Section III.2.
 The estimated true state $\bm{\hat{z}}$ in Eq.
@@
 By taking an inner product $\bm{\hat{w}}^{\text{T}}\bm{\cdot}\bm{\hat{w}}$, the two blue pixels are correlated with each other.
-4.3 Case of nonlinear SR
+IV.3 Case of nonlinear SR
 In practice, SR operators are implemented as neural networks and are thus nonlinear.
@@
 Eqs.
-(17a) and (17b) are similar to backpropagation formulas using mean squared errors as a loss function [48].
-The differences from the observations (i.e., errors), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagate backward through the SR neural network $\mathsf{F}$.
-This backpropagation updates the current solution $\bm{a}^{(n)}$ (Fig.
+(17a) and (17b) are similar to backpropagation formulas that use mean squared errors as a loss function.Goodfellow, Bengio, and Courville (2016) The differences from the observations (i.e., errors), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagate backward through the SR neural network $\mathsf{F}$.
+This backpropagation updates the current solution $\bm{a}^{(n)}$ to make $H\mathsf{F}(\bm{a}^{(n)})$ closer to $\bm{y}$ (Fig.
 4).
 Note that all parameters of $\mathsf{F}$ remain fixed because $\mathsf{F}$ is already trained.
-Since trained SR neural networks generally exhibit non-locality [30, 31, 32, 33], the backpropagation updates the analysis vector $\bm{a}$ from observations over distant grid points (Fig.
+Since trained SR neural networks generally exhibit non-locality,Dong et al.
+(2014); Gu and Dong (2021); Chen et al.
+(2023); Zhou et al.
+(2023) the backpropagation updates the analysis vector $\bm{a}$ using observations from distant grid points (Fig.
 4).
-We demonstrate that a trained SR model actually exhibits this non-locality in Section 6.2.
+We demonstrate that a trained SR model actually exhibits this non-locality in Section VI.2.
 Equation (17b) indicates that the differences from the super-resolved background state, $\mathsf{F}(\bm{a}^{(n)})-\mathsf{F}(\bm{x})$, also induce updates through backpropagation.
-The ELBO (16) determines the balance between the updates from observation and background values.
+The ELBO [Eq.
+(16)] determines the balance between the updates from observation and background values.
 Figure 4: Schematic of the non-local update by backpropagation through an SR neural network (i.e., a non-linear SR model).
-A large error (red) is propagated backward through yellow nodes in the hidden layers and updates the LR input values.
+A large error (red), $H\mathsf{F}(\bm{a}^{(n)})-\bm{y}$, propagates backward through yellow nodes in the hidden layers and updates the LR input $\bm{a}^{(n)}$ to make $H\mathsf{F}(\bm{a}^{(n)})$ closer to $\bm{y}$.
+See text for details.
 Minimizing Eq.
@@
 (17b).
-As in Section 3.3, we propose to solve the minimization problem using CVAEs.
-4.4 Loss function for SR models
+As in Section III.3, we propose to solve the minimization problem using CVAEs.
+IV.4 Loss function for SR models
 So far, we have used the trained SR model $\mathsf{F}$.
@@
 Typically, $P$ consists of interpolation and/or subsampling operations.
-The right-hand side shows that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies [27, 28, 29, 41].
-Backpropagation is feasible because $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$.
+The right-hand side shows that $l_{\text{SR}}$ is an L1 norm, which is frequently used in SR studies.Ha et al.
+(2019); Anwar, Khan, and Barnes (2020); Lepcha et al.
+(2023); Fukami, Fukagata, and Taira (2023) Backpropagation is feasible because $l_{\text{SR}}$ is differentiable with respect to $\mathsf{F}(\bm{x})$.
 The SR model $\mathsf{F}$ is trained using $\bm{x}$ as input and $\bm{y}$ as target.
-When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$ [49].
-This result suggests that the SR model becomes unbiased if the median of the observation error distribution is zero.
-4.5 Implementation of SRDA using CVAE
-An efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, as in Section 3.3, as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
+When the training data size is sufficiently large, the trained SR model infers the median of $\bm{y}$ conditioned on $\bm{x}$.Hastie, Tibshirani, and Friedman (2009) This suggests that the SR model becomes unbiased if the median of the observation error distribution is zero.
+IV.5 Implementation of SRDA using CVAE
+An efficient minimization of $l_{\text{SRDA}}$ in Eq. (13) is possible with CVAEs, as in Section III.3, as long as the statistical properties of $\bm{x}$ and $\bm{y}$ do not change significantly.
 The CVAE for SRDA consists of an encoder and a decoder (Fig.
@@
 2 is that the encoder incorporates an SR model and performs SR and DA simultaneously.
-We use the SR model in Section 4.4, which is trained with the loss function in Eq.
+We use the SR model in Section IV.4, which is trained with the loss function in Eq.
 (18).
 The encoder integrates the LR background $\bm{x}$ and the HR observations $\bm{y}$, estimating the HR analysis $\bm{a}_{\text{HR}}$ through computing the LR analysis $\bm{a}_{\text{LR}}$.
-This computation of $\bm{a}_{\text{LR}}$ [$=A_{\text{LR}}(\bm{x},\bm{y})$] emulates the result of iterative backpropagation through $\mathsf{F}$ (i.e., the result of minimizing $l_{\text{SRDA-NL}}$ in Section 4.3).
+This computation of $\bm{a}_{\text{LR}}$ [$=A_{\text{LR}}(\bm{x},\bm{y})$] emulates the result of iterative backpropagation through $\mathsf{F}$ (i.e., the result of minimizing $l_{\text{SRDA-NL}}$ in Section IV.3).
 The encoder also estimates the uncertainty of $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] by calculating the variance $V$ in Eq.
@@
 2.
-The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ as the loss function.
+The CVAE is trained in an unsupervised manner using $l_{\text{SRDA}}$ in Eq.
+(13) as the loss function.
 Figure 5: Architecture of the CVAE for SRDA.
@@
 The decoder acts as the observation operator and reconstructs the observation vector $\bm{\hat{y}}=\mathsf{H}(\bm{\hat{z}})$.
-Rigorously speaking, a different SR model should be used within the encoder because the SR model in Section 4.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$.
+Rigorously speaking, a different SR model should be used within the encoder because the SR model in Section IV.4 is trained to super-resolve the LR background $\bm{x}$, which differs from the LR analysis $\bm{a}_{\text{LR}}$.
 This difference in input may introduce bias in $\bm{a}_{\text{LR}}$.
-In contrast, the output $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] is expected to be less biased because the SR model is unbiased if the observation error distribution has a zero median (Section 4.4).
-Thus, $\bm{a}_{\text{HR}}$ is used as initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section 4.6).
+In contrast, the output $\bm{a}_{\text{HR}}$ [$=\mathsf{F}(\bm{a}_{\text{LR}})$] is expected to be less biased because the SR model is unbiased if the observation error distribution has a zero median (Section IV.4).
+Thus, $\bm{a}_{\text{HR}}$ is used as initial conditions for LR fluid models by resizing $\bm{a}_{\text{HR}}$ to the LR (Section IV.6).
 One solution for making $\bm{a}_{\text{LR}}$ less biased is to conduct fine-tuning of the SR model $\mathsf{F}$ during the training of the CVAE.
 For simplicity, this study does not use such an advanced learning method, which remains a future research topic; instead, the parameters of $\mathsf{F}$ are fixed during the CVAE training.
-4.6 Sequential SRDA using the CVAE
+IV.6 Sequential SRDA using the CVAE
 We introduce a sequential SRDA method applicable to CFD simulations.
 The background state $\bm{x}$ is hereafter referred to as the forecast state because it is a prior forecast given by an LR fluid model.
-The forecast and assimilation are executed alternately, as in other statistical DA methods [1, 2].
-This assimilation is performed using a trained CVAE’s encoder.
+The forecast and assimilation are executed alternately, as in other statistical DA methods.Asch, Bocquet, and Nodet (2016); Carrassi et al.
+(2018) This assimilation is performed using a trained CVAE’s encoder.
 Figure 6 shows a schematic of the sequential SRDA.
@@
 This SR operation is independent of assimilation, enabling the inference of HR forecasts at any time step.
-5 Methods
-We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs) [50], which are commonly used in testing DA methods for the atmosphere and ocean.
+IV.7 Related studies
+We first compare the present SRDA with previous SRDA methods.Barthélémy et al.
+(2022); Yasuda and Onishi (2023) The major differences are that the present SRDA is based on unsupervised learning and makes full use of the non-locality of SR.
+Barthélémy et al.Barthélémy et al.
+(2022) first integrated SR and DA, in which forecasts are super-resolved by a neural network trained via supervised learning, and then assimilated by an ensemble Kalman filter.
+Although this approach can be easily incorporated into existing DA frameworks due to the separation of SR and DA, the non-local nature of SR may not be fully exploited in the assimilation process.
+Our previous studyYasuda and Onishi (2023) proposed a different SRDA that performs end-to-end learning to jointly optimize SR and DA.
+This SRDA requires HR ground-truth data and does not provide a clear theoretical background for combining SR and DA.
+By contrast, the present SRDA does not rely on HR ground-truth data because it employs unsupervised learning based on the ELBO, and clarifies that the non-locality of SR benefits DA.
+Other DA studies also use ELBOs.Ghimire et al.
+(2017); Xie et al.
+(2023); Lafon, Fablet, and Naveau (2023) The most distinct point from these studies is that they do not utilize VAEs.
+Indeed, the use of ELBOs is generally distinguished from the use of VAEs.
+When VAEs are used, ELBOs or underlying probability distributions often need to be designed specifically for VAEs.Kingma et al.
+(2014); Sohn, Yan, and Lee (2015) In our case, the approximate posterior $q\left(\bm{z}\mid\bm{x},\bm{y}\right)$ is conditioned on both $\bm{x}$ and $\bm{y}$ in Eq.
+(2c), which is essential for the encoder to perform assimilation (Sections III.3 and IV.5).
+Thus, one of our contributions is to provide a basis for the combination of ELBOs and VAEs for DA.
+VAEs are already used for DA (see also Section I).
+Melinc and ZaplotnikMelinc, Boštjan and Zaplotnik, Žiga (2024) performed dimensionality reduction using a VAE and then applied the 3D-Var in the latent space.
+Glyn-Davies et al.Glyn-Davies et al.
+(2024) employed a dynamical VAE (DVAEGirin et al.
+(2021)) to encode time series of observational data, which were then assimilated using a Kalman filter.
+Xiao et al.Xiao et al.
+(2025) incorporated non-Gaussian background errors into the 3D-Var objective function by using the transformation with a VAE’s decoder.
+The key distinction of our approach is that the CVAE itself performs assimilation (i.e., integrating observations into forecasts) using the non-locality of SR.
+This integration can be understood as the update by backward propagation of discrepancies from observations (Section IV.3).
+Fundamentally, both DA and deep learning can be formulated using backpropagation with adjoint equations.Abarbanel, Rozdeba, and Shirman (2018); Chen et al.
+(2018); Geer (2021) Using this analogy, recent studies have developed neural networks that update forecasts by propagating errors from observations along time.Lafon, Fablet, and Naveau (2023); Fablet et al.
+(2021); Beauchamp et al.
+(2022) Our results reveal that SR models can act as error propagators for DA along the spatial dimensions, implying the possibility that neural networks act as propagators over the four dimensions (i.e., three spatial and one time dimensions).
+V Methods
+We evaluate the proposed SRDA method through numerical experiments similar to observing-system simulation experiments (OSSEs),Fletcher (2023) which are commonly used in testing DA methods for the atmosphere and ocean.
 The ground truth (or the nature run) is generated by an HR fluid model, from which synthetic observations are computed.
@@
 This jet system is suitable for assessing both aspects of SR and DA.
-5.1 Fluid simulation
-An idealized barotropic ocean jet, proposed by David et al.
-[51], was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi [35].
-The experimental setup is briefly explained here.
+V.1 Fluid simulation
+An idealized barotropic ocean jet, proposed by David et al.,David, Marshall, and Zanna (2017) was simulated in a two-dimensional periodic channel, using the same configuration as in Yasuda and Onishi.Yasuda and Onishi (2023) The experimental setup is briefly explained here.
 See the above references for further details.
-The governing equations in the periodic channel $(x,y)$ are as follows [51]:
+The governing equations in the periodic channel $(x,y)$ are as follows:
 $\frac{\partial\omega}{\partial t}+\bm{u\cdot\nabla}\omega+\beta\frac{\partial\psi}{\partial x}$ $=$ $-\kappa\omega-\nu\Delta^{2}\omega-\frac{d\tau(y)}{dy},$ (19a)
@@
 Equation (19a) describes the evolution of vorticity $\omega$ over time $t$.
-The left-hand side represents the advection of $\omega$ and the beta effect [52], while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$.
+The left-hand side represents the advection of $\omega$ and the beta effect,Vallis (2017a) while the right-hand side comprises linear drag, hyperviscosity, and forcing due to zonal wind stress $\tau(y)$.
 Equations (19b) and (19c) define the stream function $\psi$ and the velocity $\bm{u}$, respectively.
+All variables and parameters are non-dimensionalized following David et al.David, Marshall, and Zanna (2017)
 The parameters in Eq. (19a) were set as follows: $\beta=0.1$, $\kappa=1\times 10^{-2}$, and $\nu=1\times 10^{-5}$. The wind stress was given by
@@
 where $\tau_{0}=0.3$, $y_{0}=\pi/2$, $\delta=0.4$, and $c$ was adjusted such that the integral of $\tau(y)$ was zero.
-This setup falls within the parameter regime of mixing barriers with strong eddies [51], where coherent vortices persist in a statistically steady state (Section 6.1).
+This setup falls within the parameter regime of mixing barriers with strong eddies,David, Marshall, and Zanna (2017) where coherent vortices persist in a statistically steady state (Section VI.1).
 The initial condition was a zonal jet superimposed with random perturbations.
-The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion [53], which is a necessary condition for barotropic instability.
+The zonal jet adopts the same shape as $\tau(y)$ and satisfies the Rayleigh-Kuo inflection-point criterion,Vallis (2017b) which is a necessary condition for barotropic instability.
 Perturbations were added to the vorticity field for each wavenumber component, with phases and amplitudes drawn from Gaussian distributions.
@@
 Table 1: Fluid simulation configurations at two different spatial resolutions: LR and HR.
-The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule) [54].
+The cutoff wavenumber is the maximum wavenumber magnitude in the CFD model, which was determined by the specific condition to avoid aliasing errors in the pseudo-spectral method (i.e., the 2/3-rule).Canuto et al.
+(1988)
 Name Grid size ($x\times y$) Cutoff wavenumber Integration time step
@@
 HR (high resolution) $128\times 64$ 42 $\frac{1}{4}\times 5\times 10^{-4}$
-5.2 Synthetic observations
+V.2 Synthetic observations
 Synthetic observations were generated from the HR simulations.
@@
 To demonstrate that the SRDA can address changes in observation points, the subsampled points were varied in time by random shifts.
-For the unobserved points, Not a Numbers (NaNs) were assigned and later replaced with missing values in the preprocessing (Section 5.3).
+For the unobserved points, Not a Numbers (NaNs) were assigned and later replaced with missing values in the preprocessing (Section V.3).
 To emulate measurement errors, spatially independent Gaussian noise was added to the subsampled vorticity.
-The noise has a mean of 0 and a standard deviation of 0.1.
+This noise has a mean of 0 and a standard deviation of 0.1.
 This standard deviation is approximately 5% of the spatio-temporal average of the absolute values of vorticity.
-5.3 Neural networks
+V.3 Neural networks
 We outline the network architectures of an SR model and a CVAE for the sequential SRDA.
-Hyperparameters are provided in the Appendix, and the full implementation is available at our Zenodo repository (see Data Availability Statement).
-PyTorch 1.11.0 [55] was used for the implementation.
+Hyperparameters are provided in the Appendix, and the full implementation is available at our Zenodo repositoryYasuda and Onishi (2024) (see Data Availability Statement).
+PyTorch 1.11.0Paszke et al.
+(2019) was used for the implementation.
 The SR model architecture is shown in the top panel of Fig.
 7.
-This model resembles the neural network for SR known as VDSR, proposed by Kim et al.
-[56].
-The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR.
+This model resembles the neural network for SR known as VDSR, proposed by Kim et al.Kim, Lee, and Lee (2016) The input is the two-dimensional vorticity field obtained from the LR fluid model, namely the LR forecast state, which is super-resolved to the HR.
 First, this input is nonlinearly transformed through a stack of convolution and ReLU layers.
-Upsampling is then executed by two blocks that include pixel shuffle layers [57].
-Each block doubles the size of each dimension, resulting in a fourfold increase in resolution.
-Additionally, a skip connection [58] is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output.
+Upsampling is then executed by two blocks that include pixel shuffle layers.Shi et al.
+(2016) Each block doubles the size of each dimension, resulting in a fourfold increase in resolution.
+Additionally, a skip connectionHe et al.
+(2016) is employed, whereby the LR input, resized by bicubic interpolation, is added to the final output.
 This skip connection allows the SR model to capture the discrepancies between the LR input and the HR output.
@@
 The label “Res Block” refers to a residual block, and “Res Block x6” refers to the stacking of six Res Blocks.
-As in the typical implementation of VAEs [7], the output of variance is regarded as $\ln V$ because neural networks can generally output both positive and negative values.
-Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp(1/2\ln V)$ provides a realization of the true state $\bm{\hat{z}}$ in Eq.
+As in the typical implementation of VAEs,Kingma and Welling (2019) the output of variance is regarded as $\ln V$ because neural networks can generally output both positive and negative values.
+Consequently, $\bm{a}_{\text{HR}}+\bm{\varepsilon}^{\text{T}}\exp(1/2\log v)$ provides a realization of the true state $\bm{\hat{z}}$ in Eq.
 (14a).
@@
 7).
-The encoder performs SRDA and integrates the inputs, an LR forecast and an HR observation, into the HR analysis.
+The encoder performs SRDA and integrates the two inputs, an LR forecast and an HR observation, into the HR analysis.
 The HR observation is first downsampled through convolution to match its resolution with the LR (downsampled by a factor of 4).
-The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$.
+The effectiveness of separating forecast and observation inputs is shown in our previous study.Yasuda and Onishi (2023) The LR forecast and the downsampled observation are then concatenated and passed through residual blocks to obtain the LR analysis $\bm{a}_{\text{LR}}$.
 This $\bm{a}_{\text{LR}}$ is super-resolved by the trained SR model to the HR analysis $\bm{a}_{\text{HR}}$.
@@
 (12)] is also estimated by the encoder.
-Although $\bm{a}_{\text{HR}}$ and $V$ can be treated independently in the encoder (Section 4.5), we compute $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of parameters, using the feature extractor in the trained SR model (i.e., Feature Extractor-1 in Fig.
+Although $\bm{a}_{\text{HR}}$ and $V$ can be treated independently in the encoder (Section IV.5), we compute $V$ from $\bm{a}_{\text{LR}}$ to reduce the number of parameters, using the feature extractor in the trained SR model (i.e., Feature Extractor-1 in Fig.
 7).
 The decoder acts as an observation operator and reconstructs the observations from $\bm{\hat{z}}$.
-This operation comprises convolution and Leaky ReLU [59].
-We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), though this replacement slightly decreases the inference accuracy.
-This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section 5.2).
+This operation comprises convolution and Leaky ReLU.Maas et al.
+(2013) We confirmed that the decoder can be replaced with a subsampling operation (not a neural network), though this replacement slightly decreases the inference accuracy.
+This result is due to the experimental setup where the synthetic observations were generated by subsampling (Section V.2).
 In preprocessing, the vorticity $\omega$ at each grid point is transformed by
@@
 (21), zero is regarded as a missing value and is assigned to the HR grid points without observations.
-This clear interpretation of zero facilitates feature extraction from the input data [60].
-5.4 Training and testing methods for the neural networks
-5.4.1 Training method
-We adopted an offline strategy in which training data were pre-generated (Fig.
-8a).
-The training data consist of pairs of HR observations and LR forecasts, as discussed in Sections 4.4 and 4.5.
+This clear interpretation of zero facilitates feature extraction from the input data.Yasuda, Onishi, and Matsuda (2023)
+V.4 Training and testing methods for the neural networks
+V.4.1 Training method
+We adopted an offline strategy in which training data were pre-generated [Fig.
+8(a)].
+The training data consist of pairs of HR observations and LR forecasts, as discussed in Sections IV.4 and IV.5.
 The initial conditions for the LR fluid model were obtained by applying a low-pass filter to the HR simulation results in the wavenumber domain.
@@
 The corresponding observations were obtained by subsampling.
-Figure 8a denotes pairs of the LR forecasts and HR observations as stars.
-By varying the random initial perturbations (Section 5.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
+Figure 8(a) denotes pairs of the LR forecasts and HR observations as stars.
+By varying the random initial perturbations (Section V.1), we conducted 4,500 simulations; 70% were used as training data and the remaining 30% as validation data for hyperparameter tuning.
 Figure 8: Schematics of time series for (a) training and (b) testing phases.
@@
 Thus, the filtered HR results were used as the initial conditions.
-In realistic applications, these initial conditions can be obtained from existing assimilated data [61, 62], or online training would be effective [63, 64], where training and data generation are performed alternately or in parallel.
+In realistic applications, these initial conditions can be obtained from existing assimilated data,Hersbach et al.
+(2020); KOSAKA et al.
+(2024) or online training would be effective,Sahoo et al.
+(2018); Rasp (2020) where training and data generation are performed alternately or in parallel.
 The SR model was trained, followed by training the CVAE.
-For both trainings, the Adam optimizer [65] was used with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128.
+For both trainings, the Adam optimizerKingma and Ba (2015) was used with a learning rate of $1\times 10^{-4}$ and a mini-batch size of 128.
 Each training was terminated using early stopping with a patience parameter of 50 epochs.
 During the CVAE training, we fixed all parameters in the SR model.
+Note that VAEs typically show stable training and inference characteristicsKingma and Welling (2014) and we confirmed this behavior in our experiments (not shown).
 The loss functions of the SR model and the CVAEs are $l_{\text{SR}}$ in Eq.
@@
 Without this balancing, the performance of the CVAE would deteriorate.
-This phenomenon frequently occurs in VAEs [66, 67], and the loss balancing can be optimized during training [68].
-For simplicity, however, this advanced technique was not employed in our study.
-5.4.2 Testing method
-The sequential SRDA was tested using the method described in Section 4.6.
-Figure 8b shows a schematic of the test phase.
+This phenomenon frequently occurs in VAEs,Higgins et al.
+(2017); Alemi et al.
+(2018) and the loss balancing can be optimized during training.Asperti and Trentin (2020) For simplicity, however, this advanced technique was not employed in our study.
+V.4.2 Testing method
+The sequential SRDA was tested using the method described in Section IV.6.
+Figure 8(b) shows a schematic of the test phase.
 The only difference between Figs.
-6 and 8b is that observations are generated by subsampling the HR simulation results.
+6 and 8(b) is that observations are generated by subsampling the HR simulation results.
 Due to this difference, the HR results are regarded as the ground truth, and the SRDA is evaluated by comparing its inferences to the ground truth.
@@
 Two test metrics were used to assess pixel-wise accuracy and pattern consistency, both of which measure deviations from the ground truth.
-These two types of metrics are frequently used in SR studies [29] because they complement each other.
+These two types of metrics are frequently used in SR studiesLepcha et al.
+(2023) because they complement each other.
 The pixel-wise accuracy quantifies localized differences but may overestimate discrepancies when spatial patterns are correctly inferred with a slight shift.
@@
 where $\omega_{i}$ denotes the ground-truth vorticity at the $i$-th grid point, $\hat{\omega}_{i}$ is the corresponding inference, and the summation is taken over all grid points.
-The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns [69]:
+The mean structural similarity index measure (MSSIM) loss evaluates the consistency of spatial patterns:Wang et al. (2004)
 MSSIM loss $=$ $1-{\text{MSSIM}},$ (23)
@@
 The MSSIM loss takes a value greater than or equal to zero; a smaller value indicates spatial patterns more similar to the ground truth.
-A detailed discussion of MSSIM can be found in Wang et al.
-[69].
+A detailed discussion of MSSIM can be found in Wang et al.Wang et al.
+(2004)
 The MAE ratio and MSSIM loss were calculated at each time step in each test simulation.
@@
 Both MAE ratio and MSSIM loss are collectively called the test errors.
-5.5 Baseline model using Ensemble Kalman Filter (EnKF)
-As a baseline for comparison with the SRDA, we employed an ensemble Kalman filter (EnKF) [70] for the LR fluid model, using the perturbed observation method [71].
-The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations.
-The uncertainties estimated by the EnKF and SRDA are compared in Section 6.4.
+We discuss the results from the test data using these metrics in Section VI.
+V.5 Baseline model using Ensemble Kalman Filter (EnKF)
+As a baseline for comparison with the SRDA, we employed an ensemble Kalman filter (EnKF)Evensen (1994) for the LR fluid model, using the perturbed observation method.Burgers, van Leeuwen, and Evensen (1998) The choice of EnKF was motivated by its ability to estimate uncertainty through ensemble simulations.
+The uncertainties estimated by the EnKF and SRDA are compared in Section VI.4.
 Assimilation was conducted in the HR space after the LR forecast was super-resolved to the HR using bicubic interpolation, with the assimilation interval $\Delta T$ ($=1$) being the same as in the SRDA.
-Barthélémy et al.
-[34] argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points.
+Barthélémy et al.Barthélémy et al.
+(2022) argued that performing the EnKF in the HR space can yield more accurate inferences than those in the LR space when observations are defined at HR grid points.
 Furthermore, this approach clarifies the comparison with the SRDA because both EnKF and SRDA employ the LR fluid model and perform DA in the HR space.
-The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn [72].
-These error covariances were inflated by adding Gaussian noise to the analysis before the forecast process [73].
-This Gaussian noise incorporates spatial correlation by estimating the covariance from the training data at each time step.
+The background error covariance matrices were spatially localized using a function proposed by Gaspari and Cohn.Gaspari and Cohn (1999) These error covariances were inflated by adding Gaussian noise to the analysis before the forecast process.Whitaker et al.
+(2008) This Gaussian noise incorporates spatial correlation by estimating the covariance from the training data at each time step.
 The initial perturbations for creating ensemble members were also drawn from this Gaussian distribution.
@@
 Specifically, the following parameters were tuned: the number of ensemble members (set to 300 after tuning), the amplitude of the initial perturbation for generating ensemble members, the amplitude of the additive inflation, the localization radius, and the amplitude for perturbing observations.
-The full implementation of the EnKF is also available at our Zenodo repository (see Data Availability Statement).
-6 Results and discussion
-6.1 Time evolution of vorticity in the jet system
+The full implementation of the EnKF is also available at our Zenodo repositoryYasuda and Onishi (2024) (see Data Availability Statement).
+VI Results and discussion
+VI.1 Time evolution of vorticity in the jet system
 We first describe the time evolution of HR vorticity in the jet system and confirm the sensitivity to initial conditions (i.e., a positive Lyapunov exponent).
@@
 Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$).
-This final state is statistically steady and can be understood by Rossby wave dynamics [51].
+This final state is statistically steady and can be understood by Rossby wave dynamics.David, Marshall, and Zanna (2017)
 Figure 9: Typical vorticity evolution from the HR ground-truth data.
@@
 This indicates that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability.
-Without DA, the deviation from the ground truth continues to increase over time, as shown in Section 6.3.
-6.2 Accuracy of SR inference
+Without DA, the deviation from the ground truth continues to increase over time, as shown in Section VI.3.
+VI.2 Accuracy of SR inference
 We show the accuracy of inference by the SR model, as the SRDA is based on this model.
@@
 The input data are not indexed by time because each sample was randomly selected for testing without simulating time evolution.
-Flow-field reconstruction has been extensively studied using neural networks [74, 75, 76, 77].
-In these studies, inputs are incomplete, such as spatially sparse sensor measurements.
+Flow-field reconstruction has been extensively studied using neural networks.Lluís et al.
+(2020); Maulik et al.
+(2020); Gundersen et al.
+(2021); Wang et al.
+(2022) In these studies, inputs are incomplete, such as spatially sparse sensor measurements.
 Neural networks are then trained by comparing their outputs to complete target data.
@@
 This point-wise nature enables backpropagation even when the target data are missing.
-Furthermore, since convolution shares its kernel across space (known as weight sharing) [78], the efficient learning was realized using the small amount of non-missing observational data.
+Furthermore, since convolution shares its kernel across space (known as weight sharing),Aloysius and Geetha (2017) the efficient learning was realized using the small amount of non-missing observational data.
 Although the inference appears to be accurate, the SR model cannot adequately correct forecast errors.
@@
 The trained SR model makes non-local inference.
-This non-locality is essential for assimilating observations, as discussed in Section 4.3.
-We use two indicators for interpreting the SR model: gradients and local attribution maps (LAMs) [31].
-Gradients are expressed as $\partial\mathsf{F}(\bm{x})/\partial\bm{x}$, where $\mathsf{F}$ is the trained SR model.
-Such a simple gradient is sometimes not a good indicator for interpreting how a model output depends on its input due to nonlinearity [31, 79].
-LAM, proposed by Gu and Dong [31], combines gradient methods with path integral techniques for interpreting behaviors of nonlinear SR models.
+This non-locality is essential for assimilating observations, as discussed in Section IV.3.
+We use two indicators for interpreting the SR model: gradients and local attribution maps (LAMs).Gu and Dong (2021) Gradients are expressed as $\partial\mathsf{F}(\bm{x})/\partial\bm{x}$, where $\mathsf{F}$ is the trained SR model.
+Such a simple gradient is sometimes not a good indicator for interpreting how a model output depends on its input due to nonlinearity.Gu and Dong (2021); Sundararajan, Taly, and Yan (2017) LAM, proposed by Gu and Dong,Gu and Dong (2021) combines gradient methods with path integral techniques for interpreting behaviors of nonlinear SR models.
 Figure 11 shows gradient and LAM fields for the three SR samples previously discussed in Fig.
@@
 The gradients and LAMs indicate that colored points in the LR inputs contribute to the HR outputs in these boxes.
-If there are observed values in these boxes, the colored points in the inputs are updated by minimizing the loss function (Section 4.3).
+If there are observed values in these boxes, the colored points in the inputs are updated by minimizing the loss function (Section IV.3).
 Thus, observations have a non-local influence through the trained SR model.
@@
 10, but shown here for comparison.
-Local attribution maps (LAMs) [31] were calculated for the small square regions of $11\times 1$ 11 (i.e., black boxes in the figure).
+Local attribution maps (LAMs)Gu and Dong (2021) were calculated for the small square regions of $11\times 1$ 11 (i.e., black boxes in the figure).
 These boxes were specified on the HR grids, so they are not necessarily aligned with the LR grids.
-The gradients and LAMs were normalized by their maximum values, with ranges between 0 and 1, following Gu and Dong [31].
-6.3 Accuracy of LR forecast
+The gradients and LAMs were normalized by their maximum values, with ranges between 0 and 1, following Gu and Dong.Gu and Dong (2021)
+VI.3 Accuracy of LR forecast
 We demonstrate that the sequential SRDA can make more accurate inferences than the EnKF. The LR forecast is discussed here, and the HR analysis is examined next.
@@
 Figure 13 shows the time series of the mean values for the MAE ratio and MSSIM loss.
-Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (Section 6.1).
+Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (Section VI.1).
 In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines).
@@
 The bottom panels show the differences between the SRDA and EnKF, where the gray areas cover the 10th to 90th percentile values and the solid lines indicate the mean differences.
-In the proposed SRDA, the LR forecast can be super-resolved to the HR at an arbitrary time using the SR model (Section 4.6).
+In the proposed SRDA, the LR forecast can be super-resolved to the HR at an arbitrary time using the SR model (Section IV.6).
 Figure 14 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast.
@@
 Figure 14: Typical vortex merging in the ground truth, the LR forecasts, and the super-resolved forecasts.
-6.4 Accuracy of HR analysis
+VI.4 Accuracy of HR analysis
 Figure 15 shows snapshots of the ground truth and analyses.
@@
 Although the true value of SD is unknown, our results suggest that the SRDA tends to estimate larger SD in low-vorticity regions.
-This tendency may be due to a simplified variance estimation in the present CVAE (Section 5.3), where the variances are computed using the SR model from the LR analysis.
+This tendency may be due to a simplified variance estimation in the present CVAE (Section V.3), where the variances are computed using the SR model from the LR analysis.
 The SRDA demonstrated a potential for uncertainty quantification without relying on ensemble calculation; however, further model development is necessary.
-Recent studies demonstrate that VAEs can learn error distributions of model forecasts [22, 23, 24, 25].
-Their model architectures may be effective for the SRDA.
+Recent studies demonstrate that VAEs can learn error distributions of model forecasts.Grooms (2021); Yang and Grooms (2021); Grooms et al.
+(2023); Xiao et al.
+(2025) Their model architectures may be effective for the SRDA.
 Figure 17: Vorticity snapshots of the analyses and their standard deviations (SDs) for the EnKF and SRDA at $t=2$, 9 and 11. The analyses shown here are the same as in Fig. 15.
-6.5 Computation time of SRDA
+VI.5 Computation time of SRDA
 The computational time of the SRDA is much less than that of the EnKF.
@@
 This difference is mainly due to the calculation of the ensemble evolution, which is not required in the SRDA.
-7 Conclusions
-This study has proposed a theory of unsupervised SRDA with the CVAE.
-Using the ELBO, we have developed an unsupervised theory of DA, which is an extension of the 3D-Var (Section 3).
+VII Conclusions
+This study proposed a theory of unsupervised SRDA using the CVAE.
+By deriving an ELBO for DA, we showed that the theory is an extension of the 3D-Var (Section III).
 Since ELBOs generally serve as the loss function for CVAEs, the assimilation is feasible using the CVAE.
-In this theory, it is essential to use a non-diagonal background error covariance matrix, as in the 3D-Var.
-Instead of using such non-diagonal matrices, we have extended the theory by using the non-locality of SR to perform DA (Section 4).
-This extension provides a theoretical foundation for combining SR and DA.
-The combination of SR models and CVAEs enables us to assimilate HR observations into LR forecasts computed with LR fluid models.
-This approach eliminates the need for time integration of HR fluid models, thereby reducing computational burden.
-The effectiveness of the proposed method was evaluated via numerical experiments using an idealized barotropic ocean jet system (Sections 5 and 6).
+In this theory, it is essential to use the non-local SR operation, which makes background covariances learnable from data (Section IV).
+The proposed SRDA assimilates HR observations into LR forecasts without running HR fluid models.
+The effectiveness of the proposed method was evaluated via numerical experiments using an idealized barotropic ocean jet system (Sections V and VI).
 Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times.
-Our results suggest the potential of SRDA as a promising approach for integrating observations into CFD simulations while obtaining HR inferences with a low computational cost.
-There are several future research topics for SRDA.
+There are several future research directions for unsupervised SRDA.
+The major limitation of the proposed theory is the Gaussian assumption Eq.
+(10), which suggests that the accuracy in inference would be reduced if a system exhibited strong non-Gaussianity.
+One method to extend the theory to non-Gaussian systems is combining CVAEs with normalizing flowsRezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) to learn non-linear transformations that lead to non-Gaussian distributions.
 The proof of concept here has only examined the linear observation operator, although the proposed SRDA theoretically allows learning of nonlinear observation operators.
-Furthermore, advanced training methods, such as online training [63, 64] and loss balancing for VAEs [68], will be important to develop SRDA.
+Furthermore, advanced training methods, such as online trainingSahoo et al.
+(2018); Rasp (2020) and loss balancing for VAEs,Asperti and Trentin (2020) will be important to develop SRDA.
 As for generalizability, we have not evaluated to what extent the CVAE remains effective when the statistical properties of the input data differ.
@@
 It is necessary to evaluate the proposed SRDA for more complex systems.
-Declaration of competing interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Acknowledgements
+Acknowledgements.
 This work used computational resources of the TSUBAME3.0 supercomputer provided by the Institute of Science Tokyo through the HPCI System Research Project (Project ID: hp220102) to conduct the CFD simulations.
@@
 This paper is based on results obtained from a project, JPNP22002, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
-Appendix A Hyperparameters of the SR model and the CVAE
+Author Declarations
+Conflict of Interest
+The authors have no conflicts to disclose
+Data Availability Statement
+The data and source code that support the findings of this study are openly available at our Zenodo repository (https://doi.org/10.5281/zenodo.13958042), reference number Yasuda and Onishi, 2024.
+These files are also available at our GitHub repository (https://github.com/YukiYasuda2718/srda-cvae/releases/tag/v0.2.1).
+Appendix A Hyperparameters for the SR model and the CVAE
 We list the hyperparameters of the SR model and the CVAE shown in Fig.
@@
 The kernel size of all convolutions is 3.
-The slope of all leaky ReLU [59] is $-0.01$.
-The upscale factor of all pixel shuffle operations [57] is 2.
+The slope of all leaky ReLUMaas et al.
+(2013) is $-0.01$.
+The upscale factor of all pixel shuffle operationsShi et al.
+(2016) is 2.
 The resizing in the SR model is performed by bicubic interpolation.
-The weight parameters of all convolution layers are initialized randomly using uniform distributions [80].
+The weight parameters of all convolution layers are initialized randomly using uniform distributions.He et al.
+(2015) For the Adam optimizer, the default parametersKingma and Ba (2015) are used except for the learning rate of $1\times 10^{-4}$.
+The mini-batch size is set to 128 for all experiments.
 Table 2: Hyperparameters for the convolution layers in the SR model and the CVAE. The layer names are defined in Fig. 7.
@@
 Conv2D-17 256 256 1
-Data availability
-The data and source code that support the findings of this study are preserved and openly available at our Zenodo repository (https://doi.org/10.5281/zenodo.13958042).
-These files are also available at our GitHub repository (https://github.com/YukiYasuda2718/srda-cvae/releases/tag/v0.2.1).
-Declaration of Generative AI and AI-assisted technologies in the writing process
-During the preparation of this work, the authors used Claude 3.5 only for English editing.
-After using this service, the authors reviewed and edited the content as needed.
-The authors take full responsibility for the content of the publication.